### v2

# Non-Adaptive Multi-Stage Algorithm for Group Testing with Prior Statistics

Abstract
In this paper, we propose an efficient multi-stage algorithm for non-adaptive Group Testing (GT) with general correlated prior statistics. The proposed solution can be applied to any correlated statistical prior represented in trellis, e.g., finite state machines and Markov processes. We introduce a variation of List Viterbi Algorithm (LVA) to enable accurate recovery using much fewer tests than objectives, which efficiently gains from the correlated prior statistics structure. Our numerical results demonstrate that the proposed Multi-Stage GT (MSGT) algorithm can obtain the optimal Maximum A Posteriori (MAP) performance with feasible complexity in practical regimes, such as with COVID-19 and sparse signal recovery applications, and reduce in the scenarios tested the number of pooled tests by at least $25\%$ compared to existing classical low complexity GT algorithms. Moreover, we analytically characterize the complexity of the proposed MSGT algorithm that guarantees its efficiency.
I Introduction
Classical Group Testing (GT) aims to detect a small number of “defective” items within a large population by mixing samples into as few pooled tests as possible. The idea of GT was first introduced during World War II when it was necessary to discover soldiers infected with Syphilis. Dorfman [1] showed that the required number of tests could be reduced if multiple blood samples were grouped into pools. When the samples that participate in the next pool are selected iteratively based on the previous pool test results, the GT algorithm is called adaptive. In contrast, in non-adaptive GT, the whole process is designed in advance. Since it was first suggested, the GT problem has been investigated and generalized to many areas and applications, among them disease detection [2, 3, 4], cyber security applications [5], pattern matching algorithms [6] and communication [7, 8].
All of these applications imply a strong connection between GT and Compressed Sensing (CS) as two methods for sparse signal recovery that share common applications [9, 10, 11, 12]. The main difference between the two is that CS aims to recover a real-valued signal [13], while GT recovers a binary signal [14] or discrete-values [15, 16, 4]. Thus, one can consider GT as a Boolean CS [17, 18].
Traditional GT and its performance (i.e., the tradeoff between the number of pool tests and recovery algorithm complexity), focused on the probabilistic model in which the items are identically distributed and independent [19, 14]. Recent research explores cases where prior information about the correlation of objects is available [20, 2, 3, 21, 22]. The motivation for this approach arises from the fact that correlated prior statistics have the potential to achieve higher recovery rates and reduce the number of required tests. In disease detection, leveraging information about the proximity between individuals, represented by contact tracing information or graphs, can lead to significant savings in pool tests [20, 2, 3, 21]. However, previous GT works presented solutions designed for specific models and applications and may not be extended easily to other models and applications. In numerous signal processing applications, correlation between different frequencies, time signals, or among different sensors can also be utilized to achieve more precise estimations [23, 13, 24]. Hidden Markov Model (HMM) is a common model for many physical signals, such as speech signals [25], human motion [26], and spectrum occupancy in communication systems [27]. Infections can be also modeled as HMMs [28]. For example, [29] presents a GT solution for a specific HMM derived from contact tracing. To the best of our knowledge, no existing solution addresses the GT problem with general Markovian priors and applicable to a wide range of diverse applications.
In this work, we introduce a practical non-adaptive Multi-Stage GT (MSGT) algorithm for correlated items with prior statistics. The proposed multi-stage algorithm employs a new variation of the parallel List Viterbi Algorithm (LVA) [30, 31, 32] we designed for GT to enable accurate low complexity recovery using fewer tests. The proposed algorithm can be applied for any statistical prior represented in trellis [33], e.g., finite-state machines and Markov processes. Using LVA, MSGT leverages those statistics to estimate the defective set efficiently, even in a regime below the maximum likelihood (ML) upper bound. Furthermore, we show how the algorithm’s parameters can be tuned to achieve a maximum probability of success without exceeding the limitation of the available computational capacity. We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the MSGT computational complexity efficiency. Our numerical results demonstrate that in practical regimes for COVID-19 [34, 35] and sparse signal recovery in signal processing [9, 10, 11, 13], the low computational complexity MSGT algorithm proposed herein can reduce in the scenarios tested the number of pool tests by at least $25\%$.
The rest of this paper is organized as follows. Section II formally describes the GT model with correlated prior statistics. Section III presents the MSGT algorithm and the analytical results. Section IV details the simulation evaluation. Finally, Section V provides concluding remarks and future directions.
II Problem Formulation
Given a set of individuals $\mathcal{N}$, the objective in GT is to discover a small subset $\mathcal{K}$ of unknown defective items using the minimum number of measurements $T$. Let $N=|\mathcal{N}|$, $K=|\mathcal{K}|$ denote the total number of items and the number of defective items, respectively, where $K=\mathcal{O}(N^{\theta})$ for some $\theta\in\left[0,1\right)$. The binary vector $\mathbf{U}\in\left\{0,1\right\}^{N}$ represents the population, such that $U_{i}=1$ indicates that the $i$-th item is defective. We assume that the set of the individuals is sparse, such that $\theta\leq 1/3$[17, 19, 36, 14], and that each item can be in one of 2 states: defective and not-defective. For each item, there is some prior probability that it is defective, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}\in\left[0,1\right]^{N}$, $\pi_{i}=P\left(U_{i}=1\right)$. The correlation between the state of the $i$-th item and the states of the previous $\tau$ items is represented by $\mathbf{\Phi}_{i}\in\left[0,1\right]^{2^{\tau}\times 2}$. $\tau=1,2,\dots$ represents the number of memory steps of the underlying process considered in the recovery stage. When there is no memory in the process, $\tau=0$ and $\{\pi_{i}\}_{i=1}^{N}$ hold all the prior information. $\mathbf{\Phi}_{i}[l,k]=P\left(U_{i}=s_{k}\lvert\left(U_{i-\tau},...,U_{i-1}\right)=s_{l}\right)$, where $s_{l}$ is the binary representation of $l$ by a $\tau$-length binary vector and $s_{k}\in\{0,1\}$ (For example see Fig. 6(b)).
For the non-adaptive GT, the testing matrix $\mathbf{X}\in\left\{0,1\right\}^{T\times N}$ is defined such that each row corresponds to a single pool test, and each column corresponds to a single item. That is, the $i$-th pool test is represented as a binary row vector: $X_{i}=\left[X_{i}(1),...,X_{i}(N)\right],\,i\in\left\{1,...,T\right\}$ whose elements are defined: $X_{i}(j)=1$ if the item with an index $j\in\left\{1,...,N\right\}$ is included in the $i$-th pool test, and otherwise $X_{i}(j)=0$. Then, the outcome of the $i$-th pool test is given by
$Y_{i}=\bigvee_{j\in\mathcal{K}}X_{i}(j)=\bigvee_{j=1}^{N}X_{i}(j)U_{j}$
where $\bigvee$ is the Boolean OR function.
Given $\mathbf{X}$ and the outcome vector $Y$, the recovery success criterion in GT can be measured using various metrics [14]. The main metrics we will use herein are exact recovery and partial recovery. In terms of exact recovery, the goal is to detect the precise subset of defective items $\mathcal{K}$. Therefore, given the estimated defective set $\hat{\mathcal{K}}=\hat{\mathcal{K}}(N,K,\mathbf{X},Y)$, we define the average error probability by [^1] For simplicity of notation, $P_{s}$ and $P_{e}=1-P_{s}$ refer to success and error probabilities in the exact recovery analysis.
$P_{e}^{exact}\triangleq\frac{1}{\binom{N}{K}}\sum_{\mathcal{K}:\left|\mathcal{K}\right|=K}P\left(\hat{\mathcal{K}}\neq\mathcal{K}\right).$
In partial recovery, we allow a false positive (i.e., $|\hat{\mathcal{K}}\setminus\mathcal{K}|$) and false negative (i.e., $|\mathcal{K}\setminus\hat{\mathcal{K}}|$) detection rate. Thus, the average partial success rate is given by:
$P_{s}^{partial}\triangleq\frac{1}{\binom{N}{K}}\sum_{\mathcal{K}:\left|\mathcal{K}\right|\neq K}\frac{\left|\hat{\mathcal{K}}\cap\mathcal{K}\right|}{K}.$
To conclude, for a sparse subset of infected items of size $K$, form $N$, the goal in non-adaptive GT with correlated prior data items, is to design a $T\times N$ testing matrix and an efficient and practical recovery algorithm that can exploit correlated priors, such that by observing $Y^{T}$ we can identify the subset of infected items with high probability and with feasible computational complexity. Thus, given the knowledge of the correlated prior data items and the available computational resources, the test designer could design the testing matrix and a recovery algorithm to maximize the success probability.
III Main Results
In this section, we introduce the efficient multi-stage recovery algorithm for any statistical prior represented in a trellis diagram [33], detailed in Algorithm 1. In the first stage, standard low-complexity algorithms [19] reduce the search space independently of prior correlations. This reduction is guaranteed by new analytical results we derive. In the second stage, the algorithm employs a novel adaptation of the List Viterbi Algorithm (LVA) [32], designed for GT to enable accurate low-complexity recovery using fewer tests by exploiting the correlated prior information. Additionally, we derive a bound to ensure the low complexity of the entire algorithm. Section III-A describes the proposed algorithm. Section III-B provides analytical results, followed by a discussion in Section  III-C. For a detailed explanation of all the algorithms used as integral components of Algorithm 1, see Appendix -A.
III-A Pool-Testing Algorithm
III-A1 Testing Matrix and Pooling
The proposed multi-stage recovery algorithm is intended to work with any non-adaptive testing matrix, e.g., as given in [37]. To simplify the technical aspects and focus on the key methods, the testing matrix is generated randomly under a fixed optimal approximation with Bernoulli distribution of $p=\ln(2)/K$ [38], using classical GT methods. The pooling and its outcome are given by the process elaborated in Section II and illustrated in Fig. 1.
Figure 1: For an unknown population $\mathbf{U}\in\left\{0,1\right\}^{9}$ with $K=2$, a random testing matrix is sampled and the test result $Y$ is calculated.
Algorithm 1 Multi-Stage Recovery Algorithm
Input: $\mathbf{X},Y,\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N},K,L,\tau,\gamma$
Output: $\hat{\mathcal{K}}$
1:Stage 1: Reduction of space search
2:$\mathcal{P}^{(DND)}\leftarrow\mathbf{DND}(\mathbf{X},Y)$
3:$\mathcal{P}^{(DD)}\leftarrow\mathbf{DD}\left(\mathbf{X},Y,\mathcal{P}^{(DND)}\right)$
4:$\mathcal{P}^{(S_{1,2})}\leftarrow\mathcal{N}\setminus\left(\mathcal{P}^{(DND)}\cup\mathcal{P}^{(DD)}\right)$
5:Stage 2: Recovery exploiting prior info
6:$\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}\leftarrow\mathbf{{updatePriors}}\left(\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N},\mathcal{P}^{(S_{1,2})},\mathcal{P}^{(DD)}\right)$
7:$\mathbf{Z}\leftarrow\mathbf{LVA}\left(L,\tau,\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}\right)$
8:$\mathbf{C}\leftarrow\{\}$
9:for $l\leftarrow 1$ to $L$ do
10:    if $K\leq\sum_{i}\left(\mathbf{Z}_{l,i}\right)\leq\gamma K$ then
11:         $\mathbf{V}^{(l)}\leftarrow\{i\mid\mathbf{Z}_{l,i}=1\}$
12:         $\mathbf{C}\leftarrow\mathbf{C}\cup\mathbf{getAllCombinations}\left(\mathbf{V}^{(l)},K\right)$
13:    end if
14:end for
15:$\hat{\mathcal{K}}\leftarrow\mathbf{MAP}\left(\mathbf{X},Y,\mathbf{C},\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}\right)$
III-A2 Recovery Process
The suggested recovery algorithm operates in two main stages. In the first stage (Stage 1), to reduce the space of search (i.e., the possible defective items), the algorithm efficiently identifies non-defective and definitely defective items without considering the prior correlated information. In the first step of this stage, we use the Definitely Not Defective (DND) algorithm [39, 40, 41, 19] (line 2, Fig. 2.(a)). DND compares the columns of the testing matrix, $\mathbf{X}$, with the outcome vector, $Y^{T}$. If $Y(i)=0$ for some $i\in\{1,\ldots,T\}$, the algorithm eliminates all items participating in the $i$-th test from being defective, and outputs them as the set $\mathcal{P}^{(DND)}\subset\mathcal{N}$.
In the second step of Stage 1, we use Definite Defectives (DD) algorithm [19] (line 3, Fig. 2.(b)), which goes over the testing matrix and the test result, looking for positive pool tests that include only one possibly defective item. DD denotes those items as definitely defective items and outputs them as the set $\mathcal{P}^{(DD)}$.
Let $\mathcal{P}^{(S_{1,1})}=\mathcal{N}\setminus\mathcal{P}^{(DND)}$ and $\mathcal{P}^{(S_{1,2})}=\mathcal{N}\setminus\left(\mathcal{P}^{(DND)}\cup\mathcal{P}^{(DD)}\right)$ denote the set of items that their status is still unclear after the first step and the second step, respectively. $\mathcal{P}^{(S_{1,2})}$ holds a new space search, and $\mathcal{P}^{(DD)}$ holds the already known defectives. This knowledge is acquired without utilizing any prior data, which we reserve for the second stage.
Figure 2: First stage of MSGT. (a) The first step of Stage 1, the DND algorithm, reveals 5 DND items in $\mathbf{U}$, forming $\mathcal{P}^{(DND)}$. Since items participating in negative tests must be non-defective, we mark all the participants in the two negative test results as non-defective. (b) The second step of Stage 1, the DD algorithm, outputs $\mathcal{P}^{(DD)}$ that includes a single DD item, based on the first test result, as it is the only possibly defective item participating in this test. The two other positive test results do not contribute to our knowledge here because there is more than one possibly defective item participating in them.
In the first step of Stage 2, we translate the data we obtained in DND and DD, $\mathcal{P}^{(S_{1,2})}$ and $\mathcal{P}^{(DD)}$, into the state space in terms of transition probabilities, $\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$, and initial probabilities, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}$, so we can employ all the gathered information in the next steps (line 6, Fig. 3.(a)). In the state space, the population sequence, $\mathbf{U}$, is parallel to time steps considered traditionally in VA for communications problems [32], and there are two possible states per item, the first indicates “non-defective” and the second indicates “defective”.
In the second step, the suggested LVA for GT goes over the sequence of items and outputs $\mathbf{Z}\in\{0,1\}^{L\times N}$, a list of the $L$ most likely trajectories in the state space (line 7, Fig. 3.(b)) according to MAP decision based on the given prior information. Each trajectory is a sequence of $N$ states, representing items classified as either defective or non-defective. Thus, $\mathbf{Z}$ provides $L$ estimations of $\mathbf{U}$. In practice, the $L$ estimations may include any number of defective items and require further processing.
In the third step, we extract candidates for the defective set out of the $L$ estimated sequences $\mathbf{Z}$ (lines 9 to 14). For some $l\in\{1,\ldots,L\}$, let $\mathbf{V}^{(l)}$ denote the set of items estimated as defective in $\mathbf{Z}_{l}$, the $l$-th row of $\mathbf{Z}$ (line 11). We ignore sequences that contain less than $K$ defective items or more than $\gamma K$ defective items, for some $\gamma\geq 1$, and consider only $\mathbf{Z}_{l}$ in which $K\leq\left|\mathbf{V}^{(l)}\right|<\gamma K$ as valid sequences. For each one of the valid sequences, we refer to all the combinations of size $K$ in $\mathbf{V}^{(l)}$ as candidates for $\hat{\mathcal{K}}$, and add them to the candidates list $\mathbf{C}$ (line 12).
At this point, we have in $\mathbf{C}$ a list of candidates to be our final estimation $\hat{\mathcal{K}}$, and we can calculate the probability of each one of them using $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$. Then, in the fourth step, the estimated defective set, $\hat{\mathcal{K}}$, is finally chosen using MAP estimator out of the $\mathbf{C}$ (line 15), i.e., $\hat{\mathcal{K}}=\arg\max_{c\in\mathbf{C}}{P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)}$.
If there are no valid sequences in $\mathbf{Z}$, we consider trajectories with fewer than $K$ detections for partial recovery. We select the trajectory with the most detections and randomly complete it to form a set of size $K$ for our final estimation $\hat{\mathcal{K}}$.
Figure 3: Stage 2 of MSGT. (a) All the possible transitions in the state space that we consider in the LVA step, following the insights obtained in Stage 1. These transitions aggregate to a total of 6 trajectories. (b) The two most likely trajectories returned by LVA (assuming $L=2$). Given $K=2$, the black trajectory corresponds to a valid population vector $\mathbf{U}$ with 2 defective items, while the gray trajectory indicates an invalid population with 3 defective items instead. Consequently, in the subsequent step, MSGT will extract two optional defective sets: $\left\{U_{6},U_{8}\right\}$ and $\left\{U_{7},U_{8}\right\}$, and will finally choose the most likely one using MAP estimator. (c) Comparison of Stage 2 to ML. With $T=5$, we use the first 5 rows of the testing matrix, ignoring the last test result. This leaves 3 possibly defective items, forming two potentially defective sets of size $K=2$. Using ML, one set is chosen randomly, leading to an error probability of 0.5. With $T=6$, based on the third and sixth test results, there is only one set of size $K=2$ that matches the outcome $Y$, resulting in successful decoding with the ML decoder. As shown above, MSGT’s Stage 2 can successfully decode $\mathbf{U}$ with just $T=5$, as using the LVA step it narrows down to only 2 possible trajectories, and then the final estimation is selected based on the given prior information and the insights gained in Stage 1.
(a) Upper bound for possibly defective items after DND for $N=10000$, and $K=15$.
(b) Lower bound for definitely defective items after DD for $N=10000$, and $K=15$.
(c) Minimum $\gamma$ parameter to satisfies (3) for $N=500$, and $K=3$.
Figure 4: Numerical evaluation for theoretical results and bounds. The results in (a), (b), and (c) are over 1000 iterations. For ML Upper Bound (UB), $T_{ML}=(1+\epsilon)K\log_{2}N$, for any $\epsilon>0$ [17]. In particular, $\epsilon=0.25$ in the results presented herein.
III-B Analytical Results
Here we provide analytical results related to the proposed MSGT.
Let $P_{e,a}^{(DND)}$ and $P_{e,u}^{(DND)}$ denote the error probability of DND in the average case and its deviation from the average (“worst-case”) that still allows the success of MSGT on average in feasible computational complexity, respectively. The two following theorems give the upper bound on the expected number of possible defective items and a “worst-case” upper bound after the first step of Stage 1.
Theorem 1 ([4]).
Consider a group test with a Bernoulli testing matrix with $p=\ln{2}/K$, and $T$ tests as $K\rightarrow\infty$. Let $P_{e,a}^{(DND)}\triangleq N^{-\alpha\left(1-\ln{2}/K\right)/2}$ for $\alpha\triangleq T/K\log_{2}{N}$. The expected number of possibly defective items is bounded by
$\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]\leq K+\left(N-K\right)P_{e,a}^{(DND)}.$
Proof.
According to [42, Lemma 2]:
$\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]=K+\left(N-K\right)P_{e}^{(DND)},$ (1)
where $P_{e}^{(DND)}=\left(1-p\left(1-p\right)^{K}\right)^{T}$ is the probability of non-defective item to be hidden by another defective item during DND.
By substituting $T=\alpha K\log_{2}{N}$, $p=\ln{2}/K$ we bound the error probability as follows:
$P_{e}^{\left(DND\right)}=\left(1-\frac{\ln{2}}{K}\left(1-\frac{\ln{2}}{K}\right)^{K}\right)^{\alpha K\log_{2}{N}}$
$=\left(1-\frac{\ln{2}}{K}\left(1-\frac{\ln{2}}{K}\right)^{\left(K-1\right)}\left(1-\frac{\ln{2}}{K}\right)\right)^{\alpha K\log_{2}{N}}$
$\leq\left(1-\frac{\ln{2}}{K}e^{-\ln{2}}\left(1-\frac{\ln{2}}{K}\right)\right)^{\alpha K\log_{2}{N}}$
$=e^{-\frac{\alpha}{2}\ln{2}\left(1-\frac{\ln{2}}{K}\right)\log_{2}{N}}$
$=N^{-\frac{\alpha}{2}\left(1-\frac{\ln{2}}{K}\right)}$
Finally, we denote this bound on the average error probability in DND by $P_{e,a}^{(DND)}\triangleq N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)}$. Hence, substituting $P_{e,a}^{(DND)}$ in (1), we have:
$\mathbb{E}\left[\left|\mathcal{P}^{(DND)}\right|\right]\leq K+\left(N-K\right)P_{e,a}^{(DND)}.$
This completed the theorem proof. ∎
Theorem 2.
Consider a group test with a Bernoulli testing matrix and DND decoder. As $K\rightarrow\infty$, the worst-case error probability of DND is bounded by
$P_{e,u}^{(DND)}$ $\leq P_{e,a}^{(DND)}$
$+\left(1+\frac{\left(\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)\right)^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1},$
for any $\gamma\geq 1$.
Proof.
Let $G$ denote the number of non-defective items that were hidden and not detected in DND (i.e., the number of false positive items). We upper bound the probability of error in DND, by the sum of the average probability of error $P_{e,a}^{(DND)}$ and a concentration term as follows:
$P_{e,u}^{(DND)}\triangleq P_{e,a}^{(DND)}+P\left(G-\mathbb{E}[G]>g\right),$
for some $g>0$.
Although the variation of LVA in Stage 2 applies a maximum likelihood decision, it’s performance in the proposed algorithm, may be worse than a brute force MAP’s performance, if the number of occlusions in the test matrix exceeds the threshold $\gamma K$. For this reason we are interested in $g=\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)$.
Recall that the probability of each definitely defective item to be hidden is given by the probability of error in DND, $P_{e}^{(DND)}$, which depends only on the matrix design. Thus, in the case of Bernoulli encoder, the $N-K$ random variables which represent occlusions of non-defective items are independent. It follows that $G\sim Bin\left(N-K,P_{e}^{(DND)}\right)$, and we can apply the one-sided Chebyshev’s inequality [43] to bound the concentration term:
$P\left(G-\mathbb{E}\left[G\right]>g\right)\leq\\ \left(1+\frac{g^{2}}{\left(N-K\right)P_{e}^{(DND)}\left(1-P_{e}^{(DND)}\right)}\right)^{-1}$
For a sufficiently large $T$, $P_{e}^{(DND)}\rightarrow P_{e,a}^{(DND)}$, which concludes the proof. ∎
The following theorem provides the expected defective items detected by DD in the second step of Stage 1.
Theorem 3.
Consider a group test with a Bernoulli testing matrix with $p=\ln{2}/K$, and $T$ tests as $K\rightarrow\infty$. The expected number of defective items that DD successfully detects is bounded by
$\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]\geq K\left(1-N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}\right).$ (2)
Proof.
The detection of the $i$-th defective item ($i\in\mathcal{K}$) in the DD algorithm fails under two conditions: either it does not participate in any test, or it participates but at least one other potentially defective item is also participating. Let $A_{i}$ represent the event where at least one potentially defective item, excluding item $i$, participates in some test. Consequently, $P\left(A_{i}\right)=1-\left(1-p\right)^{\left|\mathcal{P}^{(S_{1,1})}\right|-1}$. The probability of not identifying item $i$ as definitively defective in a given test $t\in\{1,...,T\}$ is given by $1-p+pP\left(A_{i}\right)$. So the probability of not detecting a defective item in all tests is given by
$P_{e}^{(DD)}$ $=\left[1-p+p\left(1-\left(1-p\right)^{\left|\mathcal{P}^{(S_{1,1})}\right|-1}\right)\right]^{T}$
$=\left[1-p\left(1-p\right)^{\left|\mathcal{P}^{(S_{1,1})}\right|-1}\right]^{T}.$
By substituting (1) we have:
$\mathbb{E}\left[P_{e}^{(DD)}\right]$
$\leq\left(1-p\left(1-p\right)^{K+\left(N-K\right)P_{e,a}^{(DND)}-1}\right)^{T}$
$\leq\left(1-p\left(1-p\right)^{K+P_{e,a}^{(DND)}N-1}\right)^{T}$
=$\underset{(a)}{=}\left(1-\frac{\ln{2}}{K}\left(1-\frac{\ln{2}}{K}\right)^{K-1}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}\right)^{T}$
$\underset{(b)}{\leq}\left(1-\frac{\ln{2}}{K}e^{-\ln{2}}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}\right)^{\alpha K\log_{2}{N}}$
$\underset{(c)}{\leq}exp\left(-\alpha e^{-\ln{2}}\ln{2}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}\log_{2}{N}\right)$
$=N^{-\frac{\alpha}{2}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}},$
where in $(a)$ we substitute $p=\ln{2}/K$, in $(b)$ we substitute $T=\alpha K\log_{2}{N}$ and used the fact that $e^{-\ln 2}\leq\left(1-\ln{2}/n\right)^{n-1}$ for any integer $n>0$. $(c)$ follows since $e^{-x}\geq\left(1-x/n\right)^{n}$ for any integer $n>0$ and $x>0$.
Subsequently, the probability of success satisfies is given by
$\mathbb{E}\left[P_{s}^{(DD)}\right]=1-\mathbb{E}\left[P_{e}^{(DD)}\right]\geq 1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}.$
Thus, an upper bound on the expected number of defective items that the DD algorithm successfully detects can be established as follows:
$\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]=\mathbb{E}\left[P_{s}^{(DD)}\right]K\geq\\ \left(1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}\right)K.$
This completed the theorem proof. ∎
The proposition below provides a sufficient condition on $\gamma$, such that, MSGT achieves performance that we conjecture outperforms ML. This can be assumed since LVA used for this step is an optimal MAP estimator [32].
Proposition 1.
Let $\eta$ denote the average success probability of ML decoder. For GT with correlated prior information and $\gamma\geq 1$, MSGT can achieve an average success probability equal or greater than $\eta$, if
$T\geq\log{\eta}+\log{\binom{N}{K}}-\log{\binom{N-K}{K(\gamma-1)}}.$ (3)
Proof.
We follow the lines of the success probability analysis in [36, Theorem 1]. Consider a list decoder that, given the test outcome $Y$, outputs $\mathcal{L}(Y)$, a list of items estimated as defective, such that $\lvert\mathcal{L}(Y)\lvert\geq K$. Let $N(Y)$ denote the number of possible defective sets for which, given $Y$, the list decoding is successful: $N(Y)\triangleq\sum_{\mathcal{K}}\mathbb{1}_{\{\mathcal{L}(Y)\cap\mathcal{K}=K\}}$. Assuming no prior information is given, $N(Y)$ is bounded by the counting bound
$N(Y)\leq\binom{N-K}{L-K}.$ (4)
The success probability of the decoder can be written as
$P_{s}=\frac{1}{\binom{N}{K}}\sum_{Y}N(Y)=\frac{2^{T}}{\binom{N}{K}}N(Y).$ (5)
By substituting (4) into (5), taking the logarithm and rearranging, we have that
$T\geq\log{P_{s}}+\log{\binom{N}{K}}-\log{\binom{N-K}{L-K}}.$
It is important to note that unlike [36, Theorem 1], we halt the success probability analysis here without deriving a converse, and instead focus on the connection between $T$ and $\gamma$. This analysis is valid for any list decoder, particularly for an ML decoder [17, 44], allowing us to set $P_{s}=\eta$. Since MSGT uses MAP, the non-uniform and correlated prior information allows us to achieve better recovery performance for the same $T$ and $L$ compared to ML. Moreover, since for any $\gamma\geq 1$, $\gamma K$ is the maximum list size of estimated defective items we allow in MSGT. By substituting $L=\gamma K$, we establish a lower limit for $\gamma$ that guarantees MSGT’s success probability remains at least equal to that of the ML estimator. ∎
One of the key features of the proposed MSGT algorithm is its low and feasible complexity in practical regimes compared to ML or MAP-based GT decoders. Both ML and MAP involve exhaustive searches, resulting in a complexity of $\mathcal{O}(\binom{N}{K}KN\log_{2}N)$ operations [17]. The theorem below and the subsequent Remark characterize the computational complexity of MSGT.
Theorem 4.
Consider a group test for a population of size $N$ items out of which $K$ are defective and a Bernoulli testing matrix. The computational complexity of the MSGT algorithm is bounded by $\mathcal{O}\left(L\gamma^{K}KN\log_{2}N\right)$ operations.
Proof.
We begin by analyzing the complexity of each step of the proposed MSGT solution given in Algorithm 1, and finally, we sum everything up to determine the total complexity.
In Stage 1, the complexity of DND is $\mathcal{O}\left(KN\log_{2}{N}\right)$ as analyzed in [45, Remark 6]. Then, for each positive entry of the test result vector $Y^{T}$, the DD algorithm counts the number of possibly defective items that participate in the corresponding pool test. That requires $K\left|\mathcal{P}^{(S_{1,1})}\right|\log_{2}{N}$ computations, and for simplification, we bound it by the DND complexity, i.e., $\mathcal{O}\left(KN\log_{2}N\right)$.
(a) $N=500$ and $K=3$.
(b) $N=1024$ and $K=8$.
(c) $N=10000$ and $K=13$.
Figure 5: Success probability of MSGT, MAP, ML and DD over 1000 iterations. A comparison to ML and MAP is not presented in (b) and (c), as they are not feasible for populations of those sizes due to the computational complexity burden.
In Stage 2, parallel LVA requires $L$ times more computations than the VA [32]. VA calculates all the possible transition probabilities for each step in the sequence. In GT, this sequence is the order’s items sequence, where with the suggested algorithm, it is enough to consider only the $\left|\mathcal{P}^{\left(S_{1,2}\right)}\right|$ items as the sequence steps. The optional states are basically either “non-defective” or “defective”, so there are four possible transitions in each step of the trellis proposed herein for GT. Nevertheless, this algorithm can be implemented to leverage additional memory to decide the state of each item based on the preceding $\tau$ items. Consequently, LVA takes $2^{2\tau}L\left|\mathcal{P}^{(S_{1,2})}\right|$ computations. For the average case, we use the expectation bounds from Theorems 1 and 3, such that:
$\mathbb{E}\left[\left|\mathcal{P}^{\left(S_{1}\right)}\right|\right]=\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]-\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]$
$\leq\left(N-K\right)P_{e,a}^{(DND)}+N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}K.$
Accordingly, the number of the required computations for LVA step in MSGT is bounded by
$2^{2\tau}L\left[\left(N-K\right)P_{e,a}^{(DND)}+N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}K\right].$ (6)
For the worst-case, we assume that the DD step does not affect the possible detected items set, thus $\left|\mathcal{P}^{\left(S_{1,2}\right)}\right|\leq\left|\mathcal{P}^{(S_{1,1})}\right|$. To further bound the expression, we utilize both the error probability of the worst case as given in Theorem 2, thus $\left|\mathcal{P}^{\left(DND\right)}\right|\leq K+(N-K)P_{e,u}^{(DND)}$. By substituting Theorem 2 it follows that the number of possibly defective items in the worst-case is upper bounded by
$\begin{split}K&+\left(N-K\right)\left(P_{e,a}^{(DND)}+\right.\\ &\left.\left(1+\frac{g^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1}\right),\end{split}$ (7)
for $g=\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)$. Now, since the two expressions multiplying $(N-K)$ in (6) and (7) are error probabilities, both of them can be roughly bounded by one. Hence, the number of computations of LVA in the average case and worst-case becomes $\mathcal{O}(LN)$.
In the next step in Stage 2, we filter the LVA results. We sum each sequence $\mathbf{Z_{l}}$ with a complexity of $\mathcal{O}(N)$. If the result is in the range $\left[K,\gamma K\right]$, we extract all the combinations of size $K-\left|\mathcal{P}^{(DD)}\right|$. Thus, this stage is done in at most $\mathcal{O}\left(L\left(N+\binom{\gamma K}{K-\left|\mathcal{P}^{(DD)}\right|}\right)\right)$ computations.
Finally, in the MAP step of Stage 2, the algorithm goes over at most $L$ combinations of size $K-\left|\mathcal{P}^{(DD)}\right|$ out of no more than $\gamma K$ possibly defective items in each sequence. Then, for each combination, the group test is applied. Therefore, the MAP stage requires $\binom{L\gamma K}{K}KN\log_{2}{N}$ computations. Substituting the bounds of the binomial coefficient $\left(\frac{\gamma K}{K}\right)^{K}\leq\binom{\gamma K}{K}\leq\left(\frac{e\gamma K}{K}\right)^{K}$, it follows that the complexity of MAP stage in the proposed MSGT is $\mathcal{O}\left(L\gamma^{K}KN\log_{2}{N}\right)$.
To conclude, the complexity of MSGT is the sum of all the steps, i.e., $\mathcal{O}\big{(}NK\log_{2}{N}+LN+L\binom{\gamma K}{K-\left|\mathcal{P}^{(DD)}\right|}+L\gamma^{K}KN\log_{2}{N}\big{)}$. As $N$ grows, the dominant term is the complexity of the MAP step. Thus, the complexity of the MSGT algorithm is bounded by $\mathcal{O}\left(\gamma^{K}KN\log_{2}{N}\right)$ operations, which completes the proof. ∎
Remark 1.
If we skip the LVA step, MSGT converges to the MAP estimator. Thus, the MAP’s complexity is
$\mathcal{O}\left(\displaystyle\binom{\left|\mathcal{P}^{(S_{1,2})}\right|}{K-\left|\mathcal{P}^{(DD)}\right|}KN\log_{2}N\right),$
when the DND and DD are executed as prior steps, and otherwise it is $\mathcal{O}\left(\binom{N}{K}KN\log_{2}N\right)$.
Note that from Theorem 4 and Remark 1, it follows that the proposed MSGT algorithm performs $\mathcal{O}\left(\frac{1}{L}\left(\frac{N}{\gamma K}\right)^{K}\right)$, times fewer computational operations compared to MAP.
III-C Discussion
To the best of our knowledge, MSGT is the first GT algorithm to effectively leverage Markovian prior statistics. Unlike numerous previous approaches, MSGT utilizes initial probabilities and transition matrices without necessitating specific adjustments for new use cases. The algorithm offers the flexibility to be fine-tuned to optimize its performance in accordance with the available computational resources and the number of tests, $T$. The simple reduction of the search space in Stage 1 enables MSGT to handle challenging regimes with a small number of tests. Stage 2, particularly the LVA step, contributes to its high success probability. Additionally, using the parallel implementation of LVA, rather than the iterative one, keeps the complexity low [32]. It is important to note that, as explained in [32], achieving optimal results is ensured only with a very large $L$, inevitably leading to complexity equivalent to MAP’s. However, as we empirically demonstrate in the following section, results equivalent to MAP’s can be achieved with reasonable complexity. Moreover, it is shown that MSGT addresses practical regimes, e.g., in COVID-19 detection [46] ($T=48$ for $(N,K)=(384,4)$), in communications [47] ($(N,K)=(10^{5},6)$), and in GT quantizer [8] ($(N,K)=(1024,16)$ [48]).
Another aspect of novelty in this work is the integration of Viterbi Algorithm into the GT problem. In the context of Markovian priors, one can think of the population’s sequence of items as a sequence of observations stemming from a hidden Markov process within a given Markov model over $N$ steps. In that case, the selection of a Viterbi decoder becomes natural, offering an optimal and efficient decoding solution. However, the most likely sequence of items does not necessarily include $K$ defective items. Particularly, in sparse signal scenarios, which is the focus of GT, the most likely sequence typically involves the minimum number of defectives that explain the observations. As a result, VA may not necessarily detect more defective items than already known and may detect even more than $K$. To address this, we employ LVA, which produces a list of the $L$ most likely sequences, such that choosing an appropriate value for $L$ guarantees a successful recovery.
Like many previous works, MSGT assumes precise knowledge of $K$. In practical use cases, this assumption relies on using an accurate estimator for $K$ employing $\mathcal{O}\left(\log_{2}(N)\right)$ tests [49], but the estimation might be erroneous. For the sake of practical completeness, it is worth noting that modifications can be made to handle incorrect estimates of $K$, albeit with increased computational complexity. The authors of [4] suggest altering the ML estimator to consider all possible sets, without restricting the number of defective items, and show that the probability of success is almost unaffected. MSGT relies on knowing $K$ only for the MAP estimation step. Thus, a similar adaptation can be applied. In MSGT, the MAP estimator should consider all possible sets from LVA and cannot stop with the first set that explains the outcome $Y$. Hence, such modification will increase the complexity compared to [4].
(a) Number of computational operations.
(b) Correlation matrix $\mathbf{\Phi}_{i}$ for $\tau=2$.
(c) Long memory priors - exact recovery.
Figure 6: (a) Number of computational operations in MSGT and MAP as given in Theorem 4 and Remark 1, respectively. (b) Example of correlation matrix $\mathbf{\Phi}_{i}$. (c) Probability of success of MSGT with exact prior statistics of 3-memory-steps Markov process, and with limited prior statistics assuming the Markov process has only one memory step. $N=1024,K=8$, 1000 iterations.
IV Numerical Evaluation
This section assesses the performance of the proposed MSGT algorithm by numerical study. First, in Subsection IV-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III. Then, in Subsection IV-2, we contrast the performance of MSGT with the one of DD, ML, and MAP in a practical regime of $N$ and $K$. To generate the correlated prior information between adjacent items, we use the Gilbert-Elliot (GE) model [50], which is a stationary Markov chain of order 1 with two states: one representing an error phase and the other an error-free phase. The GE model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$. These characteristics align well with the inputs required by the MSGT algorithm. In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [34, 35], or in sparse signal recovery in signal-processing with fixed vector size of input samples [9, 10, 11, 13, 51]), we show that the low computational complexity MSGT algorithm can reduce the number of pool tests by at least $25\%$.
IV-1 Theoretical Analysis
In Fig. 4(a) we show the concentration of $\left|\mathcal{P}^{(S_{1,1})}\right|$, as obtained from the simulation, along with the bound on its expectation and on the worst-case that were calculated in Theorems 1 and 2 respectively. Note that the worst-case scenario regarding MSGT is when LVA filters the correct set of defective items. That may happen if the number of possibly defective items exceeds the threshold $\gamma K$. Since in MSGT, we only allow this deviation from the average and ignore the case of exceeding this threshold, our upper bound for the worst-case does not cover all potential realizations of $\left|\mathcal{P}^{(S_{1,1})}\right|$. Similarly, Fig. 4(b) demonstrates the concentration of $\left|\mathcal{P}^{(DD)}\right|$, as acquired through simulation, and the lower bound on its expectation as given in Theorem 3. Fig. 4(c) illustrates the numerically computed lower bound for $\gamma$, derived from the inequality provided in Proposition 1. For this simulation, we fix $N=500$, $K=3$, and calculate $\gamma$ value for a specific range of $T$ relative to the upper bound of ML. As explained in Section III, our conjecture asserts that any value of $\gamma$ surpassing this lower bound guarantees that MSGT performance will be at least on par with that of ML. Therefore, whenever computational resources allow, it is advisable to choose the value of $\gamma$ corresponding to the lower bound. This approach was followed in the subsequent simulations, and the practical outcomes presented in Subsection IV-2 provide empirical support for our conjecture. Fig.6(a) compares the number of potential combinations to be examined in the MAP step, with and without the execution of the LVA step in MSGT. It can be observed that the LVA step performs an extensive filtering process, which allows MSGT to remain feasible even when executing MAP is no longer possible, especially in a regime below ML upper bound.
The converse of the GT problem with general prior statistics was developed by Li et al. [52] and according to which any GT algorithm with a maximum error probability $P_{e}$ requires a number of tests that satisfies: $T\leq\left(1-P_{e}\right)H(\mathbf{U})$, where $H(\cdot)$ denotes entropy. Using the joint entropy identity we have: $\left(1-P_{e}\right)H(\mathbf{U})=\sum_{i=1}^{N}H\left(U_{i}\lvert U_{1},\ldots,U_{i-1}\right)$.
The GE model considered in our numerical evaluations is a stationary Markov chain with $\tau=1$. Thus, $P\left(U_{i}\lvert U_{0},\ldots,U_{i-1}\right)=P\left(U_{i}\lvert U_{i-1}\right)$, for $i\in\{2,\ldots,N\}$. Substituting those priors, it follows that the converse of our problem is: $T\leq H(U_{1})+\sum_{i=2}^{N}H\left(U_{i}\lvert U_{i-1}\right)$. This bound is illustrated in the practical scenarios tested in Fig. 4 and Fig. 5.
IV-2 Algorithm Evaluation
We demonstrate the performance of MSGT using simulation. The population is sampled from GE model, and the regime is $K=\mathcal{O}\left(N^{\theta}\right)$ with $\theta\leq 1/3$. The GE parameters serve as our prior statistics, but in practice, we ignore samples where the number of defective items does not match $K$. In addition, although in the complexity analysis, we considered Bernoulli encoder for simplification, here we use a near-constant column weight encoder that optimizes DND’s performance [37], with $\frac{\ln{2}}{K}T$ tests sampled randomly for each item. The parameter $\gamma$ was chosen to satisfy (3) and $L=500$ was chosen empirically. Fig. 5 compares MSGT to MAP, ML, and DD algorithms. We run DND and DD before ML and MAP for reasonable runtime and memory consumption. The population includes $N\in\{500,1024,10000\}$ items and $K\in\{3,8,13\}$ defective items, respectively, and the empirical success probability is the average over $1000$ experiments. Note that for $N=1024,10000$ it is no longer possible to compare the performance since ML and MAP become infeasible (Fig. 5(b),5(c)).
Finally, MSGT was also tested with more complex probabilistic models. We sample the population based on a 3-memory-steps Markov process ($\mathbf{\Phi}_{i}$ is a $8\times 2$ matrix for all $i$). We execute MSGT using these prior statistics and also execute it with limited prior statistics, assuming that the process has only one memory step ($\mathbf{\Phi}_{i}$ is $2\times 2$). For example of $\mathbf{\Phi}_{i}$ representing a 2-memory-steps process, see Fig 6(b). The results are shown in Fig. 6(c). It is evident that utilizing the prior of long memory improves the success probability by $10\%$ in this scenario. We do note again that for the practical regime tested as in [46, 47, 8, 48], i.e., $N=1024$ and $K=8$, it is no longer possible to compare the performance since, unlike the efficient proposed MSGT algorithm, ML and MAP decoders for GT become infeasible.
V Discussion and Future Work
To the best of our knowledge, the existing solutions in literature do not offer an efficient and general solution for the regime below $T_{ML}$. In this work, we are focusing on this regime of the number of tests (i.e., $T<T_{ML}$), where the proposed MSGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of the potential defective combinations. This approach yields an efficient computational solution that, as demonstrated in our simulation result for practical scenarios, can approach the minimum number of tests as MAP algorithm.
Future work includes the derivation of an upper bound for this problem, which holds significant importance in comprehending the algorithm’s potential.

### v3

@@
-Non-Adaptive Multi-Stage Algorithm for Group Testing with Prior Statistics
+Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics
 Abstract
@@
 We introduce a variation of List Viterbi Algorithm (LVA) to enable accurate recovery using much fewer tests than objectives, which efficiently gains from the correlated prior statistics structure.
-Our numerical results demonstrate that the proposed Multi-Stage GT (MSGT) algorithm can obtain the optimal Maximum A Posteriori (MAP) performance with feasible complexity in practical regimes, such as with COVID-19 and sparse signal recovery applications, and reduce in the scenarios tested the number of pooled tests by at least $25\%$ compared to existing classical low complexity GT algorithms.
+We also provide a sufficiency bound to the number of pooled tests required by any Maximum A Posteriori (MAP) decoder with an arbitrary correlation between infected items.
+Our numerical results demonstrate that the proposed Multi-Stage GT (MSGT) algorithm can obtain the optimal MAP performance with feasible complexity in practical regimes, such as with COVID-19 and sparse signal recovery applications, and reduce in the scenarios tested the number of pooled tests by at least $25\%$ compared to existing classical low complexity GT algorithms.
 Moreover, we analytically characterize the complexity of the proposed MSGT algorithm that guarantees its efficiency.
@@
 We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the MSGT computational complexity efficiency.
+We also provide an analytical sufficiency bound to the number of pooled tests required by any maximum a-posteriori (MAP) decoder.
+This bound applies to any arbitrary underlying correlation between infected items.
 Our numerical results demonstrate that in practical regimes for COVID-19 [34, 35] and sparse signal recovery in signal processing [9, 10, 11, 13], the low computational complexity MSGT algorithm proposed herein can reduce in the scenarios tested the number of pool tests by at least $25\%$.
@@
 Section II formally describes the GT model with correlated prior statistics.
-Section III presents the MSGT algorithm and the analytical results.
-Section IV details the simulation evaluation.
-Finally, Section V provides concluding remarks and future directions.
+Section III presents the MSGT algorithm and the analytical results, and Section IV presents a sufficiency bound for MAP decoder.
+Section V details the simulation evaluation.
+Finally, Section VI provides concluding remarks and future directions.
 II Problem Formulation
@@
 Section III-B provides analytical results, followed by a discussion in Section  III-C.
-For a detailed explanation of all the algorithms used as integral components of Algorithm 1, see Appendix -A.
+For a detailed explanation of all the algorithms used as integral components of Algorithm 1, see Appendix A.
 III-A Pool-Testing Algorithm
@@
 $(c)$ follows since $e^{-x}\geq\left(1-x/n\right)^{n}$ for any integer $n>0$ and $x>0$.
-Subsequently, the probability of success satisfies is given by
+Subsequently, the probability of success satisfies
 $\mathbb{E}\left[P_{s}^{(DD)}\right]=1-\mathbb{E}\left[P_{e}^{(DD)}\right]\geq 1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}.$
@@
 This completed the theorem proof. ∎
-The proposition below provides a sufficient condition on $\gamma$, such that, MSGT achieves performance that we conjecture outperforms ML.
+The proposition below provides a sufficient condition on $\gamma$, such that, MSGT achieves performance that outperforms ML.
 This can be assumed since LVA used for this step is an optimal MAP estimator [32].
@@
 Let $\eta$ denote the average success probability of ML decoder.
-For GT with correlated prior information and $\gamma\geq 1$, MSGT can achieve an average success probability equal or greater than $\eta$, if
+For GT with correlated prior information and $\gamma\geq 1$, MSGT achieve an average success probability equal or greater than $\eta$, if
 $T\geq\log{\eta}+\log{\binom{N}{K}}-\log{\binom{N-K}{K(\gamma-1)}}.$ (3)
@@
 Theorem 4.
-Consider a group test for a population of size $N$ items out of which $K$ are defective and a Bernoulli testing matrix.
+Consider a group test for a population of $N$ items, of which $K$ are defective and a Bernoulli testing matrix.
 The computational complexity of the MSGT algorithm is bounded by $\mathcal{O}\left(L\gamma^{K}KN\log_{2}N\right)$ operations.
@@
 $N=1024,K=8$, 1000 iterations.
-IV Numerical Evaluation
+IV MAP Analytical Bound for GT with General Correlated Prior Statistics
+In this section, we provide a sufficiency bound for the number of required pooled tests for any MAP decoder with an arbitrary underlying correlation between infected items.
+This bound also applies to the LVA-based MSGT algorithm, which can approach MAP decoder performance for a sufficiently large list [32].
+We calculate this bound by extending the known result of maximum likelihood (ML) decoding for Group Testing (GT) problems [17], which bounds the probability of error of a ML decoder using a Gallager-type bound [44] for multiple access channel (MAC) channels [50].
+Recall that a noiseless GT setup can be considered as a noiseless MAC setting with $N$ “users”, each having a single codeword.
+While these type of bounds are well explored in the literature for MAC channels with correlated prior statistics [51, 52, 53, 54, 55, 56, 57], they do not expand naturally to GT, since GT can be thought as a special case where only “$K$” users transmit at any given point.
+This key difference necessitates dedicated analysis as follows in this section.
+In particular, to gain from the correlated prior statistics information, unlike ML decoders and bounds e.g., in [17, 14].
+We first introduce new notations and definitions in Subsection IV-A, then provide the probability of error to bound the MAP performance in Subsection IV-B, characterizing in Subsection IV-C the maximal number of tests required to guarantee reliability as given in Theorem 5.
+IV-A Definitions and notations
+Let $\left[n\right]=\left\{1,\ldots,n\right\}$.
+Let $H(\cdot)$ and $I(\cdot;\cdot)$ denote Shannon’s entropy and mutual information, respectively [58].
+All vectors are column vectors and the $\log$ operator denotes $\log_{2}$, unless stated otherwise.
+Let $\omega\in\left[\binom{N}{K}\right]$ denote the index of a specific set of $K$ items out of the $N$ items.
+Throughout the paper, $\omega^{*}$ denotes the index of the set of the $K$ actual infected items.
+Let $W$ denote the random variable which points to the index of the infected items, and has a probability distribution of $P_{W}\left(\cdot\right):\left[\binom{N}{K}\right]\rightarrow\left[0,1\right]$.
+In the following, $W$ might be omitted for convenience of read.
+The $\omega$-th set is denoted by $S_{\omega}$.
+Note that $\lvert S_{\omega}\rvert=K$.
+The rows of $\mathbf{X}$ that correspond to the items in $S_{\omega}$ are denoted by $\mathbf{X}_{S_{\omega}}$.
+The output For a given $\omega\neq\omega^{*}$, the true positives (TP)/ false negatives (FN)/ false positives (FP) are denoted by $TP\left(\omega,\omega^{*}\right),FN\left(\omega,\omega^{*}\right),FP\left(\omega,\omega^{*}\right)$.
+Note that $\lvert FN\left(\omega,\omega^{\prime}\right)\rvert=\lvert FP\left(\omega,\omega^{\prime}\right)\rvert$.
+In the following, we omit the dependency on $\omega,\omega^{*}$ and write $\mathbf{X}_{TP},\mathbf{X}_{FP},\mathbf{X}_{FN}$ instead of $\mathbf{X}_{TP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FN\left(\omega,\omega^{*}\right)}$ for notational simplicity.
+We say that an error of size $i$ happens if an output set of items $\omega\neq\omega^{*}$ is chosen and $\lvert FP\left(\omega,\omega^{*}\right)\rvert=i$.
+This event is denoted by $E_{i}$.
+Assuming the event $E_{i}$ happened, the index of the FN items is denoted by $j$, i.e.
+$j\in\left[\binom{N-K}{i}\right]$.
+Let $\Omega$ denote a function that converts a $j\in\left[\binom{N-K}{i}\right]$ index to a set of size $i$ of the FN items (the dependence on $\omega^{*}$ is omitted for notational simplicity).
+A decoder is denoted by $g$ and outputs some index $\omega=g\left(Y^{T}\right)$, where the dependency on $\mathbf{X}$ is omitted in favor of notational simplicity.
+These definitions are illustrated in the following example.
+Example
+Consider a case with $N=7$ items, where $K=3$ are defective.
+Suppose the first $3$ items are defective, i.e.
+items $\left\{1,2,3\right\}$ are defective.
+In this case, $S_{\omega^{\star}}=\left\{1,2,3\right\}$.
+Suppose there exists another set that is more likely: $S_{\omega^{\prime}}=\left\{1,4,5\right\}$.
+In this case, $TP\left(\omega^{\prime},\omega^{*}\right)=\left\{1\right\},FP\left(\omega^{\prime},\omega^{*}\right)=\left\{4,5\right\},\Omega_{2}\left(j\right)=FN\left(\omega^{\prime},\omega^{*}\right)=\left\{2,3\right\}$, where $j$ is the index of the FN set.
+In this example, an error of size $i=\lvert FP\left(\omega^{\prime},\omega^{*}\right)\rvert=2$ occurred, i.e.
+the event $E_{2}$ happened.
+IV-B Probability of Error
+The optimal MAP decoding rule that minimizes the decoding error probability selects an index $\omega$ that satisfies
+$P_{W}\left(S_{\omega}\right)P(Y^{T}|\mathbf{X}_{S_{\omega}})\geq P_{W}\left(S_{\omega^{\prime}}\right)P\left(Y^{T}|\mathbf{X}_{S_{\omega^{\prime}}}\right);\hskip 5.69046pt\forall\omega^{\prime}\neq\omega.$ (8)
+Recall that the pooling matrix $\mathbf{X}$ is chosen according to a Bernoulli distribution with parameter $p$ in an i.i.d.
+fashion.
+An error occurs if there exists $\omega^{\prime}\neq\omega^{*}$ such that $P_{W}\left(\omega^{\prime}\right)P\left(Y^{T}\lvert\mathbf{X}_{S_{\omega^{\prime}}}\right)\geq P_{W}\left(\omega^{*}\right)P\left(Y^{T}\lvert\mathbf{X}_{S_{\omega^{\star}}}\right)$.
+This event is denoted by $\zeta_{\omega^{\prime}}$.
+The average probability of error, in terms of exact recovery as defined in Section II, is defined as
+$P_{e}=\sum_{\omega}P_{W}\left(S_{\omega}\right)P\left(g\left(Y^{T}\right)\neq\omega\lvert\omega\right).$ (9)
+Let $E_{i}^{j}$ denote the event where an error of size $i$ at index $j\in\left[\binom{N}{i}\right]$ happens. The error probability for this event is
+$P\left(E_{i}^{j}\right)=\sum_{\omega\in\left[\binom{N}{K}\right]}P\left(w\right)P\left(E_{i}^{j}\lvert w\right).$ (10)
+By plugging in (10) into (9), we get
+$P_{e}$ $=\sum_{i=1}^{K}\sum_{j=1}^{\binom{N-K}{i}}P\left(E_{i}^{j}\right)$
+$=\sum_{i=1}^{K}\sum_{j=1}^{\binom{N-K}{i}}\sum_{\omega=1}^{\binom{N}{K}}P_{W}\left(\omega\right)P\left(E_{i}^{j}\lvert W=\omega\right).$
+IV-C Sufficiency GT MAP Bound
+Let $(S_{2},S_{1})$ denote a partition of the defective set $S$ into disjoint sets $S_{2}$ and $S_{1}$, with cardinalities $i$ and $K-i$, respectively.
+This partition allows for categorizing error occurrences into distinct groups, where in group $i$, $K-i$ defective items are already identified correctly, and the primary error scenario arises from trying to detect the remaining $i$ defective items.
+Theorem 5.
+Consider a group test with a MAP decoder for the population with correlated prior statistics. For $K=O(1)$, if the number of the tests satisfies
+$T\geq(1+\varepsilon)\max_{i=1,\ldots,K}\frac{K}{i}H\left(P_{S_{2}|S_{1}}\right)$
+then, as $N\rightarrow\infty$ the average error probability approaches zero.
+The sufficiency proof of Theorem 5 is twofold.
+We first bound in the following lemma the error of probability of the optimal MAP decoder as defined in (8).
+We then use this error of probability analysis to bound the required number of tests required by any decoder.
+Lemma 1.
+Consider a group test with a MAP decoder for a population with correlated prior statistics. The error probability $P(E_{i}^{j})$ in (10) is bounded by
+$P(E_{i}^{j})\leq 2^{-T\left(E_{0}(\rho)-\frac{\frac{1}{\binom{K}{i}}E_{s,j}(\rho,P_{W})}{T}\right)},$
+where the error exponent $E_{0}(\rho)$ is given by
+$\begin{split}&E_{0}(\rho)=\\ &-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$
+the correlated prior statistics function of the items $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$ is given by
+$\begin{split}E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)=\log\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}},\end{split}$
+for $0\leq\rho\leq 1$.
+Proof of Lemma 1: The proof is given in Appendix B. ∎
+Proof of Theorm 5.
+Define
+$\emph{f}(\rho)=E_{o}(\rho)-\frac{E_{s,j}(\rho,P_{W})}{T}.$
+Since $0\leq\rho\leq 1$ can be optimized, we aim to show that $Tf(\rho)\to\infty$ as $N\to\infty$ for some $\rho$ within this range.
+If this holds for all $E_{i}^{j}$, then, due to the resulting exponential decay of $P(E_{i}^{j})$, applying a simple union bound will demonstrate that the total error probability remains small, thus completing the proof of Theorem 5.
+Since the function $f(\rho)$ is differentiable and admits a power series expansion, we use a Taylor series expansion in the neighborhood of $\rho=0$, yielding
+$f(\rho)=f(0)+\rho f^{\prime}(0)+\frac{\rho^{2}}{2}f^{\prime\prime}(\psi)$
+for some $\psi\in(0,\rho)$.
+For completeness, we show as in [17] that
+$\frac{\partial E_{o}}{\partial\rho}|_{\rho=0}$
+$=$ $\sum_{Y}\sum_{X_{TP}}[\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\log P\left(Y,X_{TP}|X_{FP}\right)$
+-- $\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})]$
+$=$ $\sum_{Y}\sum_{X_{TP}}\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})$
+$\log\frac{p(Y,X_{TP}|X_{FP})}{\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})}=I(X_{FP};X_{TP},Y).$
+And also, in Appedix C, that
+$\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}=H(P_{S_{2}|S_{1}}).$
+This is obtained directly from [59] as also elaborated in [57, Chap. 1]. Note that $E_{o}(0)=0$ and $E_{s,j}(0,P_{W})=0$, hence we have
+$T\emph{f}(\rho)=$ $T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)-\frac{H\left(P_{S_{2}|S_{1}}\right)}{T}\Bigg{)}$
+-- $\log\frac{1}{\binom{K}{i}}+T\frac{\rho^{2}}{2}E_{o}^{\prime\prime}\left(\psi\right)-T\frac{\rho^{2}}{2}E_{s,j}^{\prime\prime}\left(\psi\right),$
+and, if
+$T\geq\left(1+\varepsilon\right)\max_{i=1,\ldots,K}\frac{H\left(P_{S_{2}|S_{1}}\right)}{I\left(X_{FP};X_{TP},Y\right)}$ (11)
+for some $\varepsilon>0$, we have
+$T\emph{f}(\rho)\geq-\log\frac{1}{\binom{K}{i}}$
+$+T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)+\frac{\rho}{2}\left(E_{o}^{\prime\prime}(\psi)-E_{s,j}^{\prime\prime}\left(\psi\right)\right)\Bigg{)}$
+Moreovver, note that $E_{o}^{\prime\prime}\left(\psi\right)$ is negative [44]. However, it is independent of the other constants and $T$, hence choosing
+$0<\rho<\frac{2I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)}{|E_{o}^{\prime\prime}\left(\psi\right)-E_{s,j}^{\prime\prime}\left(\psi\right)|},$
+and as $\frac{1}{\binom{K}{i}}$ is fixed and independent of $T$, we have $T\emph{f}(\rho)\rightarrow\infty$ as $N\rightarrow\infty$.
+In practice, the expression $I(X_{FP};X_{TP},Y)$ in (11) is important to understand how many tests are required, yet it is not a function of the problem parameters and is bounded in [17] and [45] to get a better handle on $T$.
+Claim 1 ([17] and [45]).
+For large $K$, and under a fixed input distribution for the testing matrix $\left(\frac{\ln(2)}{K},1-\frac{\ln(2)}{K}\right)$, the mutual information between $X_{FP}$ and $(X_{TP},Y)$ is bounded by $I(X_{FP};X_{TP},Y)\geq\frac{i}{K}$.
+Hence, by applying Clain 1 in Eq. (11), we achieves the sufficiency bound on $T$ as provided in Theorem 5,
+$(1+\varepsilon)\max_{i=1,\ldots,K}\frac{K}{i}H\left(P_{S_{2}|S_{1}}\right),$
+such that, as $N\rightarrow\infty$ the average error probability approaches zero, and reliability is established.
+V Numerical Evaluation
 This section assesses the performance of the proposed MSGT algorithm by numerical study.
-First, in Subsection IV-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III.
-Then, in Subsection IV-2, we contrast the performance of MSGT with the one of DD, ML, and MAP in a practical regime of $N$ and $K$.
-To generate the correlated prior information between adjacent items, we use the Gilbert-Elliot (GE) model [50], which is a stationary Markov chain of order 1 with two states: one representing an error phase and the other an error-free phase.
+First, in Subsection V-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III.
+Then, in Subsection V-2, we contrast the performance of MSGT with the one of DD, ML, and MAP in a practical regime of $N$ and $K$.
+To generate the correlated prior information between adjacent items, we use the Gilbert-Elliot (GE) model [60] (also considered very recently analytically in [61]), which is a stationary Markov chain of order 1 with two states: one representing an error phase and the other an error-free phase.
 The GE model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$.
@@
 In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [34, 35], or in sparse signal recovery in signal-processing with fixed vector size of input samples [9, 10, 11, 13, 62]), we show that the low computational complexity MSGT algorithm can reduce the number of pool tests by at least $25\%$.
-IV-1 Theoretical Analysis
+V-1 Theoretical Analysis
 In Fig.
@@
 Therefore, whenever computational resources allow, it is advisable to choose the value of $\gamma$ corresponding to the lower bound.
-This approach was followed in the subsequent simulations, and the practical outcomes presented in Subsection IV-2 provide empirical support for our conjecture.
+This approach was followed in the subsequent simulations, and the practical outcomes presented in Subsection V-2 provide empirical support for our conjecture.
 Fig.6(a) compares the number of potential combinations to be examined in the MAP step, with and without the execution of the LVA step in MSGT.
@@
 The converse of the GT problem with general prior statistics was developed by Li et al.
-[52] and according to which any GT algorithm with a maximum error probability $P_{e}$ requires a number of tests that satisfies: $T\leq\left(1-P_{e}\right)H(\mathbf{U})$, where $H(\cdot)$ denotes entropy.
-Using the joint entropy identity we have: $\left(1-P_{e}\right)H(\mathbf{U})=\sum_{i=1}^{N}H\left(U_{i}\lvert U_{1},\ldots,U_{i-1}\right)$.
+[63] and according to which any GT algorithm with a maximum error probability $P_{e}$ requires a number of tests that satisfies
+$T\geq\left(1-P_{e}\right)H(\mathbf{U}).$
+Using the joint entropy identity we have
+$\left(1-P_{e}\right)H(\mathbf{U})=\sum_{i=1}^{N}H\left(U_{i}\lvert U_{1},\ldots,U_{i-1}\right).$
 The GE model considered in our numerical evaluations is a stationary Markov chain with $\tau=1$.
 Thus, $P\left(U_{i}\lvert U_{0},\ldots,U_{i-1}\right)=P\left(U_{i}\lvert U_{i-1}\right)$, for $i\in\{2,\ldots,N\}$.
-Substituting those priors, it follows that the converse of our problem is: $T\leq H(U_{1})+\sum_{i=2}^{N}H\left(U_{i}\lvert U_{i-1}\right)$.
-This bound is illustrated in the practical scenarios tested in Fig.
-4 and Fig.
-5.
-IV-2 Algorithm Evaluation
+Substituting those priors, it follows that the converse of our problem is
+$T\geq H(U_{1})+\sum_{i=2}^{N}H\left(U_{i}\lvert U_{i-1}\right).$
+This bound is illustrated in the practical scenarios tested in Fig. 4 and Fig. 5.
+V-2 Algorithm Evaluation
 We demonstrate the performance of MSGT using simulation.
@@
 We do note again that for the practical regime tested as in [46, 47, 8, 48], i.e., $N=1024$ and $K=8$, it is no longer possible to compare the performance since, unlike the efficient proposed MSGT algorithm, ML and MAP decoders for GT become infeasible.
-V Discussion and Future Work
-To the best of our knowledge, the existing solutions in literature do not offer an efficient and general solution for the regime below $T_{ML}$.
-In this work, we are focusing on this regime of the number of tests (i.e., $T<T_{ML}$), where the proposed MSGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of the potential defective combinations.
+VI Discussion and Future Work
+To the best of our knowledge, the existing solutions in the literature do not offer an efficient and general solution for the regime below $T_{ML}$.
+In this work, we focus on this regime of the number of tests (i.e., $T<T_{ML}$), where the proposed MSGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of the potential defective combinations.
 This approach yields an efficient computational solution that, as demonstrated in our simulation result for practical scenarios, can approach the minimum number of tests as MAP algorithm.
-Future work includes the derivation of an upper bound for this problem, which holds significant importance in comprehending the algorithm’s potential.
+VII Acknowledgment
+The authors would like to thank Arezou Rezazadeh for fruitful discussions and suggestions.
+Appendix A Recovery Algorithms
+For the completeness of our proposed Algorithm 1, we present here all the algorithms that we use as integral components within MSGT, and provide detailed explanations for each.
+A-A Definitely Not Defective (DND)
+DND algorithm [39, 40, 41, 19], which is shown in Algorithm 2 takes the testing matrix $\mathbf{X}$ and the results vector $Y$ as inputs.
+It systematically compares each column of the $\mathbf{X}$ with $Y^{T}$.
+If a specific column contains a value of $1$ while its corresponding test result is 00, we say that this column cannot explain the test result.
+Since each column represents one item in the population, if $\mathbf{X}(i)$ cannot explain $Y^{T}$, then DND marks this item as definitely non-defective.
+The output is the set of all such items, denoted by $\mathcal{P}^{(DND)}$.
+Algorithm 2 Definite Not Defectives (DND) [39, 40, 41, 19]
+Input: $\mathbf{X},Y$
+Output: $\mathcal{P}^{(DND)}$
+1:$\mathcal{P}^{(DND)}\leftarrow\emptyset$
+2:for $i$ s.t. $\mathbf{Y}_{i}=0$ do $\mathcal{P}^{(DND)}\leftarrow\mathcal{P}^{(DND)}\cup\{j\lvert X_{i,j}=1\}$
+3:end for
+4:return $\mathcal{P}^{(DND)}$
+A-B Definite Defectives (DD)
+DD algorithm [19], as shown in Algorithm 3 takes as inputs $\mathcal{P}^{(S_{1,1})}$, the testing matrix $\mathbf{X}$ and the results vector $Y$.
+DD’s objective is to identify test results that can be explained by a single possibly defective item.
+In practice, DD examines the positive entries of $Y$.
+For each positive test, if there is only one possibly defective item participating in it, DD marks that item as definitely defective.
+We skip the last step of DD in [19], and in our implementation, DD returns only the set of definitely defective items, denoted by $\mathcal{P}^{(DD)}$.
+Algorithm 3 Definite Defective (DD) [19]
+Input: $\mathbf{X},Y,\mathcal{P}^{(S_{1,1})}$
+Output: $\mathcal{P}^{(DD)}$
+1:$\mathcal{P}^{(DD)}\leftarrow\emptyset$
+2:for $i$ s.t. $\mathbf{Y}_{i}=1$ do
+3:    for $j\in\mathcal{P}^{(S_{1,1})}$ do
+4:         if $\mathbf{X}_{i,j}=1$ and $\mathbf{X}_{i,j^{\prime}}=0,$ $\forall j^{\prime}\neq j\in\mathcal{P}^{(S_{1,1})}$ then
+5:             $\mathcal{P}^{(DD)}\leftarrow\mathcal{P}^{(DD)}\cup\left\{j\right\}$
+6:         end if
+7:    end for
+8:end for
+9:return $\mathcal{P}^{(DD)}$
+A-C List Viterbi Algorithm (LVA)
+Our variations of the LVA algorithm [32], outlined in Algorithm 4, is designed to return the $L$ most likely sequences, representing the estimation of the whole population, for a given $L\geq 1$.
+In the algorithm we suggest herein, the key differences are: (1) the population sequence in GT replaces the time sequence in classical LVA as given in [32], and (2) we use the aggregated sequence likelihood instead of the general cost function presented in the original paper.
+In particular, in the suggested algorithm, as we traverse the trellis diagram, we are iteratively maximizing the likelihood of the sequence representing the status of the population, tested with correlated and non-uniform prior statistics, denoted by $\Psi$ (see line 15).
+Algorithm 4 List Viterbi Algorithm [32]
+Input: $L,\tau,\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N},\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$
+Output: ${\mathbf{Z}}$
+1:Initialization:
+2:for $s\leftarrow 1$ to $2^{\tau}$ do
+3:    $\Psi(1,s,1)\leftarrow\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}\left(s\right)$triangleright initial probability for each
+                                     state
+4:    $\xi\left(n,s,l\right)\leftarrow s$triangleright previous state of each state
+5:    for $l\leftarrow 2$ to $L$ do
+6:         $\Psi(1,s,l)\leftarrow 0$
+7:         $\xi\left(n,s,l\right)\leftarrow s$
+8:    end for
+9:end for
+10:Recursion:
+11:for $n\leftarrow 2$ to $N$ do
+12:    for $s_{2}\leftarrow 1$ to $2^{\tau}$ do
+13:         for $l\leftarrow 1$ to $L$ do
+14:             $\triangleright$ find the probability of the $l$-most likely
+                   previous state and its rank
+15:             $\Psi(n,s_{2},l)=\max^{(l)}_{{\begin{subarray}{c}s_{1}\in\{1,...,2^{\tau}\}\\ k\in\{1,..,L\}\end{subarray}}}\newline \hskip 103.50017pt\left\{\Psi(n-1,s_{1},k)\Phi_{n}\left(s_{1},s_{2}\right)\right\}$
+16:             $\left({s_{1}}^{*},k^{*}\right)=\arg\max^{(l)}_{{\begin{subarray}{c}s_{1}\in\{1,...,2^{\tau}\}\\ k\in\{1,..,L\}\end{subarray}}}\newline \hskip 103.50017pt\left\{\Psi(n-1,s_{1},k)\Phi_{n}\left(s_{1},s_{2}\right)\right\}$
+17:             $\xi\left(n,s_{2},l\right)\leftarrow{s_{1}}^{*}$
+18:             $\chi\left(n,l\right)\leftarrow k^{*}$
+19:         end for
+20:    end for
+21:end for
+22:Backtracking:
+23:for $l\leftarrow 1$ to $L$ do
+24:    $\mathbf{Z}_{l,N}\leftarrow\arg\max_{s\in\{1,...,2^{\tau}\}}\left\{\Psi(N,s,l)\right\}$
+25:    $l_{N}\leftarrow\chi\left(n,l\right)$
+26:    for $n\leftarrow N-1$ to $1$ do
+27:         $\mathbf{Z}_{l,n}\leftarrow\xi\left(n+1,\mathbf{Z}_{l_{n+1},n+1},l_{n+1}\right)$
+28:         $l_{n}\leftarrow\chi\left(n+1,l_{n+1}\right)$
+29:    end for
+30:end for
+31:return $\mathbf{Z}$
+This algorithm operates in three primary steps: first initialization of the setup, then recursion using a trellis structure to compute probabilities for all possible sequences while eliminating unlikely ones in each step, and finally, backtracking to reconstruct the most probable sequences.
+The algorithm inputs are the number of the most likely list $L$, the number of memory steps to consider, $\tau$, and the prior statistics $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N},\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$.
+Let $\Psi\in[0,1]^{N\times 2^{\tau}\times L}$ denote the probabilities of the $L$ most likely states along the trellis.
+Let $\xi\in\{1,...,2^{\tau}\}^{N\times 2^{\tau}\times L}$ denote the corresponding previous state of each state along the trellis, and let $\chi\in\{1,...,L\}^{N\times 2^{\tau}\times L}$ denote the corresponding rank of the current state, among the $L$ most likely options.
+In the initialization stage, we fill the given initial probabilities $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}$ in the corresponding entries of $\Psi$, and we set each state to be its previous state (lines 3-4).
+In the recursion stage, we iterate from the second item to the last, and for each item, compute all the $2^{\tau}\times 2^{\tau}$ possible transition probabilities between all the possible states he could be in and the states of its predecessor.
+For each item, state, and rank, we set in $\Psi$ the $L$-most likely probabilities of the overall sequences from the first item until the current item and state (line 15), and we set in $\xi$ the corresponding previous state and in $\chi$ the corresponding rank (lines 17-18).
+Here, $\max^{(l)}$ denotes the $l$-largest value.
+In the backtracking stage, we identify the $L$-most likely sequences based on the probabilities in the entries of the $\Psi$ that correspond to the last item (line 24), and then backtrack the steps of these sequences using the information in $\xi$ (line 27).
+The algorithm returns a list of these sequences, denoted by $\mathbf{Z}\in\left\{1,...,2^{\tau}\right\}^{L\times N}$.
+If $\tau>1$, and a further processing is applied to map the states $\{1,...,2^{\tau}\}$ to the binary states $\{0,1\}$, representing “defective” and “non-defective”.
+A-D Maximum A Posteriori (MAP)
+The MAP estimator [44], as shown in Algorithm 5, returns the set $c\in\mathbf{C}$ with the highest maximum a posteriori probability among all the sets, that is, the set $c$ that obtains the maximum $P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)$.
+In this expression, the first probability represents the likelihood of obtaining the results $Y$ from a group test using the testing matrix $\mathbf{X}$ and the given set $c$ as the defective set.
+If the set $c$ and the given testing matrix $\mathbf{X}$ cannot explain for $Y$, then this probability is equal to zero.
+The second probability corresponds to the prior probability of the overall defective set $c$, calculated using $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N},\left\{\mathbf{\Phi}_{i}\right\}_{i=1}^{N}$.
+The MAP algorithm returns the estimated defective set, denoted by $\hat{\mathcal{K}}$.
+Algorithm 5 Maximum A Posteriori (MAP) [44]
+Input: $\mathbf{X},Y,\mathbf{C},\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N},\left\{\mathbf{\Phi}_{i}\right\}_{i=1}^{N}$
+Output: $\hat{\mathcal{K}}$
+1:return $\arg\max_{c\in\mathbf{C}}{P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)}$
+Appendix B Probability of Error (Proof of Lemma 1)
+Let $i\in\left[K\right],j\in\left[\binom{N-K}{i}\right]$. Let ~$\tilde{\mathcal{I}}^{j}_{dif}$ denote the set of all indices $\omega$ that result in a the set of FN of size $i$ as in $j$, i.e.
+~$\tilde{\mathcal{I}}^{j}_{dif}=\left\{\omega\in\left[\binom{N}{K}\right]:\quad FN\left(\omega,\omega^{*}\right)=\Omega_{i}\left(j\right)\right\}.$ (12)
+Recall $\omega^{*}$ denotes the index of the set of the $K$ actual infected items.
+To visualize this definition, consider the example in Section IV-A.
+For $S_{\omega^{\star}}=\left\{1,2,3\right\}$, let $j$ be the index such that $\Omega_{2}\left(j\right)=\left\{2,3\right\}$.
+In this case, ~$\tilde{\mathcal{I}}^{j}_{dif}$ has all indices that point to $\left\{1,4,5\right\},\left\{1,4,6\right\},\left\{1,4,7\right\},\left\{1,5,6\right\},\left\{1,5,7\right\},\left\{1,6,7\right\}$.
+Let $\zeta_{\omega^{\prime}},\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}$ denote the event that $S_{\omega^{\prime}}$ is more likely than $S_{\omega^{\star}}$ (see Eq.
+(8)).
+We use the Union Bound to bound $P\left(E_{i}^{j}\lvert W=\omega^{*}\right)$, such that
+$P\left(E_{i}^{j}\lvert W=\omega^{*}\right)$ $\leq P\left(\bigcup_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}\zeta_{\omega^{\prime}}\right)\leq\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P\left(\zeta_{\omega^{\prime}}\right)$
+For MAP decoder the probability of $\zeta_{\omega^{\prime}}$ is given by
+$P\left(\zeta_{\omega^{\prime}}\right)=\sum_{\mathbf{X}_{S_{\omega^{\prime}}}:P\left(\mathbf{X}_{S_{\omega^{\prime}}}\lvert Y^{T}\right)\geq P\left(\mathbf{X}_{S_{\omega^{\star}}}\lvert Y^{T}\right)}P\left(\mathbf{X}_{S_{\omega^{\prime}}}\lvert\mathbf{X}_{S_{\omega^{\star}}}\right)$
+$\overset{(a)}{\leq}\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\frac{P\left(Y^{T}\lvert\mathbf{X}_{S_{\omega^{\prime}}}\right)^{s}P_{W}\left(S_{\omega^{\prime}}\right)^{s}}{P\left(Y^{T}\lvert\mathbf{X}_{S_{\omega^{\star}}}\right)^{s}P_{W}\left(S_{\omega^{\star}}\right)^{s}}$
+$=\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\cdot$
+$\hskip 14.22636pt\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
+$\overset{(b)}{\leq}\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}},\hskip 5.69046pt\forall s>0.$
+$(a)$ follows that given $\mathbf{X}_{S_{\omega^{\star}}}$, $P(\mathbf{X}_{TP})$ is known and then only then probability of the uncommon codewords left: $P\left(\mathbf{X}_{S_{\omega^{\prime}}}\lvert\mathbf{X}_{S_{\omega^{\star}}}\right)=P\left(\mathbf{X}_{FP}\right)$ and from the fact that $d\triangleq\frac{P\left(Y^{T}\lvert\mathbf{X}_{S_{\omega^{\prime}}}\right)}{P\left(Y^{T}\lvert\mathbf{X}_{S_{\omega^{\star}}}\right)}\geq 1$, so we upperbound the whole expression by multiplying each term in the sum in $d^{s}\geq 1$ and sum over more test matrices.
+$(b)$ is due to the fact that $\mathbf{X}_{TP}$ and $\mathbf{X}_{FN}$ are independent, because they contain two distinct sets of codewords, as are $\mathbf{X}_{TP}$ and $\mathbf{X}_{FP}$.
+Therefore,
+$P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)=P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)=P\left(\mathbf{X}_{TP}\right).$
+By substituting Eq. (14) in Eq. (B), it follows that
+$P\left(E_{i}^{j}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$ (15)
+$\leq\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
+$=\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}}{P_{W}\left(\omega^{*}\right)^{s}}\cdot$
+$\hskip 71.13188pt\frac{\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}$
+$\overset{(a)}{\leq}\Biggl{[}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)}{P_{W}\left(\omega^{*}\right)^{s}}\cdot$
+$\hskip 71.13188pt\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}\Biggr{]}^{\rho}$
+$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot$
+$\hskip 71.13188pt\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}},$
+$\qquad\forall s>0,\rho\in\left[0,1\right]$
+where $(a)$ follows from the fact that the probability $P\left(E_{i}\lvert w^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)\leq 1$, hence it holds that if it is upperbounded by $U$, then $P\left(E_{i}\lvert w^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)\leq U^{\rho}$ for any $\rho\in[0,1]$.
+Using the law of total probability, we have
+$P\left(E_{i}^{j}\lvert W=\omega^{*}\right)$ (16)
+$=\sum_{\mathbf{X}_{S_{\omega^{\star}}}}\sum_{Y^{T}}P\left(\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
+$=\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(\mathbf{X}_{TP},Y^{T}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot$
+$\hskip 22.76228ptP\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
+$\overset{(a)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot$
+$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)\cdot$
+$\hskip 39.83368ptP\left(\mathbf{X}_{FN}\right)\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}}$
+$\overset{(b)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right))}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot$
+$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot$
+$\hskip 45.52458pt\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right))^{\frac{\rho}{1+\rho}}}$
+$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot$
+$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot$
+$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot$
+$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},$
+in $(a)$ we substitute Eq.
+(15), and $(b)$ follows as $\mathbf{X}_{FN}$ and $\mathbf{X}_{FP}$ are two independent matrices of $i$ codewords with the same size and distribution.
+Thus, $P\left(\mathbf{X}_{FN}\right)=P\left(\mathbf{X}_{FP}\right)$ and $P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)=P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)$, and the inequality holds for any $s>0$.
+In particular, for $s=1/\left(1+\rho\right)$.
+We first provide the following lemma to then continue with the error probability analysis using Eq. (16).
+Lemma 2.
+Let $i\in\left[K\right],j\in\left[\binom{N}{i}\right],\alpha\geq 0$. Then
+$\begin{split}\sum_{\omega\in\left[\binom{N}{K}\right]}P_{W}\left(\omega\right)^{\alpha}&=\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega\right)^{\alpha}\end{split}$ (17)
+where
+$\begin{split}&\mathcal{T}\mathcal{P}_{i}=\left\{S_{1}:\lvert S_{1}\rvert=K-i\right\},\\ &\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)=\left\{S_{2}:\lvert S_{2}\rvert=i,\;S_{2}\cap S_{1}=\emptyset\right\},\\ &\text{and}\\ &\omega\in\left[\binom{N}{K}\right]\;s.t.\;S_{\omega}=S_{1}\cup S_{2}.\end{split}$
+Proof.
+The left hand side (LHS) of Eq.
+(17) in Lemma 2 has $\binom{N}{K}$ elements.
+We count the number of elements of $\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}$ in the right hand side (RHS).
+The set $\mathcal{T}\mathcal{P}_{i}$ contains $\lvert\mathcal{T}\mathcal{P}_{i}\rvert=\binom{N}{K-i}$ items.
+For a given $S_{1}\in\mathcal{T}\mathcal{P}_{i}$, the set $\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)$ contains $\lvert\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)\rvert=\binom{N-\left(K-i\right)}{i}$ items.
+The overall number of combinations is given by
+$\binom{N}{K-i}\cdot\binom{N-\left(K-i\right)}{i}$
+$=$ $\frac{N!}{\left(K-i\right)!\left(N-\left(K-i\right)\right)!}\cdot\frac{\left(N-\left(K-i\right)\right)!}{i!\left(N-K\right)!}$
+$=$ $\frac{N!}{K!\left(N-K\right)!}\cdot\frac{K!}{\left(K-i\right)!i!}$
+$=$ $\binom{N}{K}\cdot\binom{K}{i}$
+By symmetry, every index $\omega$ in the RHS of (17) in Lemma 2 appears the same number of times, so an equality is achieved by dividing by the RHS by $\binom{K}{i}$. ∎
+One way to interpret the sets $\mathcal{T}\mathcal{P}_{i}$ and $\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)$ can be thought of as sets that iterate over TPs and FPs and the FN for a fixed set of FN items.
+Following the example in Section IV-A, the set number of errors is $i=2$.
+In this case, $\mathcal{T}\mathcal{P}_{i}=\left\{\left\{1\right\},\left\{4\right\},\left\{5\right\},\left\{6\right\},\left\{7\right\}\right\}$.
+For a given item $S_{1}\in\mathcal{T}\mathcal{P}_{i}$, $S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)$ runs over the remaining $i$ items that do not overlap with $S_{1}$.
+For instance, for $S_{1}=\left\{1\right\}$, we get $\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)=\left\{\left\{2,3\right\},\ldots,\left\{6,7\right\}\right\}$.
+Note that a specific set $S_{\omega}=S_{1}\cup S_{2}$ can be chosen $\binom{K}{i}=3$ different times.
+For instance, $S_{\omega}=\left\{1,2,3\right\}$ can be chosen $3$ times.
+Each $S_{1}\in\left\{\left\{1\right\},\left\{2\right\},\left\{3\right\}\right\}$ yields an $S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)$ that results in $S_{1}\cup S_{2}=\left\{1,2,3\right\}$.
+Thus, as demonstrated in the proof of Lemma 2, equality in (17) is achieved by dividing the RHS by $\binom{K}{i}=3$.
+Now, by substituting Eq. (16) in Eq. (10), it follows that
+$P\left(E_{i}^{j}\right)=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)P\left(E_{i}^{j}|W=\omega^{*}\right)$
+$\leq\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}$
+$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{1-\frac{\rho}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}$
+$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}$
+$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+=$\overset{(a)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}\right]$
+$\hskip 22.76228pt{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}$
+$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$\overset{(b)}{\leq}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}$
+$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+=$\overset{(c)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}$
+$\hskip 19.91684pt\left[\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}\right]^{T}$
+=$\overset{(d)}{=}2^{-T\left(E_{0}(\rho)-\frac{\frac{1}{\binom{K}{i}}E_{s,j}(\rho,P_{W})}{T}\right)},$
+where (a) follows from Lemma 2 with $\alpha=\frac{1}{1+\rho}$, (b) follows from the fact that for a fixed $S_{1}\in\mathcal{T}\mathcal{P}_{i}$, ~$\tilde{\mathcal{I}}^{j}_{dif}\subset\left\{\omega=\left(S_{1},S_{2}\right):S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)\right\}$, (c) follows since $X_{FP},X_{TP}$ represent a single column of $\mathbf{X}_{FP},\mathbf{X}_{TP}$ respectively, as the elements of $\mathbf{X}$ are drawn in i.i.d.
+fashion, and (d) follows for
+$\begin{split}&E_{0}(\rho)=\\ &-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$
+and
+$\begin{split}E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)=\log\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}.\end{split}$
+This completed the error probability proof.
+Appendix C Derivative of $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$
+In this section, we show the derivative of $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$ for $\rho=0$.
+This result is obtained directly from [59] as also elaborated in [57, Chap.
+1].
+Let $\tilde{f}\left(\rho,S_{1}\right)\triangleq\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}=\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}$, $A=\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}=\exp{\log\left({\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}}\right)}$.
+Thus,
+$\frac{\partial}{\partial\rho}\left(A\right)=A\frac{\partial}{\partial\rho}\left(\log\left(A\right)\right)=A\frac{\partial}{\partial\rho}\left(\left(1+\rho\right)\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)\right)$
+$=A\left(\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)+(1+\rho)\frac{\partial\tilde{f}\left(\rho,S_{1}\right)}{\partial\rho}\right)$
+=$\overset{(a)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\left(\log(\tilde{f}\left(\rho,S_{1}\right))+(1+\rho)\frac{\partial\tilde{f}\left(\rho,S_{1}\right)}{\partial\rho}\right)$
+=$\overset{(b)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))$
+$\hskip 17.07182pt-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
+=$\overset{(c)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))$
+$\hskip 17.07182pt-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
+=$\overset{(d)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}$
+$\hskip 17.07182ptP_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
+where (a) follows replacing $A$, (b) by derivative of $\tilde{f}\left(\rho,S_{1}\right)$, and (c) replacing $\tilde{f}\left(\rho,S_{1}\right)$.
+(d) follows by applying that $\log a-\frac{1}{\left(1+\rho\right)}\log b=\log\frac{a}{b^{\frac{1}{\left(1+\rho\right)}}}$.
+Now, we recall that $E_{s,j}(\rho,P_{W})=\log\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\right)$, such that $E_{s,j}(\rho,P_{W})=\log\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A$.
+Hence.
+from Eq.(C), we have
+$\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}=\frac{\partial\log\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A}{\partial\rho}=\frac{\partial\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A}{\partial\rho}$
+$=$ $\frac{1}{\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}$
+$\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}$
+$\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
+Next, setting $\rho=0$ in Eq. (C), and since
+$\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A|_{\rho=0}=\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)=1,$
+we have
+$\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}$
+$=$ $\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)$
+$\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)}{P_{W}\left(S_{1},S_{2}\right)}\Biggr{)}$
+$=$ $\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)\log\frac{P\left(S_{1}\right)}{P_{W}\left(S_{1},S_{2}\right)}\right)$
+$=$ $H\left(P_{S_{2}|S_{1}}\right).$

### v4

@@
 Thus, one can consider GT as a Boolean CS [17, 18].
-Traditional GT and its performance (i.e., the tradeoff between the number of pool tests and recovery algorithm complexity), focused on the probabilistic model in which the items are identically distributed and independent [19, 14].
+Traditional GT and its performance (i.e., the tradeoff between the number of pool tests and recovery algorithm complexity), focuses on the probabilistic model in which the items are identically distributed and independent [19, 14].
 Recent research explores cases where prior information about the correlation of objects is available [20, 2, 3, 21, 22].
@@
 To the best of our knowledge, no existing solution addresses the GT problem with general Markovian priors and applicable to a wide range of diverse applications.
-In this work, we introduce a practical non-adaptive Multi-Stage GT (MSGT) algorithm for correlated items with prior statistics.
+In this work, we introduce a practical non-adaptive Multi-Stage Group Testing (MSGT) algorithm for correlated items with prior statistics.
 The proposed multi-stage algorithm employs a new variation of the parallel List Viterbi Algorithm (LVA) [30, 31, 32] we designed for GT to enable accurate low complexity recovery using fewer tests.
@@
 We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the MSGT computational complexity efficiency.
-We also provide an analytical sufficiency bound to the number of pooled tests required by any maximum a-posteriori (MAP) decoder.
-This bound applies to any arbitrary underlying correlation between infected items.
+We provide an analytical sufficiency bound for the number of pooled tests needed by any maximum a-posteriori (MAP) decoder.
+This bound holds for any correlation between infected items.
+It also applies to multiple access communication systems, where $K$ out of $N$ users transmit information simultaneously.
+This is considered in [34, 35, 15] without prior statistical information.
+In contrast, existing analyses with prior statistics assume all $N$ users transmit at the same time [36, 37, 38, 39, 40, 41, 42].
 Our numerical results demonstrate that in practical regimes for COVID-19 [43, 44] and sparse signal recovery in signal processing [9, 10, 11, 13], the low computational complexity MSGT algorithm proposed herein can reduce in the scenarios tested the number of pool tests by at least $25\%$.
@@
 When there is no memory in the process, $\tau=0$ and $\{\pi_{i}\}_{i=1}^{N}$ hold all the prior information.
-$\mathbf{\Phi}_{i}[l,k]=P\left(U_{i}=s_{k}\lvert\left(U_{i-\tau},...,U_{i-1}\right)=s_{l}\right)$, where $s_{l}$ is the binary representation of $l$ by a $\tau$-length binary vector and $s_{k}\in\{0,1\}$ (For example see Fig.
-6(b)).
+$\mathbf{\Phi}_{i}[l,k]=P\left(U_{i}=s_{k}\lvert\left(U_{i-\tau},...,U_{i-1}\right)=s_{l}\right)$, where $s_{l}$ is the binary representation of $l$ by a $\tau$-length binary vector and $s_{k}\in\{0,1\}$ (for example, see Fig.
+7(b)).
 For the non-adaptive GT, the testing matrix $\mathbf{X}\in\left\{0,1\right\}^{T\times N}$ is defined such that each row corresponds to a single pool test, and each column corresponds to a single item.
@@
 $Y_{i}=\bigvee_{j\in\mathcal{K}}X_{i}(j)=\bigvee_{j=1}^{N}X_{i}(j)U_{j}$
-where $\bigvee$ is the Boolean OR function.
+where $\bigvee$ is the Boolean OR function. All vectors are column vectors and the $\log$ operator denotes $\log_{2}$, unless stated otherwise. These definitions are illustrated in Fig. 1.
 Given $\mathbf{X}$ and the outcome vector $Y$, the recovery success criterion in GT can be measured using various metrics [14].
@@
 $P_{s}^{partial}\triangleq\frac{1}{\binom{N}{K}}\sum_{\mathcal{K}:\left|\mathcal{K}\right|\neq K}\frac{\left|\hat{\mathcal{K}}\cap\mathcal{K}\right|}{K}.$
-To conclude, for a sparse subset of infected items of size $K$, form $N$, the goal in non-adaptive GT with correlated prior data items, is to design a $T\times N$ testing matrix and an efficient and practical recovery algorithm that can exploit correlated priors, such that by observing $Y^{T}$ we can identify the subset of infected items with high probability and with feasible computational complexity.
-Thus, given the knowledge of the correlated prior data items and the available computational resources, the test designer could design the testing matrix and a recovery algorithm to maximize the success probability.
+In conclusion, the goal of non-adaptive GT with correlated prior data is to identify the $K$ defective items out of the $N$ items, where there is correlation between different infected items.
+To achieve this, one needs to design a $T\times N$ testing matrix and an efficient and practical recovery algorithm that can exploit correlated priors.
+The algorithm can identify the subset of infected items by observing $Y^{T}$ with high probability and with feasible computational complexity.
+Given the knowledge of the correlated prior data items and the available computational resources, one can design the testing matrix and a recovery algorithm to maximize the success probability.
 III Main Results
@@
 This reduction is guaranteed by new analytical results we derive.
-In the second stage, the algorithm employs a novel adaptation of the List Viterbi Algorithm (LVA) [32], designed for GT to enable accurate low-complexity recovery using fewer tests by exploiting the correlated prior information.
+In the second stage, the algorithm employs a novel adaptation of the LVA [32], designed for GT to enable accurate low-complexity recovery using fewer tests by exploiting the correlated prior information.
 Additionally, we derive a bound to ensure the low complexity of the entire algorithm.
@@
 If $Y(i)=0$ for some $i\in\{1,\ldots,T\}$, the algorithm eliminates all items participating in the $i$-th test from being defective, and outputs them as the set $\mathcal{P}^{(DND)}\subset\mathcal{N}$.
-In the second step of Stage 1, we use Definite Defectives (DD) algorithm [19] (line 3, Fig.
+In the second step of Stage 1, we use the Definitely Defectives (DD) algorithm [19] (line 3, Fig.
 2.(b)), which goes over the testing matrix and the test result, looking for positive pool tests that include only one possibly defective item.
@@
 3.(a)).
-In the state space, the population sequence, $\mathbf{U}$, is parallel to time steps considered traditionally in VA for communications problems [32], and there are two possible states per item, the first indicates “non-defective” and the second indicates “defective”.
+In the state space, the population sequence, $\mathbf{U}$, is parallel to time steps considered traditionally in Viterbi Algorithm (VA) for communications problems [32], and there are two possible states per item, the first indicates “non-defective” and the second indicates “defective”.
 In the second step, the suggested LVA for GT goes over the sequence of items and outputs $\mathbf{Z}\in\{0,1\}^{L\times N}$, a list of the $L$ most likely trajectories in the state space (line 7, Fig.
@@
 As shown above, MSGT’s Stage 2 can successfully decode $\mathbf{U}$ with just $T=5$, as using the LVA step it narrows down to only 2 possible trajectories, and then the final estimation is selected based on the given prior information and the insights gained in Stage 1.
-(a) Upper bound for possibly defective items after DND for $N=10000$, and $K=15$.
-(b) Lower bound for definitely defective items after DD for $N=10000$, and $K=15$.
-(c) Minimum $\gamma$ parameter to satisfies (3) for $N=500$, and $K=3$.
-Figure 4: Numerical evaluation for theoretical results and bounds.
-The results in (a), (b), and (c) are over 1000 iterations.
-For ML Upper Bound (UB), $T_{ML}=(1+\epsilon)K\log_{2}N$, for any $\epsilon>0$ [17].
-In particular, $\epsilon=0.25$ in the results presented herein.
 III-B Analytical Results
@@
 The two following theorems give the upper bound on the expected number of possible defective items and a “worst-case” upper bound after the first step of Stage 1.
-Theorem 1 ([4]).
+Theorem 1 (Cohen et al. [4]).
 Consider a group test with a Bernoulli testing matrix with $p=\ln{2}/K$, and $T$ tests as $K\rightarrow\infty$.
@@
 Proof.
-According to [42, Lemma 2]:
+According to [51, Lemma 2]:
 $\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]=K+\left(N-K\right)P_{e}^{(DND)},$ (1)
@@
 Consider a group test with a Bernoulli testing matrix and DND decoder. As $K\rightarrow\infty$, the worst-case error probability of DND is bounded by
-$P_{e,u}^{(DND)}$ $\leq P_{e,a}^{(DND)}$
-$+\left(1+\frac{\left(\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)\right)^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1},$
+$P_{e,u}^{(DND)}$ $\leq P_{e,a}^{(DND)}+\left(1+\frac{\left(\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)\right)^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1},$
 for any $\gamma\geq 1$.
@@
 Thus, an upper bound on the expected number of defective items that the DD algorithm successfully detects can be established as follows:
-$\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]=\mathbb{E}\left[P_{s}^{(DD)}\right]K\geq\\ \left(1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}\right)K.$
+$\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]=\mathbb{E}\left[P_{s}^{(DD)}\right]K\geq\left(1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}\right)K.$
 This completed the theorem proof. ∎
@@
 Proof.
-We follow the lines of the success probability analysis in [36, Theorem 1].
+We follow the lines of the success probability analysis in [45, Theorem 1].
 Consider a list decoder that, given the test outcome $Y$, outputs $\mathcal{L}(Y)$, a list of items estimated as defective, such that $\lvert\mathcal{L}(Y)\lvert\geq K$.
@@
 $T\geq\log{P_{s}}+\log{\binom{N}{K}}-\log{\binom{N-K}{L-K}}.$
-It is important to note that unlike [36, Theorem 1], we halt the success probability analysis here without deriving a converse, and instead focus on the connection between $T$ and $\gamma$.
+It is important to note that unlike [45, Theorem 1], we halt the success probability analysis here without deriving a converse, and instead focus on the connection between $T$ and $\gamma$.
 This analysis is valid for any list decoder, particularly for an ML decoder [17, 53], allowing us to set $P_{s}=\eta$.
-Since MSGT uses MAP, the non-uniform and correlated prior information allows us to achieve better recovery performance for the same $T$ and $L$ compared to ML.
+Since MSGT uses MAP, the non-uniform and correlated prior information allows us to achieve better recovery performance for the same $T$ and $L$ compared to ML [^2] In Section IV-C, we provide sufficiency bound for any MAP decoder with an arbitrary underlying correlation between infected items that outperform the ML sufficiency bound and illustrate it for practical scenarios in Section V..
 Moreover, since for any $\gamma\geq 1$, $\gamma K$ is the maximum list size of estimated defective items we allow in MSGT.
-By substituting $L=\gamma K$, we establish a lower limit for $\gamma$ that guarantees MSGT’s success probability remains at least equal to that of the ML estimator.
+By substituting $L=\gamma K$, we establish a lower bound for $\gamma$ that guarantees MSGT’s success probability remains at least equal to that of the ML estimator.
 ∎
@@
 We begin by analyzing the complexity of each step of the proposed MSGT solution given in Algorithm 1, and finally, we sum everything up to determine the total complexity.
-In Stage 1, the complexity of DND is $\mathcal{O}\left(KN\log_{2}{N}\right)$ as analyzed in [45, Remark 6].
+In Stage 1, the complexity of DND is $\mathcal{O}\left(KN\log_{2}{N}\right)$ as analyzed in [54, Remark 6].
 Then, for each positive entry of the test result vector $Y^{T}$, the DD algorithm counts the number of possibly defective items that participate in the corresponding pool test.
 That requires $K\left|\mathcal{P}^{(S_{1,1})}\right|\log_{2}{N}$ computations, and for simplification, we bound it by the DND complexity, i.e., $\mathcal{O}\left(KN\log_{2}N\right)$.
-(a) $N=500$ and $K=3$.
-(b) $N=1024$ and $K=8$.
-(c) $N=10000$ and $K=13$.
-Figure 5: Success probability of MSGT, MAP, ML and DD over 1000 iterations.
-A comparison to ML and MAP is not presented in (b) and (c), as they are not feasible for populations of those sizes due to the computational complexity burden.
 In Stage 2, parallel LVA requires $L$ times more computations than the VA [32].
-VA calculates all the possible transition probabilities for each step in the sequence.
+The VA calculates all the possible transition probabilities for each step in the sequence.
 In GT, this sequence is the order’s items sequence, where with the suggested algorithm, it is enough to consider only the $\left|\mathcal{P}^{\left(S_{1,2}\right)}\right|$ items as the sequence steps.
@@
 For the average case, we use the expectation bounds from Theorems 1 and 3, such that:
-$\mathbb{E}\left[\left|\mathcal{P}^{\left(S_{1}\right)}\right|\right]=\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]-\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]$
-$\leq\left(N-K\right)P_{e,a}^{(DND)}+N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}K.$
+$\mathbb{E}\left[\left|\mathcal{P}^{\left(S_{1}\right)}\right|\right]=\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]-\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]\leq\left(N-K\right)P_{e,a}^{(DND)}+N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}K.$
 Accordingly, the number of the required computations for LVA step in MSGT is bounded by
@@
 By substituting Theorem 2 it follows that the number of possibly defective items in the worst-case is upper bounded by
-$\begin{split}K&+\left(N-K\right)\left(P_{e,a}^{(DND)}+\right.\\ &\left.\left(1+\frac{g^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1}\right),\end{split}$ (7)
+$\begin{split}K&+\left(N-K\right)\left(P_{e,a}^{(DND)}+\left(1+\frac{g^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1}\right),\end{split}$ (7)
 for $g=\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)$.
@@
 when the DND and DD are executed as prior steps, and otherwise it is $\mathcal{O}\left(\binom{N}{K}KN\log_{2}N\right)$.
-Note that from Theorem 4 and Remark 1, it follows that the proposed MSGT algorithm performs $\mathcal{O}\left(\frac{1}{L}\left(\frac{N}{\gamma K}\right)^{K}\right)$, times fewer computational operations compared to MAP.
+Note that from Theorem 4 and Remark 1, it follows that the proposed MSGT algorithm performs $\mathcal{O}\left(\frac{1}{L}\left(\frac{N}{\gamma K}\right)^{K}\right)$ times fewer computational operations compared to MAP.
 III-C Discussion
@@
 The simple reduction of the search space in Stage 1 enables MSGT to handle challenging regimes with a small number of tests.
-Stage 2, particularly the LVA step, contributes to its high success probability.
-Additionally, using the parallel implementation of LVA, rather than the iterative one, keeps the complexity low [32].
+Stage 2, and the LVA step in particular, contributes to its high success probability.
+Furthermore, using the parallel implementation of LVA, rather than the iterative one, keeps the complexity low [32].
 It is important to note that, as explained in [32], achieving optimal results is ensured only with a very large $L$, inevitably leading to complexity equivalent to MAP’s.
@@
 Moreover, it is shown that MSGT addresses practical regimes, e.g., in COVID-19 detection [55] ($T=48$ for $(N,K)=(384,4)$), in communications [56] ($(N,K)=(10^{5},6)$), and in GT quantizer [8] ($(N,K)=(1024,16)$ [57]).
-Another aspect of novelty in this work is the integration of Viterbi Algorithm into the GT problem.
+Another aspect of novelty in this work is the integration of the VA into the GT problem.
 In the context of Markovian priors, one can think of the population’s sequence of items as a sequence of observations stemming from a hidden Markov process within a given Markov model over $N$ steps.
@@
 However, the most likely sequence of items does not necessarily include $K$ defective items.
-Particularly, in sparse signal scenarios, which is the focus of GT, the most likely sequence typically involves the minimum number of defectives that explain the observations.
+Particularly in sparse signal scenarios, which is the focus of GT, the most likely sequence typically involves the minimum number of defectives that explain the observations.
 As a result, VA may not necessarily detect more defective items than already known and may detect even more than $K$.
@@
 Hence, such modification will increase the complexity compared to [4].
-(a) Number of computational operations.
-(b) Correlation matrix $\mathbf{\Phi}_{i}$ for $\tau=2$.
-(c) Long memory priors - exact recovery.
-Figure 6: (a) Number of computational operations in MSGT and MAP as given in Theorem 4 and Remark 1, respectively.
-(b) Example of correlation matrix $\mathbf{\Phi}_{i}$.
-(c) Probability of success of MSGT with exact prior statistics of 3-memory-steps Markov process, and with limited prior statistics assuming the Markov process has only one memory step.
-$N=1024,K=8$, 1000 iterations.
 IV MAP Analytical Bound for GT with General Correlated Prior Statistics
@@
 This bound also applies to the LVA-based MSGT algorithm, which can approach MAP decoder performance for a sufficiently large list [32].
-We calculate this bound by extending the known result of maximum likelihood (ML) decoding for Group Testing (GT) problems [17], which bounds the probability of error of a ML decoder using a Gallager-type bound [44] for multiple access channel (MAC) channels [50].
+We calculate this bound by extending the known result of ML decoding for GT problems [17], which bounds the probability of error of a ML decoder using a Gallager-type bound [53] for multiple access channel (MAC) channels [59].
 Recall that a noiseless GT setup can be considered as a noiseless MAC setting with $N$ “users”, each having a single codeword.
@@
 We first introduce new notations and definitions in Subsection IV-A, then provide the probability of error to bound the MAP performance in Subsection IV-B, characterizing in Subsection IV-C the maximal number of tests required to guarantee reliability as given in Theorem 5.
+Finally, in Subsection IV-D, we evaluate this bound in Gilbert-Elliott (GE) model [60] and illustrate it for a practical regime in Section V.
 IV-A Definitions and notations
@@
 Let $H(\cdot)$ and $I(\cdot;\cdot)$ denote Shannon’s entropy and mutual information, respectively [61].
-All vectors are column vectors and the $\log$ operator denotes $\log_{2}$, unless stated otherwise.
 Let $\omega\in\left[\binom{N}{K}\right]$ denote the index of a specific set of $K$ items out of the $N$ items.
@@
 where the error exponent $E_{0}(\rho)$ is given by
-$\begin{split}&E_{0}(\rho)=\\ &-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$
+$\begin{split}&E_{0}(\rho)=-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$
 the correlated prior statistics function of the items $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$ is given by
 $\begin{split}E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)=\log\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}},\end{split}$
-for $0\leq\rho\leq 1$.
+for $0\leq\rho\leq 1$, where $\mathcal{T}\mathcal{P}_{i}=\left\{S_{1}:\lvert S_{1}\rvert=K-i\right\}$, $\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)=\left\{S_{2}:\lvert S_{2}\rvert=i,\;S_{2}\cap S_{1}=\emptyset\right\}$, and $S_{\omega}=S_{1}\cup S_{2}$.
 Proof of Lemma 1: The proof is given in Appendix B. ∎
@@
 $\frac{\partial E_{o}}{\partial\rho}|_{\rho=0}$
-$=$ $\sum_{Y}\sum_{X_{TP}}[\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\log P\left(Y,X_{TP}|X_{FP}\right)$
--- $\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})]$
-$=$ $\sum_{Y}\sum_{X_{TP}}\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})$
-$\log\frac{p(Y,X_{TP}|X_{FP})}{\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})}=I(X_{FP};X_{TP},Y).$
+$=$ $\sum_{Y}\sum_{X_{TP}}[\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\log P\left(Y,X_{TP}|X_{FP}\right)-\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})]$
+$=$ $\sum_{Y}\sum_{X_{TP}}\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})\log\frac{p(Y,X_{TP}|X_{FP})}{\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})}$
+$=$ $I(X_{FP};X_{TP},Y).$
 And also, in Appedix C, that
 $\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}=H(P_{S_{2}|S_{1}}).$
-This is obtained directly from [59] as also elaborated in [57, Chap. 1]. Note that $E_{o}(0)=0$ and $E_{s,j}(0,P_{W})=0$, hence we have
-$T\emph{f}(\rho)=$ $T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)-\frac{H\left(P_{S_{2}|S_{1}}\right)}{T}\Bigg{)}$
--- $\log\frac{1}{\binom{K}{i}}+T\frac{\rho^{2}}{2}E_{o}^{\prime\prime}\left(\psi\right)-T\frac{\rho^{2}}{2}E_{s,j}^{\prime\prime}\left(\psi\right),$
+This is obtained directly from [62] as also elaborated in [42, Chap. 1]. Note that $E_{o}(0)=0$ and $E_{s,j}(0,P_{W})=0$, hence we have
+$T\emph{f}(\rho)=$ $T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)-\frac{H\left(P_{S_{2}|S_{1}}\right)}{T}\Bigg{)}-\log\frac{1}{\binom{K}{i}}+T\frac{\rho^{2}}{2}E_{o}^{\prime\prime}\left(\psi\right)-T\frac{\rho^{2}}{2}E_{s,j}^{\prime\prime}\left(\psi\right),$
 and, if
@@
 for some $\varepsilon>0$, we have
-$T\emph{f}(\rho)\geq-\log\frac{1}{\binom{K}{i}}$
-$+T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)+\frac{\rho}{2}\left(E_{o}^{\prime\prime}(\psi)-E_{s,j}^{\prime\prime}\left(\psi\right)\right)\Bigg{)}$
+$T\emph{f}(\rho)\geq-\log\frac{1}{\binom{K}{i}}+T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)+\frac{\rho}{2}\left(E_{o}^{\prime\prime}(\psi)-E_{s,j}^{\prime\prime}\left(\psi\right)\right)\Bigg{)}$
 Moreovver, note that $E_{o}^{\prime\prime}\left(\psi\right)$ is negative [53]. However, it is independent of the other constants and $T$, hence choosing
@@
 In practice, the expression $I(X_{FP};X_{TP},Y)$ in (11) is important to understand how many tests are required, yet it is not a function of the problem parameters and is bounded in [17] and [54] to get a better handle on $T$.
-Claim 1 ([17] and [45]).
-For large $K$, and under a fixed input distribution for the testing matrix $\left(\frac{\ln(2)}{K},1-\frac{\ln(2)}{K}\right)$, the mutual information between $X_{FP}$ and $(X_{TP},Y)$ is bounded by $I(X_{FP};X_{TP},Y)\geq\frac{i}{K}$.
+Claim 1 ([17]).
+For large $K$, and under a fixed input distribution for the testing matrix $\left(\frac{\ln(2)}{K},1-\frac{\ln(2)}{K}\right)$, the mutual information between $X_{FP}$ and $(X_{TP},Y)$ is bounded by
+$I(X_{FP};X_{TP},Y)\geq\frac{i}{K}.$
 Hence, by applying Clain 1 in Eq. (11), we achieves the sufficiency bound on $T$ as provided in Theorem 5,
@@
 such that, as $N\rightarrow\infty$ the average error probability approaches zero, and reliability is established.
+(a) Upper bound for possibly defective items after DND for $N=10000$, and $K=15$.
+(b) Lower bound for definitely defective items after DD for $N=10000$, and $K=15$.
+(c) Minimum $\gamma$ parameter to satisfies (3) for $N=500$, and $K=3$.
+Figure 4: Numerical evaluation for theoretical results and bounds.
+The results in (a), (b), and (c) are over 1000 iterations.
+For ML Upper Bound (UB), $T_{ML}=(1+\epsilon)K\log_{2}N$, for any $\epsilon>0$ [17].
+In particular, $\epsilon=0.25$ in the results presented herein.
+IV-D Gilbert Elliott calculations
+Consider a GE model [60].
+It is a Markov process with two states.
+In the context of GT, as also recently considered in [63], we call them infected (state 1) and not infected (state 0).
+The probability of transition from state 0,1 to state 1,0 is denoted by $q,s$, respectively.
+The probability that the first state is in state 1 is denoted by $\pi\triangleq P\left(U_{i}=1\right)$.
+Unless stated otherwise, we assume the system is in steady state, that is, the probability of the system to be at state 1 at any point is given by $\pi=\frac{q}{q+s}$.
+To evaluate Theorem 5, we need to evaluate
+$H\left(P_{S_{2}|S_{1}}\right)={\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)\log\frac{P\left(S_{1}\right)}{P_{W}\left(S_{1},S_{2}\right)}\right),$
+where $P\left(S_{1}\right)=\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)$.
+Let $S_{\omega}=\left(S_{1},S_{2}\right)$ denote a specific set of $K$ items.
+Recall that $P_{W}\left(S_{\omega}\right)$ calculates the probability to get $S_{\omega}$ as the infected set, given that there are $K$ infected items.
+Let $\mathcal{G}\mathcal{E}_{N}$ denote a GE process that runs for $N$ steps.
+Let $\left\{U_{i}\right\}_{i=1}^{N}$ denote the states of the process during the $N$ steps.
+Let $\lvert\mathcal{G}\mathcal{E}_{N}\rvert\triangleq\sum_{i=1}^{N}U_{i}$.
+While our analysis is conditioned on the assumption that exactly $K$ items are infected, a general $\mathcal{G}\mathcal{E}_{N}$ process may result in any number of infected items between 0 and $N$.
+We therefore analyze and simulate a special class of GE processes; GE processes that reach the infected state exactly $K$ times after $N$ steps.
+Such a GE process that is guaranteed to produce exactly $K$ infected items is denoted by $\mathcal{G}\mathcal{E}_{N,K}$.
+Then $P_{W}\left(S_{\omega}\right)$ of $\mathcal{G}\mathcal{E}_{N,K}$ is given as follows
+$P_{W}\left(S_{\omega}\right)$ $=$ $P\left(\mathcal{G}\mathcal{E}_{N,K}\text{ results in }S_{\omega}\right)$ (12)
+=$\overset{(a)}{=}$ $P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\lvert\;\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)$
+=$\overset{(b)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega},\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}$
+=$\overset{(c)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)},$
+where (a) follows from the definition of $\mathcal{G}\mathcal{E}_{N,K}$, (b) from the definition of conditional probability, and (c) follows since $S_{\omega}$ contains exactly $K$ infected items.
+IV-D1 Numerator of Eq.(12)
+The numerator evaluates the probability that a standard GE process results in a specific pattern of states, where state one is achieved only at the states in $S_{\omega}$.
+Recall that GE is a Markov chain, and so $P\left(S_{\omega}\right)=P\left(U_{1}\right)\cdot\prod_{i=2}^{N}P\left(U_{i}\lvert U_{i-1}\right)$.
+This can be readily calculated using the definitions of $q,s,\pi$, as described at the beginning of this section
+$P\left(U_{1}\right)=\begin{cases}1-\pi&\text{if }1\not\in S_{\omega}\\ \pi&\text{if }1\in S_{\omega}\end{cases},$
+and
+$P\left(U_{i}\lvert U_{i-1}\right)=\begin{cases}1-q&\text{if }i\not\in S_{\omega},i-1\not\in S_{\omega}\\ q&\text{if }i\in S_{\omega},i-1\not\in S_{\omega}\\ s&\text{if }i\not\in S_{\omega},i-1\in S_{\omega}\\ 1-s&\text{if }i\in S_{\omega},i-1\in S_{\omega}\end{cases}.$
+IV-D2 Denominator of Eq.(12)
+We create a new Markov chain that represents $\sum_{i=1}^{n}U_{i}$ as $n$ advances from $1$ to $N$.
+This Markov chain keeps track of how many times the GE process was in state 1.
+Hence this Markov chain has $2\left(N+1\right)$ states.
+State $n,n^{\prime}\in\left\{0,\ldots,N\right\}$ denotes a state where the GE process was in the infected state (state 1) $n$ times, and the current state is 0,1, respectively.
+An example of this Markov chain is shown in Fig.
+6.
+For convenience, the states $0^{\prime},N$ are also listed, even though they cannot be achieved.
+This Markov chain’s states can be arranged in a vector, where the $j$-th state represents $\sum_{i=1}^{N}U_{i}=\left\lfloor j/2\right\rfloor$, and even/odd $j$ represents the current state is not/is 00.
+For instance, in the example of Fig.
+6, the states $6/7$ represent states 2/$2/2^{\prime}$ in the figure.
+The initial state vector is given as by $\left(1-\pi,0,0,\pi,0,\ldots,0\right)$.
+The transition matrix is given as follows
+$P=\begin{bmatrix}1-q&0&0&q&0&0&\dots&0\\ 0&0&0&0&0&0&\dots&0\\ 0&0&1-q&0&0&q&\dots&0\\ 0&0&s&0&1-s&0&\dots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&0&0&\dots&0\end{bmatrix}.$
+(a) $N=500$ and $K=3$.
+(b) $N=1024$ and $K=8$.
+(c) $N=10000$ and $K=13$.
+Figure 5: Success probability of MSGT, MAP, ML and DD over 1000 iterations.
+A comparison to ML and MAP is not presented in (b) and (c), as they are not feasible for populations of those sizes due to the computational complexity burden.
+The probabilities to be in any state after $N$ steps is therefore given by
+$\left(1-\pi,0,0,\pi,0,\ldots,0\right)\cdot P^{N-1}.$ (13)
+To conclude the calculation, the probability to have exactly $K$ infected items after $N$ steps can be obtained by summing items $12K+$ 1 and $22K+$ 2 of the vector in (13).
+Figure 6: Markov chain for the number of infected items of a GE model, $N=4$.
+In the next section, we evaluate the proposed sufficient bound given in Theorem 5 for MAP decoder in the GE setting considered herein.
 V Numerical Evaluation
@@
 First, in Subsection V-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III.
-Then, in Subsection V-2, we contrast the performance of MSGT with the one of DD, ML, and MAP in a practical regime of $N$ and $K$.
-To generate the correlated prior information between adjacent items, we use the Gilbert-Elliot (GE) model [60] (also considered very recently analytically in [61]), which is a stationary Markov chain of order 1 with two states: one representing an error phase and the other an error-free phase.
+Then, in Subsection V-2, we contrast the performance of MSGT with those of DD, ML, and MAP in a practical regime of $N$ and $K$.
+To generate the correlated prior information between adjacent items, we use the special GE process that reaches the infected state exactly $K$ times after $N$ steps, as described and analyzed in Section IV-D.
 The GE model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$.
@@
 In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [43, 44], or in sparse signal recovery in signal-processing with fixed vector size of input samples [9, 10, 11, 13, 64]), we show that the low computational complexity MSGT algorithm can reduce the number of pool tests by at least $25\%$.
+(a) Number of computational operations.
+(b) Correlation matrix $\mathbf{\Phi}_{i}$ for $\tau=2$.
+(c) Long memory priors - exact recovery.
+Figure 7: (a) Number of computational operations in MSGT and MAP as given in Theorem 4 and Remark 1, respectively.
+(b) Example of correlation matrix $\mathbf{\Phi}_{i}$.
+(c) Probability of success of MSGT with exact prior statistics of 3-memory-steps Markov process, and with limited prior statistics assuming the Markov process has only one memory step.
+$N=1024,K=8$, 1000 iterations.
 V-1 Theoretical Analysis
@@
 Substituting those priors, it follows that the converse of our problem is
-$T\geq H(U_{1})+\sum_{i=2}^{N}H\left(U_{i}\lvert U_{i-1}\right).$
-This bound is illustrated in the practical scenarios tested in Fig. 4 and Fig. 5.
+$T\geq H(U_{1})+\sum_{i=2}^{N}H\left(U_{i}\lvert U_{i-1}\right).$ (14)
+This bound is illustrated in the practical scenarios tested in Fig.
+4 and Fig.
+5.
+In Fig 8, we compare the lower and the upper bounds with GE model.
+The lower and upper bounds for GT with correlated prior statistics are given in (14) and in Section IV-D, respectively, while for GT without prior statistical information, the bounds are as given in [17].
+Figure 8: Lower and upper bounds for GT with GE model (left). MSGT and corresponding MAP upper bound for GT with GE model (right). For MAP and ML upper bounds $\varepsilon=0.6$.
 V-2 Algorithm Evaluation
@@
 5 compares MSGT to MAP, ML, and DD algorithms.
-We run DND and DD before ML and MAP for reasonable runtime and memory consumption.
+We run DND and DD before ML and MAP for feasible runtime and memory consumption.
 The population includes $N\in\{500,1024,10000\}$ items and $K\in\{3,8,13\}$ defective items, respectively, and the empirical success probability is the average over $1000$ experiments.
@@
 The results are shown in Fig.
-6(c).
+7(c).
 It is evident that utilizing the prior of long memory improves the success probability by $10\%$ in this scenario.
 We do note again that for the practical regime tested as in [55, 56, 8, 57], i.e., $N=1024$ and $K=8$, it is no longer possible to compare the performance since, unlike the efficient proposed MSGT algorithm, ML and MAP decoders for GT become infeasible.
-VI Discussion and Future Work
-To the best of our knowledge, the existing solutions in the literature do not offer an efficient and general solution for the regime below $T_{ML}$.
-In this work, we focus on this regime of the number of tests (i.e., $T<T_{ML}$), where the proposed MSGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of the potential defective combinations.
+VI Conclusions and Future Work
+In this work, we propose a MAP decoder for GT problems with arbitrary correlated sources.
+We analyze the performance of the MAP decoders for this case and provide, to the best of our knowledge, the only analysis that goes below $T_{ML}$, as the existing solutions in the literature do not offer an efficient and general solution for this regime.
+Our proposed MSGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of potential defective combinations.
 This approach yields an efficient computational solution that, as demonstrated in our simulation result for practical scenarios, can approach the minimum number of tests as MAP algorithm.
+Future work includes the study of the proposed algorithm and analytical bounds in communication systems [36, 37, 38, 39, 40, 41, 42] and signal processing applications [23, 13, 24], in which correlation between different frequencies, time signals, or among different sensors, etc., can also be utilized to achieve more precise estimations.
 VII Acknowledgment
@@
 9:return $\mathcal{P}^{(DD)}$
-A-C List Viterbi Algorithm (LVA)
+A-C Priors Update
+The algorithm translates the information gathered during the first stage of MSGT into the state space.
+In particular, if an element is detected as defective, its probability of being defective is updated to $1$.
+Additionally, the transition probabilities are modified so that the transition to the defective state becomes deterministic.
+Algorithm 4 updatePriors
+Input: $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N},\mathcal{P}^{(S_{1,2})},\mathcal{P}^{(DD)}$
+Output: $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$
+1:for each $i\in\mathcal{P}^{(\mathrm{DD})}$ do
+2:    $\pi_{i}\leftarrow 0$
+3:    if $i\neq 1$ then
+4:         $\boldsymbol{\Phi}_{i-1}[:,0]\leftarrow\boldsymbol{0}$
+5:         $\boldsymbol{\Phi}_{i-1}[:,1]\leftarrow\boldsymbol{1}$
+6:    end if
+7:end for
+8:return $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$
+A-D List Viterbi Algorithm (LVA)
 Our variations of the LVA algorithm [32], outlined in Algorithm 5, is designed to return the $L$ most likely sequences, representing the estimation of the whole population, for a given $L\geq 1$.
@@
 If $\tau>1$, and a further processing is applied to map the states $\{1,...,2^{\tau}\}$ to the binary states $\{0,1\}$, representing “defective” and “non-defective”.
-A-D Maximum A Posteriori (MAP)
+A-E Maximum A Posteriori (MAP)
 The MAP estimator [53], as shown in Algorithm 6, returns the set $c\in\mathbf{C}$ with the highest maximum a posteriori probability among all the sets, that is, the set $c$ that obtains the maximum $P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)$.
@@
 $\overset{(a)}{\leq}\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\frac{P\left(Y^{T}\lvert\mathbf{X}_{S_{\omega^{\prime}}}\right)^{s}P_{W}\left(S_{\omega^{\prime}}\right)^{s}}{P\left(Y^{T}\lvert\mathbf{X}_{S_{\omega^{\star}}}\right)^{s}P_{W}\left(S_{\omega^{\star}}\right)^{s}}$
-$=\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\cdot$
-$\hskip 14.22636pt\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
+$=\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\cdot\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
 $\overset{(b)}{\leq}\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}},\hskip 5.69046pt\forall s>0.$
@@
 $\leq\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
-$=\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}}{P_{W}\left(\omega^{*}\right)^{s}}\cdot$
-$\hskip 71.13188pt\frac{\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}$
-$\overset{(a)}{\leq}\Biggl{[}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)}{P_{W}\left(\omega^{*}\right)^{s}}\cdot$
-$\hskip 71.13188pt\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}\Biggr{]}^{\rho}$
-$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot$
-$\hskip 71.13188pt\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}},$
+$=\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}$
+$\overset{(a)}{\leq}\left[\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}\right]^{\rho}$
+$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}},$
 $\qquad\forall s>0,\rho\in\left[0,1\right]$
@@
 $=\sum_{\mathbf{X}_{S_{\omega^{\star}}}}\sum_{Y^{T}}P\left(\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
-$=\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(\mathbf{X}_{TP},Y^{T}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot$
-$\hskip 22.76228ptP\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
-$\overset{(a)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot$
-$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)\cdot$
-$\hskip 39.83368ptP\left(\mathbf{X}_{FN}\right)\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}}$
-$\overset{(b)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right))}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot$
-$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot$
-$\hskip 45.52458pt\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right))^{\frac{\rho}{1+\rho}}}$
-$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot$
-$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot$
-$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot$
-$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},$
+$=\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(\mathbf{X}_{TP},Y^{T}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
+$\overset{(a)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)\cdot P\left(\mathbf{X}_{FN}\right)\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}}$
+$\overset{(b)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right))}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right))^{\frac{\rho}{1+\rho}}}$
+$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},$
 in $(a)$ we substitute Eq.
@@
 $P\left(E_{i}^{j}\right)=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)P\left(E_{i}^{j}|W=\omega^{*}\right)$
-$\leq\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}$
-$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{1-\frac{\rho}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}$
-$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}$
-$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-=$\overset{(a)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}\right]$
-$\hskip 22.76228pt{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}$
-$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-$\overset{(b)}{\leq}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}$
-$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-=$\overset{(c)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}$
-$\hskip 19.91684pt\left[\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}\right]^{T}$
+$\leq\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{1-\frac{\rho}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+=$\overset{(a)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}\right]{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$\overset{(b)}{\leq}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+=$\overset{(c)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}\left[\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}\right]^{T}$
 =$\overset{(d)}{=}2^{-T\left(E_{0}(\rho)-\frac{\frac{1}{\binom{K}{i}}E_{s,j}(\rho,P_{W})}{T}\right)},$
-where (a) follows from Lemma 2 with $\alpha=\frac{1}{1+\rho}$, (b) follows from the fact that for a fixed $S_{1}\in\mathcal{T}\mathcal{P}_{i}$, ~$\tilde{\mathcal{I}}^{j}_{dif}\subset\left\{\omega=\left(S_{1},S_{2}\right):S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)\right\}$, (c) follows since $X_{FP},X_{TP}$ represent a single column of $\mathbf{X}_{FP},\mathbf{X}_{TP}$ respectively, as the elements of $\mathbf{X}$ are drawn in i.i.d.
-fashion, and (d) follows for
-$\begin{split}&E_{0}(\rho)=\\ &-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$
+where (a) follows from Lemma 2 with $\alpha=\frac{1}{1+\rho}$, (b) follows from the fact that for a fixed
+$S_{1}\in\mathcal{T}\mathcal{P}_{i},\quad\tilde{\mathcal{I}}^{j}_{dif}\subset\left\{\omega=\left(S_{1},S_{2}\right):S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)\right\},$
+(c) follows since $X_{FP},X_{TP}$ represent a single column of $\mathbf{X}_{FP},\mathbf{X}_{TP}$ respectively, as the elements of $\mathbf{X}$ are drawn in i.i.d. fashion, and (d) follows for
+$\begin{split}&E_{0}(\rho)=-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$
 and
@@
 In this section, we show the derivative of $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$ for $\rho=0$.
-This result is obtained directly from [59] as also elaborated in [57, Chap.
+This result is obtained directly from [62] as also elaborated in [42, Chap.
 1].
@@
 Thus,
-$\frac{\partial}{\partial\rho}\left(A\right)=A\frac{\partial}{\partial\rho}\left(\log\left(A\right)\right)=A\frac{\partial}{\partial\rho}\left(\left(1+\rho\right)\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)\right)$
+$\frac{\partial}{\partial\rho}\left(A\right)=A\frac{\partial}{\partial\rho}\left(\log\left(A\right)\right)$
+$=A\frac{\partial}{\partial\rho}\left(\left(1+\rho\right)\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)\right)$
 $=A\left(\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)+(1+\rho)\frac{\partial\tilde{f}\left(\rho,S_{1}\right)}{\partial\rho}\right)$
 =$\overset{(a)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\left(\log(\tilde{f}\left(\rho,S_{1}\right))+(1+\rho)\frac{\partial\tilde{f}\left(\rho,S_{1}\right)}{\partial\rho}\right)$
-=$\overset{(b)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))$
-$\hskip 17.07182pt-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
-=$\overset{(c)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))$
-$\hskip 17.07182pt-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
-=$\overset{(d)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}$
-$\hskip 17.07182ptP_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
+=$\overset{(b)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
+=$\overset{(c)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
+=$\overset{(d)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
 where (a) follows replacing $A$, (b) by derivative of $\tilde{f}\left(\rho,S_{1}\right)$, and (c) replacing $\tilde{f}\left(\rho,S_{1}\right)$.
@@
 $=$ $\frac{1}{\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}$
-$\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}$
-$\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
+$\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
 Next, setting $\rho=0$ in Eq. (C), and since
@@
 $\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}$
-$=$ $\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)$
-$\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)}{P_{W}\left(S_{1},S_{2}\right)}\Biggr{)}$
+$=$ $\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)}{P_{W}\left(S_{1},S_{2}\right)}\Biggr{)}$
 $=$ $\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)\log\frac{P\left(S_{1}\right)}{P_{W}\left(S_{1},S_{2}\right)}\right)$

### v5

@@
-Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics
+arXiv:2402.10018v5[cs.IT] 28 Feb 2026
 Abstract
-In this paper, we propose an efficient multi-stage algorithm for non-adaptive Group Testing (GT) with general correlated prior statistics.
+In this paper, we propose an efficient two-stage decoding algorithm for non-adaptive Group Testing (GT) with general correlated prior statistics.
 The proposed solution can be applied to any correlated statistical prior represented in trellis, e.g., finite state machines and Markov processes.
 We introduce a variation of List Viterbi Algorithm (LVA) to enable accurate recovery using much fewer tests than objectives, which efficiently gains from the correlated prior statistics structure.
-We also provide a sufficiency bound to the number of pooled tests required by any Maximum A Posteriori (MAP) decoder with an arbitrary correlation between infected items.
-Our numerical results demonstrate that the proposed Multi-Stage GT (MSGT) algorithm can obtain the optimal MAP performance with feasible complexity in practical regimes, such as with COVID-19 and sparse signal recovery applications, and reduce in the scenarios tested the number of pooled tests by at least $25\%$ compared to existing classical low complexity GT algorithms.
-Moreover, we analytically characterize the complexity of the proposed MSGT algorithm that guarantees its efficiency.
+We also provide a sufficiency bound to the number of pooled tests required by any Maximum A Posteriori (MAP) decoder with an arbitrary correlation, i.e., dependence between infected items.
+Our numerical results demonstrate that the proposed two-stage decoding GT (2SDGT) algorithm can obtain the optimal MAP performance with feasible complexity in practical regimes, such as with COVID-19 and sparse signal recovery applications, and reduce in the scenarios tested the number of pooled tests by at least $25\%$ compared to existing classical low complexity GT algorithms.
+Moreover, we analytically characterize the complexity of the proposed 2SDGT algorithm that guarantees its efficiency.
 I Introduction
@@
 In contrast, in non-adaptive GT, the whole process is designed in advance.
-Since it was first suggested, the GT problem has been investigated and generalized to many areas and applications, among them disease detection [2, 3, 4], cyber security applications [5], pattern matching algorithms [6] and communication [7, 8].
-All of these applications imply a strong connection between GT and Compressed Sensing (CS) as two methods for sparse signal recovery that share common applications [9, 10, 11, 12].
-The main difference between the two is that CS aims to recover a real-valued signal [13], while GT recovers a binary signal [14] or discrete-values [15, 16, 4].
-Thus, one can consider GT as a Boolean CS [17, 18].
-Traditional GT and its performance (i.e., the tradeoff between the number of pool tests and recovery algorithm complexity), focuses on the probabilistic model in which the items are identically distributed and independent [19, 14].
-Recent research explores cases where prior information about the correlation of objects is available [20, 2, 3, 21, 22].
+Since it was first suggested, the GT problem has been investigated and generalized to many areas and applications, among them disease detection [34, 61, 22, 60], cyber security applications [64], pattern matching algorithms [35] and communication [65, 20].
+All of these applications imply a strong connection between GT and Compressed Sensing as two methods for sparse signal recovery that share common applications [32, 63, 1, 47].
+The main difference between the two is that Compressed Sensing aims to recover a real-valued signal [26], while GT recovers a binary signal [4] or discrete-values [18, 21, 22].
+Thus, one can consider GT as a Boolean Compressed Sensing [8, 55].
+Traditional GT and its performance (i.e., the tradeoff between the number of pooled tests and recovery algorithm complexity), focuses on the probabilistic model in which the items are identically distributed and independent [3, 4].
+Recent research explores cases where prior information about the correlation (i.e., dependence) of objects is available [48, 34, 61, 39, 6].
 The motivation for this approach arises from the fact that correlated prior statistics have the potential to achieve higher recovery rates and reduce the number of required tests.
-In disease detection, leveraging information about the proximity between individuals, represented by contact tracing information or graphs, can lead to significant savings in pool tests [20, 2, 3, 21].
-However, previous GT works presented solutions designed for specific models and applications and may not be extended easily to other models and applications.
+In disease detection, leveraging information about the proximity between individuals, represented by contact tracing information or graphs, can lead to significant savings in pooled tests [48, 34, 61, 39].
+However, previous GT works presented solutions designed for specific models and applications and may not extend easily to other models and applications.
 In numerous signal processing applications, correlation between different frequencies, time signals, or among different sensors can also be utilized to achieve more precise estimations [68, 26, 7].
-Hidden Markov Model (HMM) is a common model for many physical signals, such as speech signals [25], human motion [26], and spectrum occupancy in communication systems [27].
-Infections can be also modeled as HMMs [28].
-For example, [29] presents a GT solution for a specific HMM derived from contact tracing.
+Hidden Markov Model is a common model for many physical signals, such as speech signals [28], human motion [62], and spectrum occupancy in communication systems [31].
+Infections can be also modeled as Hidden Markov Models [14].
+For example, [13] presents a GT solution for a specific Hidden Markov Model derived from contact tracing.
 To the best of our knowledge, no existing solution addresses the GT problem with general Markovian priors and applicable to a wide range of diverse applications.
-In this work, we introduce a practical non-adaptive Multi-Stage Group Testing (MSGT) algorithm for correlated items with prior statistics.
-The proposed multi-stage algorithm employs a new variation of the parallel List Viterbi Algorithm (LVA) [30, 31, 32] we designed for GT to enable accurate low complexity recovery using fewer tests.
-The proposed algorithm can be applied for any statistical prior represented in trellis [33], e.g., finite-state machines and Markov processes.
-Using LVA, MSGT leverages those statistics to estimate the defective set efficiently, even in a regime below the maximum likelihood (ML) upper bound.
+In this work, we addresses the GT problem with general prior statistics.
+We adopt a more general formulation in which the number of defective items is not assumed to be fixed a priori, but is instead induced by the underlying correlated prior distribution.
+In particular, the number of defectives is treated as a random variable, and the system design is based on a typical cardinality interval of the number of defective items.
+We introduce a practical non-adaptive Two-Stage Decoding GT (2SDGT) algorithm for correlated items with prior statistics.
+The proposed two-stage decoding algorithm employs a new variation of the parallel List Viterbi Algorithm (LVA) [5, 43, 56] we designed for GT to enable accurate low complexity recovery using fewer tests.
+The proposed algorithm can be applied to general Markovian correlated priors, which may be represented using trellis structures [46], e.g., finite-state machines and Markov processes.
+While classical group testing assumes exchangeable items, many practical settings admit structured dependencies arising from spatial, temporal, or graph-based relationships.
+Rather than prescribing a specific physical model, this work adopts a general trellis-based prior as a unifying abstraction for correlated item statistics.
+This formulation subsumes the i.i.d.
+model as a special case and enables a single decoding framework to operate across diverse applications where correlation information is available or can be estimated.
+The use of the LVA is motivated by the observation that, under the Markovian prior model, the population sequence admits a natural trellis representation, and that approximate MAP decoding over this trellis can be performed efficiently by enumerating a small list of the most likely trajectories rather than a single path.
+Using LVA, 2SDGT leverages those statistics to estimate the defective set efficiently, even in a regime below the maximum likelihood (ML) upper bound.
 Furthermore, we show how the algorithm’s parameters can be tuned to achieve a maximum probability of success without exceeding the limitation of the available computational capacity.
-We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the MSGT computational complexity efficiency.
+We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the 2SDGT computational complexity efficiency.
 We provide an analytical sufficiency bound for the number of pooled tests needed by any maximum a-posteriori (MAP) decoder.
-This bound holds for any correlation between infected items.
+This bound holds for any dependence between infected items.
 It also applies to multiple access communication systems, where $K$ out of $N$ users transmit information simultaneously.
@@
 In contrast, existing analyses with prior statistics assume all $N$ users transmit at the same time [58, 69, 70, 12, 11, 51, 52].
-Our numerical results demonstrate that in practical regimes for COVID-19 [43, 44] and sparse signal recovery in signal processing [9, 10, 11, 13], the low computational complexity MSGT algorithm proposed herein can reduce in the scenarios tested the number of pool tests by at least $25\%$.
+Our numerical results demonstrate that in practical regimes for COVID-19 [44, 9] and sparse signal recovery in signal processing [32, 63, 1, 26], the low computational complexity 2SDGT algorithm proposed herein can reduce in the scenarios tested the number of pooled tests by at least $25\%$.
 The rest of this paper is organized as follows.
 Section II formally describes the GT model with correlated prior statistics.
-Section III presents the MSGT algorithm and the analytical results, and Section IV presents a sufficiency bound for MAP decoder.
+Section III presents the 2SDGT algorithm and the analytical results, and Section IV presents a sufficiency bound for MAP decoder.
 Section V details the simulation evaluation.
@@
 II Problem Formulation
-Given a set of individuals $\mathcal{N}$, the objective in GT is to discover a small subset $\mathcal{K}$ of unknown defective items using the minimum number of measurements $T$.
-Let $N=|\mathcal{N}|$, $K=|\mathcal{K}|$ denote the total number of items and the number of defective items, respectively, where $K=\mathcal{O}(N^{\theta})$ for some $\theta\in\left[0,1\right)$.
-The binary vector $\mathbf{U}\in\left\{0,1\right\}^{N}$ represents the population, such that $U_{i}=1$ indicates that the $i$-th item is defective.
-We assume that the set of the individuals is sparse, such that $\theta\leq 1/3$[17, 19, 45, 14], and that each item can be in one of 2 states: defective and not-defective.
-For each item, there is some prior probability that it is defective, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}\in\left[0,1\right]^{N}$, $\pi_{i}=P\left(U_{i}=1\right)$.
-The correlation between the state of the $i$-th item and the states of the previous $\tau$ items is represented by $\mathbf{\Phi}_{i}\in\left[0,1\right]^{2^{\tau}\times 2}$.
-$\tau=1,2,\dots$ represents the number of memory steps of the underlying process considered in the recovery stage.
-When there is no memory in the process, $\tau=0$ and $\{\pi_{i}\}_{i=1}^{N}$ hold all the prior information.
-$\mathbf{\Phi}_{i}[l,k]=P\left(U_{i}=s_{k}\lvert\left(U_{i-\tau},...,U_{i-1}\right)=s_{l}\right)$, where $s_{l}$ is the binary representation of $l$ by a $\tau$-length binary vector and $s_{k}\in\{0,1\}$ (for example, see Fig.
-7(b)).
+Given a set of individuals $\mathcal{N}$, the objective in GT is to discover a small random subset $\mathcal{K}$ of unknown defective items using the minimum number of measurements $T$.
+Let $N=|\mathcal{N}|$ denote the total number of items.
+Let $\mathcal{U}=\{0,1\}$ be the state alphabet, and let $\mathbf{U}=(U_{1},\ldots,U_{N})\in\mathcal{U}^{N}$ represent the random population vector, where $U_{i}=1$ indicates that the $i$-th item is defective.
+Under this model, the random defective set is $\mathcal{K}\triangleq\{i\in\mathcal{N}:U_{i}=1\}$ and the number of defective items is the random variable $K\triangleq|\mathcal{K}|=\sum_{i=1}^{N}U_{i}$.
+We denote by $\mu\triangleq\mathbb{E}[K]=\sum_{i=1}^{N}\Pr(U_{i}=1)$ the expected number of defective items.
+We model the prior over the sequence as a $\tau$-order Markov process.
+The joint law of the first $\tau$ variables is specified by an initial distribution $\pi(u_{1},\ldots,u_{\tau})\triangleq P(U_{1}=u_{1},\ldots,U_{\tau}=u_{\tau})$, for $(u_{1},\ldots,u_{\tau})\in\mathcal{U}^{\tau}$.
+This initial distribution generalizes the marginal parameters $\{\pi_{i}\}$ stated earlier: for $\tau=1$ it reduces to $\pi(u_{1})=P(U_{1}=u_{1})$ (equivalently, $\pi_{1}(u_{1})$), whereas for $\tau>1$ the model requires specifying the joint initialization of the first $\tau$ states.
+For $i\geq\tau+1$, the dependence on the previous $\tau$ states is specified by the transition kernels $\{\Phi_{i}\}_{i=\tau+1}^{N}$, defined by
+$\Phi_{i}\!\left[u_{i}\mid u_{i-\tau},\ldots,u_{i-1}\right]\triangleq P\!\left(U_{i}=u_{i}\,\big|\,U_{i-1}=u_{i-1},\ldots,U_{i-\tau}=u_{i-\tau}\right).$
+Equivalently, the transition of the $i$-th item can be represented in matrix form by
+$\boldsymbol{\Phi}_{i}\in[0,1]^{2^{\tau}\times 2},\quad\text{with transition probability}\quad\boldsymbol{\Phi}_{i}[\ell,k]=P\!\big(U_{i}=s_{k}\,\big|\,(U_{i-\tau},\ldots,U_{i-1})=s_{\ell}\big),$
+where $s_{\ell}\in\{0,1\}^{\tau}$ is the $\tau$-bit binary representation of $\ell\in\{0,\ldots,2^{\tau}-1\}$ and $s_{k}\in\{0,1\}$ (see Fig. 1). The distribution of the entire chain is then
+$P\left(U_{1}=u_{1},\ldots,U_{N}=u_{N}\right)=\pi(u_{1},\ldots,u_{\tau})\prod_{i=\tau+1}^{N}\Phi_{i}\!\left[u_{i}\mid u_{i-\tau},\ldots,u_{i-1}\right].$
+When there is no memory ($\tau=0$), the process reduces to independent (but not necessarily identical) marginals, and the prior information is fully captured by the marginal probabilities
+$P(U_{1}=u_{1},\ldots,U_{N}=u_{N})=\prod_{i=1}^{N}P(U_{i}=u_{i}).$
+Figure 1: Correlation (i.e., dependence of objects) matrix $\mathbf{\Phi}_{i}$ for $\tau=2$.
+Instead of conditioning on a fixed realization of $K$ [4], we introduce design sparsity bounds $K_{\min}$ and $K_{\max}$ that are used for test-matrix design and algorithmic thresholds, and are not assumed to equal the realized value of $K$.
+A convenient choice is $K_{\min}\triangleq\lfloor(1-\varepsilon)\mu\rfloor$ and $K_{\max}\triangleq\lceil(1+\varepsilon)\mu\rceil$, for some $\varepsilon\in(0,1)$.
+In subsequent sections, the recovery performance is analyzed by conditioning on the typical event $\color[rgb]{0,0,0}\{K_{\min}\leq K\leq K_{\max}\}$}, while accounting separately for the tail probability $\Pr(K<K_{\min})+\Pr(K>K_{\max})$ via concentration bounds on $K$.
+However, it is important to emphasize that the random variable $K$ is not restricted a priori and may take any value in $\{0,\ldots,N\}$.
+We focus on sparse regimes with $\mu=\mathcal{O}(N^{\theta})$ for some $\theta\in[0,1)$, and in particular $\theta\leq 1/3$ [8, 3, 54, 4].
 For the non-adaptive GT, the testing matrix $\mathbf{X}\in\left\{0,1\right\}^{T\times N}$ is defined such that each row corresponds to a single pool test, and each column corresponds to a single item.
@@
 Then, the outcome of the $i$-th pool test is given by
-$Y_{i}=\bigvee_{j\in\mathcal{K}}X_{i}(j)=\bigvee_{j=1}^{N}X_{i}(j)U_{j}$
-where $\bigvee$ is the Boolean OR function. All vectors are column vectors and the $\log$ operator denotes $\log_{2}$, unless stated otherwise. These definitions are illustrated in Fig. 1.
+$Y_{i}=\bigvee_{j\in\mathcal{K}}X_{i}(j)=\bigvee_{j=1}^{N}X_{i}(j)U_{j},$
+where $\bigvee$ is the Boolean OR function.
 Given $\mathbf{X}$ and the outcome vector $Y$, the recovery success criterion in GT can be measured using various metrics [4].
 The main metrics we will use herein are exact recovery and partial recovery.
+We note that the trellis (Markov) prior adopted herein should be interpreted as a modeling abstraction rather than as an assumption of a fixed physical ordering of items.
+The index $i$ does not necessarily correspond to spatial or temporal adjacency; instead, it represents an ordering induced by latent structure such as clustering, spatial proximity, temporal evolution, graph embeddings, or other side information available to the designer.
+When the items are exchangeable and no correlation information is available, the proposed model naturally reduces to the independent prior as a special case ($\tau=0$).
+When correlations are present, an appropriate permutation of item indices allows the dependency structure to be represented using a low-order Markov model.
+Under such correlated priors, the test matrix need not be permutation-invariant, as exploiting prior information inherently breaks symmetry across items.
+This behavior is not a limitation of the method but rather a consequence of leveraging structured side information to improve recovery performance.
+The proposed framework does not assume that such priors are always observable in practice; instead, it provides a general recovery architecture capable of exploiting correlated priors whenever they are available or can be inferred from data.
+We consider the prior probability of a defective set $\mathcal{K}$ through the population vector $\mathbf{U}=\mathbf{1}_{\mathcal{K}}$: $P(\mathcal{K})\ \triangleq\ P(\mathbf{U}=\mathbf{1}_{\mathcal{K}})$, where $P(\mathbf{U})$ is induced by the $\tau$-order Markov model described above in this section.
+Here $P(\mathcal{K})$ denotes the probability of the defective set induced by the prior distribution on the population vector $\mathbf{U}$, without conditioning on a fixed sparsity level.
+Note that this probability is defined over all subsets $\mathcal{K}\subseteq\mathcal{N}$ and does not restrict the cardinality $|\mathcal{K}|$.
 In terms of exact recovery, the goal is to detect the precise subset of defective items $\mathcal{K}$.
-Therefore, given the estimated defective set $\hat{\mathcal{K}}=\hat{\mathcal{K}}(N,K,\mathbf{X},Y)$, we define the average error probability by [^1] For simplicity of notation, $P_{s}$ and $P_{e}=1-P_{s}$ refer to success and error probabilities in the exact recovery analysis.
-$P_{e}^{exact}\triangleq\frac{1}{\binom{N}{K}}\sum_{\mathcal{K}:\left|\mathcal{K}\right|=K}P\left(\hat{\mathcal{K}}\neq\mathcal{K}\right).$
-In partial recovery, we allow a false positive (i.e., $|\hat{\mathcal{K}}\setminus\mathcal{K}|$) and false negative (i.e., $|\mathcal{K}\setminus\hat{\mathcal{K}}|$) detection rate.
-Thus, the average partial success rate is given by:
-$P_{s}^{partial}\triangleq\frac{1}{\binom{N}{K}}\sum_{\mathcal{K}:\left|\mathcal{K}\right|\neq K}\frac{\left|\hat{\mathcal{K}}\cap\mathcal{K}\right|}{K}.$
-In conclusion, the goal of non-adaptive GT with correlated prior data is to identify the $K$ defective items out of the $N$ items, where there is correlation between different infected items.
-To achieve this, one needs to design a $T\times N$ testing matrix and an efficient and practical recovery algorithm that can exploit correlated priors.
-The algorithm can identify the subset of infected items by observing $Y^{T}$ with high probability and with feasible computational complexity.
-Given the knowledge of the correlated prior data items and the available computational resources, one can design the testing matrix and a recovery algorithm to maximize the success probability.
+Therefore, given the estimated defective set $\hat{\mathcal{K}}=\hat{\mathcal{K}}\!\left(N,K_{\min},K_{\max},\mathbf{X},Y,\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}\right)$, we define the average exact error probability by [^1] For simplicity of notation, $P_{s}$ and $P_{e}=1-P_{s}$ refer to success and error probabilities in the exact recovery analysis.
+$P_{e}^{\mathrm{exact}}\triangleq\sum_{\mathcal{K}\subseteq\mathcal{N}}P(\mathcal{K})\;\Pr\!\left(\hat{\mathcal{K}}\neq\mathcal{K}\,\big|\,\mathcal{K}\right)=\Pr\!\left(\hat{\mathcal{K}}\neq\mathcal{K}\right).$
+Moreover, for any design sparsity bounds $(K_{\min},K_{\max})$,
+$P_{e}^{\mathrm{exact}}\leq\Pr\!\left(\hat{\mathcal{K}}\neq\mathcal{K}\mid{\color[rgb]{0,0,0}K_{\min}\leq K\leq K_{\max}}\right){\color[rgb]{0,0,0}+\Pr(K<K_{\min})+\Pr(K>K_{\max}).}$ (1)
+Here $(K_{\min},K_{\max})$ are design parameters rather than the realized sparsity $K$.
+[…]
+$P_{s}^{\mathrm{partial}}\triangleq\sum_{\mathcal{K}\subseteq\mathcal{N}}P(\mathcal{K})\;\mathbb{E}\left[\left.{|\hat{\mathcal{K}}\cap\mathcal{K}|}/{|\mathcal{K}|}\,\right|\mathcal{K}\right].$
+To conclude, for a sparse subset of infected items with random size $K$ (with $\mathbb{E}[K]=\mu$) out of $N$, the goal in non-adaptive GT with correlated prior data items is to design a $T\times N$ testing matrix and an efficient and practical recovery algorithm that can exploit correlated priors, such that by observing $Y^{T}$ we can identify the subset of infected items with high probability and with feasible computational complexity.
+Thus, given the knowledge of the correlated prior data items and the available computational resources, the test designer could design the testing matrix and a recovery algorithm to maximize the success probability.
 III Main Results
-In this section, we introduce the efficient multi-stage recovery algorithm for any statistical prior represented in a trellis diagram [33], detailed in Algorithm 1.
+In this section, we introduce the efficient two-stage [^2] Although we refer to the decoding procedure as consisting of two stages, the testing process itself is fully non-adaptive: the considered testing matrix is fixed in advance (see Section III-A), and no additional tests are designed based on intermediate decoding results.
+The term “stage” is used here only to distinguish between successive computational steps of the proposed decoder herein.
+recovery algorithm for any statistical prior represented in a trellis diagram [46], detailed in Algorithm 1.
 In the first stage, standard low-complexity algorithms [3] reduce the search space independently of prior correlations.
@@
 Additionally, we derive a bound to ensure the low complexity of the entire algorithm.
+Under the correlated prior (See Section II), the number of defectives $K=\sum_{i=1}^{N}U_{i}$ is a random variable.
+The algorithm is not assumed to know the realized $K$.
+Whenever a step requires a cardinality (filtering the LVA output and enumerating candidate supports), we use $(K_{\min},K_{\max})$ as the design bounds.
+Performance under random $K$ is handled by conditioning on the typical event $\{K_{\min}\leq K\leq K_{\max}\}$ and adding the tail probability $\Pr(K<K_{\min})+\Pr(K>K_{\max})$, as defined in (1) and elaborated in Section III-C.
+A key design parameter of the proposed 2SDGT algorithm presented in this section is the constant $\gamma\geq 1$, which controls the maximal candidate support size admitted from the LVA stage to the final MAP decoding stage.
+Specifically, only trajectories whose estimated number of defective items lies within the interval $K_{\min},\lfloor\gamma K_{\max}\rfloor]$ are retained for candidate generation.
+The parameter $\gamma$ therefore directly governs the tradeoff between computational complexity and recovery performance: larger values of $\gamma$ increase robustness to model mismatch and estimation errors at the cost of increased candidate enumeration, while smaller values enforce stricter complexity constraints.
 Section III-A describes the proposed algorithm.
@@
 III-A1 Testing Matrix and Pooling
-The proposed multi-stage recovery algorithm is intended to work with any non-adaptive testing matrix, e.g., as given in [46].
-To simplify the technical aspects and focus on the key methods, the testing matrix is generated randomly under a fixed optimal approximation with Bernoulli distribution of $p=\ln(2)/K$ [47], using classical GT methods.
+The proposed two-stage recovery algorithm is intended to work with any non-adaptive testing matrix, e.g., as given in [37].
+To simplify the technical aspects and focus on the key methods, the testing matrix is generated randomly under design sparsity bounds $(K_{\min},K_{\max})$ (as defined in Section II) with Bernoulli distribution of $p=\ln(2)/K_{\max}$ [2], using classical GT methods.
 The pooling and its outcome are given by the process elaborated in Section II and illustrated in Fig.
-1.
-Figure 1: For an unknown population $\mathbf{U}\in\left\{0,1\right\}^{9}$ with $K=2$, a random testing matrix is sampled and the test result $Y$ is calculated.
-Algorithm 1 Multi-Stage Recovery Algorithm
-Input: $\mathbf{X},Y,\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N},K,L,\tau,\gamma$
+2.
+We emphasize that the testing matrix is intentionally chosen to be a standard Bernoulli design, so that the impact of exploiting correlated prior information is isolated to the decoding stage [^3] An interesting direction for future work is the design of non-adaptive testing matrices that explicitly exploit non-uniform or correlated priors, for example via item-dependent testing probabilities..
+Figure 2: For an unknown population $\mathbf{U}\in\left\{0,1\right\}^{9}$ a realized defective set of size $K=2$ and design interval $K_{\min},K_{\max}]=[2,2]$, a random testing matrix is sampled and the test result $Y$ is calculated.
+Algorithm 1 Two-Stage Recovery Algorithm
+Input: $\mathbf{X},Y,{\color[rgb]{0,0,0}\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}},{\color[rgb]{0,0,0}K_{\min},K_{\max}},L,\tau,\gamma$
 Output: $\hat{\mathcal{K}}$
 1:Stage 1: Reduction of space search
-2:$\mathcal{P}^{(DND)}\leftarrow\mathbf{DND}(\mathbf{X},Y)$
-3:$\mathcal{P}^{(DD)}\leftarrow\mathbf{DD}\left(\mathbf{X},Y,\mathcal{P}^{(DND)}\right)$
-4:$\mathcal{P}^{(S_{1,2})}\leftarrow\mathcal{N}\setminus\left(\mathcal{P}^{(DND)}\cup\mathcal{P}^{(DD)}\right)$
+[…]
+[…]
+[…]
 5:Stage 2: Recovery exploiting prior info
-6:$\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}\leftarrow\mathbf{{updatePriors}}\left(\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N},\mathcal{P}^{(S_{1,2})},\mathcal{P}^{(DD)}\right)$
-7:$\mathbf{Z}\leftarrow\mathbf{LVA}\left(L,\tau,\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}\right)$
-8:$\mathbf{C}\leftarrow\{\}$
+[…]
+[…]
+[…]
 9:for $l\leftarrow 1$ to $L$ do
-10:    if $K\leq\sum_{i}\left(\mathbf{Z}_{l,i}\right)\leq\gamma K$ then
+10:  if ${\color[rgb]{0,0,0}K_{\min}\leq}\sum_{i}\left(\mathbf{Z}_{l,i}\right)\leq{\color[rgb]{0,0,0}\lfloor}\gamma{\color[rgb]{0,0,0}K_{\max}}{\color[rgb]{0,0,0}\rfloor}$ then
 11:    $\mathbf{V}^{(l)}\leftarrow\{i\mid\mathbf{Z}_{l,i}=1\}$
-12:         $\mathbf{C}\leftarrow\mathbf{C}\cup\mathbf{getAllCombinations}\left(\mathbf{V}^{(l)},K\right)$
+12:    $\mathbf{C}\leftarrow\mathbf{C}\cup\mathbf{getAllCombinations}\left(\mathbf{V}^{(l)},{\color[rgb]{0,0,0}K_{\min}},{\color[rgb]{0,0,0}K_{\max}}\right)$
 13:  end if
 14:end for
-15:$\hat{\mathcal{K}}\leftarrow\mathbf{MAP}\left(\mathbf{X},Y,\mathbf{C},\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}\right)$
+[…]
 III-A2 Recovery Process
@@
 In the first step of this stage, we use the Definitely Not Defective (DND) algorithm [38, 16, 15, 3] (line 2, Fig.
-2.(a)).
+3.(a), and Algorithm 2).
 DND compares the columns of the testing matrix, $\mathbf{X}$, with the outcome vector, $Y^{T}$.
@@
 In the second step of Stage 1, we use the Definitely Defectives (DD) algorithm [3] (line 3, Fig.
-2.(b)), which goes over the testing matrix and the test result, looking for positive pool tests that include only one possibly defective item.
+3.(b), and Algorithm 3), which goes over the testing matrix and the test result, looking for positive pooled tests that include only one possibly defective item.
 DD denotes those items as definitely defective items and outputs them as the set $\mathcal{P}^{(DD)}$.
-Let $\mathcal{P}^{(S_{1,1})}=\mathcal{N}\setminus\mathcal{P}^{(DND)}$ and $\mathcal{P}^{(S_{1,2})}=\mathcal{N}\setminus\left(\mathcal{P}^{(DND)}\cup\mathcal{P}^{(DD)}\right)$ denote the set of items that their status is still unclear after the first step and the second step, respectively.
+Let $\mathcal{P}^{(S_{1,1})}=\mathcal{N}\setminus\mathcal{P}^{(DND)}$ and $\mathcal{P}^{(S_{1,2})}=\mathcal{N}\setminus\left(\mathcal{P}^{(DND)}\cup\mathcal{P}^{(DD)}\right)$ denote the set of items that their status is still unclear after the first step and the second step, respectively, where we use the notation $S_{i,j}$ with the first index $i$ indicating the stage and the second index $j$ indicating the step within that stage.
 $\mathcal{P}^{(S_{1,2})}$ holds a new space search, and $\mathcal{P}^{(DD)}$ holds the already known defectives.
 This knowledge is acquired without utilizing any prior data, which we reserve for the second stage.
-Figure 2: First stage of MSGT.
+Figure 3: First stage of 2SDGT.
 (a) The first step of Stage 1, the DND algorithm, reveals 5 DND items in $\mathbf{U}$, forming $\mathcal{P}^{(DND)}$.
 Since items participating in negative tests must be non-defective, we mark all the participants in the two negative test results as non-defective.
-(b) The second step of Stage 1, the DD algorithm, outputs $\mathcal{P}^{(DD)}$ that includes a single DD item, based on the first test result, as it is the only possibly defective item participating in this test.
+(b) The second step of Stage 1, the DD algorithm, outputs $\mathcal{P}^{(DD)}$ that includes a single $D$ item, based on the first test result, as it is the only possibly defective item participating in this test.
 The two other positive test results do not contribute to our knowledge here because there is more than one possibly defective item participating in them.
-In the first step of Stage 2, we translate the data we obtained in DND and DD, $\mathcal{P}^{(S_{1,2})}$ and $\mathcal{P}^{(DD)}$, into the state space in terms of transition probabilities, $\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$, and initial probabilities, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}$, so we can employ all the gathered information in the next steps (line 6, Fig.
-3.(a)).
-In the state space, the population sequence, $\mathbf{U}$, is parallel to time steps considered traditionally in Viterbi Algorithm (VA) for communications problems [32], and there are two possible states per item, the first indicates “non-defective” and the second indicates “defective”.
+In the first step of Stage 2, we translate the data we obtained in DND and DD, $\mathcal{P}^{(S_{1,2})}$ and $\mathcal{P}^{(DD)}$, into the state space in terms of transition probabilities, $\color[rgb]{0,0,0}\mathbf{\pi}$}, and initial probabilities, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}$, so we can employ all the gathered information in the next steps (line 6, Fig.
+4.(a), and Algorithm 4).
+In the state space, the population sequence, $\mathbf{U}$, considered herein, is replaced with time steps considered traditionally in VA for communications problems [56], in which we have in GT two possible states per item, the first indicates “non-defective” and the second indicates “defective”.
+Remark 1.
+Under the Markovian prior model described in Section II, the population vector $\mathbf{U}=(U_{1},\ldots,U_{N})$ admits a natural representation as a path in a trellis, where each state corresponds to the defectivity status of an item (or a $\tau$-length state for $\tau$-order memory), and the transition probabilities are given by the prior parameters $(\pi,\{\Phi_{i}\})$.
+In this representation, exact MAP decoding corresponds to finding the most likely path through the trellis given the observed test outcomes.
+However, exact MAP decoding over this trellis requires evaluating an exponential number of paths and is therefore computationally infeasible for large $N$.
+The LVA provides an efficient approximation by returning a list of the $L$ most likely trellis trajectories, rather than committing to a single path.
+This is particularly well suited to the proposed framework, since after the Stage 1 reduction (DND and DD), multiple high-probability trajectories may remain consistent with the observations.
+By producing a small set of candidate trajectories, LVA enables systematic exploitation of the correlated prior structure while maintaining feasible computational complexity.
 In the second step, the suggested LVA for GT goes over the sequence of items and outputs $\mathbf{Z}\in\{0,1\}^{L\times N}$, a list of the $L$ most likely trajectories in the state space (line 7, Fig.
-3.(b)) according to MAP decision based on the given prior information.
+4.(b), and Algorithm 5) according to MAP decision based on the given prior information.
 Each trajectory is a sequence of $N$ states, representing items classified as either defective or non-defective.
@@
 In practice, the $L$ estimations may include any number of defective items and require further processing.
+The list size parameter $L$ controls the tradeoff between recovery performance and computational complexity.
+Larger values of $L$ increase the probability that the true defective pattern is contained in the list of candidate trajectories, at the cost of increased computation in subsequent candidate enumeration and MAP selection.
+This tunable tradeoff is a key advantage of using LVA in the proposed decoding architecture.
 In the third step, we extract candidates for the defective set out of the $L$ estimated sequences $\mathbf{Z}$ (lines 9 to 14).
 For some $l\in\{1,\ldots,L\}$, let $\mathbf{V}^{(l)}$ denote the set of items estimated as defective in $\mathbf{Z}_{l}$, the $l$-th row of $\mathbf{Z}$ (line 11).
-We ignore sequences that contain less than $K$ defective items or more than $\gamma K$ defective items, for some $\gamma\geq 1$, and consider only $\mathbf{Z}_{l}$ in which $K\leq\left|\mathbf{V}^{(l)}\right|<\gamma K$ as valid sequences.
-For each one of the valid sequences, we refer to all the combinations of size $K$ in $\mathbf{V}^{(l)}$ as candidates for $\hat{\mathcal{K}}$, and add them to the candidates list $\mathbf{C}$ (line 12).
-At this point, we have in $\mathbf{C}$ a list of candidates to be our final estimation $\hat{\mathcal{K}}$, and we can calculate the probability of each one of them using $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$.
-Then, in the fourth step, the estimated defective set, $\hat{\mathcal{K}}$, is finally chosen using MAP estimator out of the $\mathbf{C}$ (line 15), i.e., $\hat{\mathcal{K}}=\arg\max_{c\in\mathbf{C}}{P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)}$.
-If there are no valid sequences in $\mathbf{Z}$, we consider trajectories with fewer than $K$ detections for partial recovery.
-We select the trajectory with the most detections and randomly complete it to form a set of size $K$ for our final estimation $\hat{\mathcal{K}}$.
-Figure 3: Stage 2 of MSGT.
+We ignore sequences that contain fewer than $K_{\min}$ defective items or more than ${\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$ defective items, for some $\gamma\geq 1$, and consider only $\mathbf{Z}_{l}$ in which ${\color[rgb]{0,0,0}K_{\min}\leq}\left|\mathbf{V}^{(l)}\right|<{\color[rgb]{0,0,0}\lfloor}\gamma{\color[rgb]{0,0,0}K_{\max}{\color[rgb]{0,0,0}\rfloor}}$ as valid sequences.
+For each one of the valid sequences, we refer to all the combinations of size $k$ in $\mathbf{V}^{(l)}$, for every $k\in\{K_{\min},\ldots,\min(K_{\max},|\mathbf{V}^{(l)}|)\}$, as candidates for $\hat{\mathcal{K}}$ and add them to the candidates list $\mathbf{C}$ (line 12).
+In particular, for a given set of indices $\mathbf{V}^{(l)}\subseteq\mathcal{N}$ and integers $1\leq K_{\min}\leq K_{\max}$, we define $\mathbf{getAllCombinations}(\mathbf{V}^{(l)},K_{\min},K_{\max})\;\triangleq\;\bigcup_{k=K_{\min}}^{\min(K_{\max},|\mathbf{V}^{(l)}|)}\left\{\mathcal{S}\subseteq\mathbf{V}^{(l)}\;:\;|\mathcal{S}|=k\right\}$.
+Thus, $\mathbf{getAllCombinations}$ $(\mathbf{V}^{(l)},K_{\min},K_{\max})$ returns the collection of all subsets of $\mathbf{V}^{(l)}$ whose cardinality lies in the range $K_{\min}$ and $K_{\max}$.
+The operator $\mathbf{getAllCombinations}(\cdot)$ as detailed in Algorithm 6 is used here to generate candidate defective sets without assuming knowledge of the exact sparsity level.
+At this point, we have in $\mathbf{C}$ a list of sets of size $k\in\{K_{\min},\ldots,K_{\max}\}$, that are candidates to be our final estimation $\hat{\mathcal{K}}$, and we can calculate the probability of each one of them using $\color[rgb]{0,0,0}\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}$} (see Algorithm 4).
+Then, in the fourth step, the estimated defective set, $\hat{\mathcal{K}}$, is finally chosen using MAP estimator out of the $\mathbf{C}$ (line 15), i.e., $\hat{\mathcal{K}}=\arg\max_{c\in\mathbf{C}}{P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)}$ (see Algorithm 7).
+If there are no valid sequences in $\mathbf{Z}$, we consider trajectories with fewer than $K_{\min}$ detections for partial recovery.
+We select the trajectory with the most detections and randomly complete it to form a set of size $K_{\min}$ for our final estimation $\hat{\mathcal{K}}$.
+Figure 4: Stage 2 of 2SDGT.
 (a) All the possible transitions in the state space that we consider in the LVA step, following the insights obtained in Stage 1.
@@
 Given $K=2$, the black trajectory corresponds to a valid population vector $\mathbf{U}$ with 2 defective items, while the gray trajectory indicates an invalid population with 3 defective items instead.
-Consequently, in the subsequent step, MSGT will extract two optional defective sets: $\left\{U_{6},U_{8}\right\}$ and $\left\{U_{7},U_{8}\right\}$, and will finally choose the most likely one using MAP estimator.
+Consequently, in the subsequent step, 2SDGT will extract two optional defective sets: $\left\{U_{6},U_{8}\right\}$ and $\left\{U_{7},U_{8}\right\}$, and will finally choose the most likely one using MAP estimator.
 (c) Comparison of Stage 2 to ML.
@@
 With $T=6$, based on the third and sixth test results, there is only one set of size $K=2$ that matches the outcome $Y$, resulting in successful decoding with the ML decoder.
-As shown above, MSGT’s Stage 2 can successfully decode $\mathbf{U}$ with just $T=5$, as using the LVA step it narrows down to only 2 possible trajectories, and then the final estimation is selected based on the given prior information and the insights gained in Stage 1.
+As shown above, 2SDGT’s Stage 2 can successfully decode $\mathbf{U}$ with just $T=5$, as using the LVA step it narrows down to only 2 possible trajectories, and then the final estimation is selected based on the given prior information and the insights gained in Stage 1.
 III-B Analytical Results
-Here we provide analytical results related to the proposed MSGT.
-Let $P_{e,a}^{(DND)}$ and $P_{e,u}^{(DND)}$ denote the error probability of DND in the average case and its deviation from the average (“worst-case”) that still allows the success of MSGT on average in feasible computational complexity, respectively.
-The two following theorems give the upper bound on the expected number of possible defective items and a “worst-case” upper bound after the first step of Stage 1.
-Theorem 1 (Cohen et al. [4]).
-Consider a group test with a Bernoulli testing matrix with $p=\ln{2}/K$, and $T$ tests as $K\rightarrow\infty$.
-Let $P_{e,a}^{(DND)}\triangleq N^{-\alpha\left(1-\ln{2}/K\right)/2}$ for $\alpha\triangleq T/K\log_{2}{N}$.
+In this section, we provide analytical results related to the proposed 2SDGT. To support our theoretical results, we demonstrate the derived bounds and compare them to simulation results in Fig. 5.
+In the correlated-prior model, the number of defectives $K=\sum_{i=1}^{N}U_{i}$ is random but is assumed to satisfy the known bound $K_{\min}\leq K\leq K_{\max}$ (see Section II).
+Accordingly, throughout this section we analyze Stage 1 under a Bernoulli design with parameter $p=\ln(2)/K_{\max}$ and $T$ tests, where $K_{\max}$ is the design upper bound.
+Theorem 1 bounds the expected number of items whose status remains unknown after the DND step (equivalently, it bounds the expected number of non-defectives that survive as “possibly defective”), as a function of the realized $K=k$, and we denote the corresponding average contribution by $P_{e,a}^{(DND)}$.
+Complementarily, Theorem 2 upper bounds the probability that this number exceeds its expectation by a margin, and we denote this upper-tail deviation term by $P_{e,u}^{(DND)}$.
+These bounds control the size of the remaining search space passed to Stage 2, and in particular ensure that the Stage 2 list-size budget (parameterized by $\gamma$ and the design upper bound $K_{\max}$) is met.
+Theorem 1 ([60]).
+Consider a group test with a Bernoulli testing matrix with $p=\ln{2}/K_{\max}$ and $T$ tests, and suppose the (random) number of defectives $K$ satisfies $K_{\min}\leq K\leq K_{\max}$.
+Fix any realization $K=k$ and let $P_{e,a}^{(DND)}\triangleq N^{-\alpha\left(1-\ln{2}/K_{\max}\right)/2}$ for $\alpha\triangleq T/(K_{\max}\log_{2}N)$.
 The expected number of possibly defective items is bounded by
-$\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]\leq K+\left(N-K\right)P_{e,a}^{(DND)}.$
+$\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\,\big|\,K={\color[rgb]{0,0,0}k}\right]\leq{\color[rgb]{0,0,0}k+\left(N-k\right)}P_{e,a}^{(DND)}.$
+The proof of Theorem 1 follows directly from the definition of the DND decoding rule and the Bernoulli test design [22, 60].
+A non-defective item is eliminated unless it appears only in tests whose outcomes are positive; equivalently, it survives the DND step if and only if all tests in which it participates are occluded by at least one defective item.
+Since each item participates in tests independently with probability $p$, the survival event can be characterized by conditioning on the number of tests involving the item and on whether those tests include at least one defective.
+In particular, under a Bernoulli(p)(p) design with $k$ defectives, $(1-p)^{k}$ denotes the probability that a test is negative, and $p(1-p)^{k-1}$ denotes the probability that a test contains defective $i$ and no other defectives.
+This leads to the stated expression for the probability of a non-defective item surviving the DND step, which is then used in the subsequent expectation and concentration analysis.
 Proof.
-According to [51, Lemma 2]:
-$\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]=K+\left(N-K\right)P_{e}^{(DND)},$ (1)
-where $P_{e}^{(DND)}=\left(1-p\left(1-p\right)^{K}\right)^{T}$ is the probability of non-defective item to be hidden by another defective item during DND.
-By substituting $T=\alpha K\log_{2}{N}$, $p=\ln{2}/K$ we bound the error probability as follows:
-$P_{e}^{\left(DND\right)}=\left(1-\frac{\ln{2}}{K}\left(1-\frac{\ln{2}}{K}\right)^{K}\right)^{\alpha K\log_{2}{N}}$
-$=\left(1-\frac{\ln{2}}{K}\left(1-\frac{\ln{2}}{K}\right)^{\left(K-1\right)}\left(1-\frac{\ln{2}}{K}\right)\right)^{\alpha K\log_{2}{N}}$
-$\leq\left(1-\frac{\ln{2}}{K}e^{-\ln{2}}\left(1-\frac{\ln{2}}{K}\right)\right)^{\alpha K\log_{2}{N}}$
-$=e^{-\frac{\alpha}{2}\ln{2}\left(1-\frac{\ln{2}}{K}\right)\log_{2}{N}}$
-$=N^{-\frac{\alpha}{2}\left(1-\frac{\ln{2}}{K}\right)}$
-Finally, we denote this bound on the average error probability in DND by $P_{e,a}^{(DND)}\triangleq N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)}$.
-Hence, substituting $P_{e,a}^{(DND)}$ in (1), we have:
-$\mathbb{E}\left[\left|\mathcal{P}^{(DND)}\right|\right]\leq K+\left(N-K\right)P_{e,a}^{(DND)}.$
-This completed the theorem proof. ∎
+According to [10, Lemma 2], conditioning on $\{K={\color[rgb]{0,0,0}k}\}$ yields
+$\mathbb{E}\!\left[\left|\mathcal{P}^{(S_{1,1})}\right|\,\big|\,K={\color[rgb]{0,0,0}k}\right]={\color[rgb]{0,0,0}k+\left(N-k\right)}P_{e}^{(\mathrm{DND})},\quad\text{where}\quad P_{e}^{(\mathrm{DND})}=\bigl(1-p(1-p)^{{\color[rgb]{0,0,0}k}}\bigr)^{T}.$ (2)
+By substituting the design choice $p=\ln 2/K_{\max}$, we bound $P_{e}^{(\mathrm{DND})}$ as follows
+$P_{e}^{(\mathrm{DND})}$ $=\left(1-p(1-p)^{{\color[rgb]{0,0,0}k}}\right)^{T}$
+$=\left(1-p(1-p)^{{\color[rgb]{0,0,0}k-1}}(1-p)\right)^{T}$
+$\leq\left(1-p\,e^{-\ln 2}\,(1-p)\right)^{T},$
+where the last step follows since $e^{-\ln 2}\leq(1-\ln 2/m)^{m-1}$ for any integer $m\geq 0$, such that, together with the bound ${\color[rgb]{0,0,0}k}\leq K_{\max}$ and with $p=\ln 2/K_{\max}$, we have
+$(1-p)^{{\color[rgb]{0,0,0}k-1}}\;\geq\;(1-p)^{K_{\max}-1}\;\geq\;e^{-\ln 2}\;=\;\frac{1}{2},$
+as the first inequality holds since 0<1p<10<1-p<1 and ${\color[rgb]{0,0,0}k-1}\leq K_{\max}-1$, and the second from the exponential bound.
+Now, applying $(1-z)^{T}\leq e^{-zT}$ with
+$z\triangleq p\cdot\frac{1}{2}\cdot(1-p)=\frac{\ln 2}{K_{\max}}\cdot\frac{1}{2}\cdot\left(1-\frac{\ln 2}{K_{\max}}\right),\quad\text{and}\quad T=\alpha K_{\max}\log_{2}N,$
+we obtain
+$P_{e}^{(\mathrm{DND})}\leq\exp\!\left(-z\cdot\alpha K_{\max}\log_{2}N\right)=N^{-\frac{\alpha}{2}\left(1-\frac{\ln 2}{K_{\max}}\right)}.$
+Finally, substituting $P_{e,a}^{(\mathrm{DND})}\triangleq N^{-\frac{\alpha}{2}\left(1-\frac{\ln 2}{K_{\max}}\right)}$ into (2) yields
+$\mathbb{E}\!\left[\left|\mathcal{P}^{(S_{1,1})}\right|\,\big|\,K={\color[rgb]{0,0,0}k}\right]\leq{\color[rgb]{0,0,0}k+(N-k)}\,P_{e,a}^{(\mathrm{DND})}.$
 Theorem 2.
-Consider a group test with a Bernoulli testing matrix and DND decoder. As $K\rightarrow\infty$, the worst-case error probability of DND is bounded by
-$P_{e,u}^{(DND)}$ $\leq P_{e,a}^{(DND)}+\left(1+\frac{\left(\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)\right)^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1},$
-for any $\gamma\geq 1$.
+Consider a group test with $T$ tests, a Bernoulli testing matrix with the design parameter $p=\ln(2)/K_{\max}$, and the DND decoder.
+Fix any realization $K=k\in[K_{\min},K_{\max}]$ and let $G$ denote the number of non-defective items marked as possibly defective by DND.
+Then, for any threshold $g>0$,
+$P_{e,u}^{(\mathrm{DND})}(g)\;\triangleq\;P\!\left(G-\mathbb{E}[G]>g\,\big|\,K={\color[rgb]{0,0,0}k}\right)\;\leq\;\frac{\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})}{\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})+g^{2}},$ (3)
+where
+$\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})=(N-{\color[rgb]{0,0,0}k})(\bar{r}_{1}-\bar{r}_{2})\;+\;(N-{\color[rgb]{0,0,0}k})^{2}(\bar{r}_{2}-\bar{r}_{1}^{\,2}),$ (4)
+with $\bar{r}_{1}=\big(1-pq_{0}\big)^{T}$ and $\bar{r}_{2}=\big(1-(2p-p^{2})q_{0}\big)^{T}$ for $q_{0}\triangleq(1-p)^{{\color[rgb]{0,0,0}k}}$.
 Proof.
 Let $G$ denote the number of non-defective items that were hidden and not detected in DND (i.e., the number of false positive items).
-We upper bound the probability of error in DND, by the sum of the average probability of error $P_{e,a}^{(DND)}$ and a concentration term as follows:
-$P_{e,u}^{(DND)}\triangleq P_{e,a}^{(DND)}+P\left(G-\mathbb{E}[G]>g\right),$
-for some $g>0$.
-Although the variation of LVA in Stage 2 applies a maximum likelihood decision, it’s performance in the proposed algorithm, may be worse than a brute force MAP’s performance, if the number of occlusions in the test matrix exceeds the threshold $\gamma K$.
-For this reason we are interested in $g=\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)$.
-Recall that the probability of each definitely defective item to be hidden is given by the probability of error in DND, $P_{e}^{(DND)}$, which depends only on the matrix design.
-Thus, in the case of Bernoulli encoder, the $N-K$ random variables which represent occlusions of non-defective items are independent.
-It follows that $G\sim Bin\left(N-K,P_{e}^{(DND)}\right)$, and we can apply the one-sided Chebyshev’s inequality [52] to bound the concentration term:
-$P\left(G-\mathbb{E}\left[G\right]>g\right)\leq\\ \left(1+\frac{g^{2}}{\left(N-K\right)P_{e}^{(DND)}\left(1-P_{e}^{(DND)}\right)}\right)^{-1}$
-For a sufficiently large $T$, $P_{e}^{(DND)}\rightarrow P_{e,a}^{(DND)}$, which concludes the proof. ∎
+We introduce the following upper-tail deviation probability for the number of false positives
+$P_{e,u}^{(\mathrm{DND})}\;\triangleq\;P\!\left(G-\mathbb{E}[G]>g\,\big|\,K={\color[rgb]{0,0,0}k}\right),\quad\text{for}\quad g>0.$
+Under the defective model given in Section II with $K_{\min}\leq K\leq K_{\max}$, Stage 2 enforces only an upper budget ${\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$ on the support size (Algorithm 1).
+Accordingly, we choose $g$ to control the event that the number of possibly defective items after DND exceeds this budget, e.g.,
+$g\;=\;{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}\;-\;\mathbb{E}\!\left[\left|\mathcal{P}^{(S_{1,1})}\right|\mid K={\color[rgb]{0,0,0}k}\right]\;=\;{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}\;-\;\Big({\color[rgb]{0,0,0}k+(N-k)}P_{e,a}^{(\mathrm{DND})}\Big),$
+whenever this quantity is positive. Applying the one-sided Chebyshev’s inequality [27] to bound the concentration term gives
+$P\!\big(G-\mathbb{E}[G]>g\,\big|\,K={\color[rgb]{0,0,0}k}\big)\;\leq\;\frac{\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})}{\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})+g^{2}}.$ (5)
+Let $M_{0}$ denote the number of negative tests, $q_{0}\triangleq(1-p)^{{\color[rgb]{0,0,0}k}}$, and $n^{\prime}\triangleq N-{\color[rgb]{0,0,0}k}$.
+Each test is negative independently with probability $q_{0}$, hence $M_{0}\sim Bin(T,q_{0})$.
+Conditioned on $M_{0}=m_{0}$, a fixed non-defective item appears in none of the $m_{0}$ negative tests with probability $(1-p)^{m_{0}}$.
+Consequently, $G\,\big|\,M_{0}=m_{0},\ K={\color[rgb]{0,0,0}k}\;\sim\;Bin\big(n^{\prime},\ (1-p)^{m_{0}}\big)$ [3].
+For $M\sim Bin(T,q)$ and any $a\in\mathbb{R}$,
+$\mathbb{E}\!\left[a^{M}\right]\;=\;\big((1-q)+q\,a\big)^{T}.$ (6)
+Using the law of total variance and $r\triangleq\left(1-p\right)^{M_{0}}$,
+$\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})$ $=\mathbb{E}\!\left[\mathrm{Var}(G\mid M_{0},K={\color[rgb]{0,0,0}k})\right]+\mathrm{Var}\!\left(\mathbb{E}[G\mid M_{0},K={\color[rgb]{0,0,0}k}]\right)$
+$=\mathbb{E}\!\left[n^{\prime}r(1-r)\right]+\mathrm{Var}(n^{\prime}r)$
+$=n^{\prime}\big(\mathbb{E}[r]-\mathbb{E}[r^{2}]\big)+n^{\prime 2}\big(\mathbb{E}[r^{2}]-(\mathbb{E}[r])^{2}\big).$ (7)
+Using (6), we obtain
+$\bar{r}_{1}=\mathbb{E}[r]=(1-pq_{0})^{T},\quad\text{and}\quad\bar{r}_{2}=\mathbb{E}[r^{2}]=\big(1-(2p-p^{2})q_{0}\big)^{T}.$
+Substituting these expressions into the variance decomposition (7) yields the variance formula stated in (4).
+Finally, combining (5) with (4) establishes the upper-tail bound in (3), completing the proof.
 The following theorem provides the expected defective items detected by DD in the second step of Stage 1.
 Theorem 3.
-Consider a group test with a Bernoulli testing matrix with $p=\ln{2}/K$, and $T$ tests as $K\rightarrow\infty$. The expected number of defective items that DD successfully detects is bounded by
-$\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]\geq K\left(1-N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}\right).$ (2)
+Consider noiseless nonadaptive group testing with $N$ items, a random number of defectives $K$ (conditioned on its realization), and $T$ tests, under a Bernoulli(p)(p) design with design parameter $p=\ln(2)/K_{\max}$.
+Fix any realization $k\in[K_{\min},K_{\max}]$ and consider performance conditioned on $\{K=k\}$.
+The number of defective items detected by the DD decoder is bounded by
+$\mathbb{E}\big[\,|\mathcal{P}^{(\mathrm{DD})}|\,\big|\,K={\color[rgb]{0,0,0}k}\big]\;\geq\;{\color[rgb]{0,0,0}k}\!\left\{a\,T(1-q_{0})\,(1-p)^{N-{\color[rgb]{0,0,0}k}}-\frac{a^{2}}{2}\,T^{2}\right\}^{+},$ (8)
+where $\{x\}^{+}\triangleq\max\{x,0\}$, all expectations are taken conditionally on the realized $K$, and $q_{0}\triangleq(1-p)^{{\color[rgb]{0,0,0}k}}$, $q_{1}\triangleq p(1-p)^{{\color[rgb]{0,0,0}k-1}}$, and $a\triangleq q_{1}/(1-q_{0})$.
+Proof:.
+The proof of Theorem 3 in given in Appendix B. ∎
+Remark 2.
+Under the Bernoulli design used in this work, the testing matrix is parameterized by a design sparsity range $K_{\min},K_{\max}]$, and the realized number of defectives $K$ is random and unknown to the decoder.
+Accordingly, the lower bound in (8), which is stated conditionally on a fixed realization of $K$, may evaluate to zero for some values of $K\in[K_{\min},K_{\max}]$ after applying the positive-part operator.
+As a result, the bound is not uniformly informative over the entire admissible sparsity range, but remains valid for all realizations of $K$.
+Since we always have the trivial bound $\mathbb{E}[\,|\mathcal{P}^{(\mathrm{DD})}|\,]\geq 0$, this behavior is consistent with correctness and does not indicate a failure of the analysis.
+Moreover, the proposed 2SDGT algorithm does not rely on the magnitude of $|\mathcal{P}^{(\mathrm{DD})}|$ for correctness or feasibility.
+In particular, Stage 2 applies a deterministic filtering rule that retains only candidate supports whose cardinality lies in the range $K_{\min}\leq|\mathcal{V}^{(l)}|\leq\lfloor\gamma K_{\max}\rfloor$, independently of the quantitative DD bound.
+For this reason, the complexity analysis of 2SDGT in Theorem 4 relies only on the elementary inequalities $0\leq|\mathcal{P}^{(\mathrm{DD})}|\leq|\mathcal{P}^{(S_{1,1})}|\leq N$, rather than on the quantitative lower bound (8).
+The proposition below states a prior-free counting bound relating the Stage 2 candidate-set budget and the best achievable uniformly averaged probability of exact support recovery at a fixed sparsity level $k$, where the averaging is uniform over all supports of cardinality $k$ (see Remark 3).
+It is used to motivate how the Stage 2 parameter $\gamma$ should scale with $T$ and $(N,K_{\max})$.
+Although the 2SDGT and MAP decoders explicitly depend on the prior $P(\mathcal{K})$, this proposition is a prior-free counting bound that applies to any decoding rule.
+In particular, the bound remains valid when MAP is used with an arbitrary correlated prior, since it constrains only the number of candidate supports consistent with a given outcome.
+Proposition 1.
+Fix any $k\in\{K_{\min},\ldots,K_{\max}\}$.
+Let $\eta_{k}\triangleq\left({1}/{\binom{N}{k}}\right)\sum_{\mathcal{K}:\,|\mathcal{K}|=k}P\!\left(\hat{\mathcal{K}}=\mathcal{K}\right)$ denote the uniformly averaged (over all $k$-subsets) probability of exact support recovery at sparsity level $k$.
+Consider any decoding procedure that, given $Y$, outputs a candidate item set $\mathcal{L}(Y)\subseteq[N]$ satisfying $|\mathcal{L}(Y)|\leq{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$.
+Then, achieving $\eta_{k}\geq\bar{\eta}_{k}$ requires
+$T\;\geq\;\log_{2}\bar{\eta}_{k}+\log_{2}\binom{N}{k}-\log_{2}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k}$. (9)
 Proof.
-The detection of the $i$-th defective item ($i\in\mathcal{K}$) in the DD algorithm fails under two conditions: either it does not participate in any test, or it participates but at least one other potentially defective item is also participating.
-Let $A_{i}$ represent the event where at least one potentially defective item, excluding item $i$, participates in some test.
-Consequently, $P\left(A_{i}\right)=1-\left(1-p\right)^{\left|\mathcal{P}^{(S_{1,1})}\right|-1}$.
-The probability of not identifying item $i$ as definitively defective in a given test $t\in\{1,...,T\}$ is given by $1-p+pP\left(A_{i}\right)$.
-So the probability of not detecting a defective item in all tests is given by
-$P_{e}^{(DD)}$ $=\left[1-p+p\left(1-\left(1-p\right)^{\left|\mathcal{P}^{(S_{1,1})}\right|-1}\right)\right]^{T}$
-$=\left[1-p\left(1-p\right)^{\left|\mathcal{P}^{(S_{1,1})}\right|-1}\right]^{T}.$
-By substituting (1) we have:
-$\mathbb{E}\left[P_{e}^{(DD)}\right]$
-$\leq\left(1-p\left(1-p\right)^{K+\left(N-K\right)P_{e,a}^{(DND)}-1}\right)^{T}$
-$\leq\left(1-p\left(1-p\right)^{K+P_{e,a}^{(DND)}N-1}\right)^{T}$
-=$\underset{(a)}{=}\left(1-\frac{\ln{2}}{K}\left(1-\frac{\ln{2}}{K}\right)^{K-1}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}\right)^{T}$
-$\underset{(b)}{\leq}\left(1-\frac{\ln{2}}{K}e^{-\ln{2}}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}\right)^{\alpha K\log_{2}{N}}$
-$\underset{(c)}{\leq}exp\left(-\alpha e^{-\ln{2}}\ln{2}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}\log_{2}{N}\right)$
-$=N^{-\frac{\alpha}{2}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}},$
-where in $(a)$ we substitute $p=\ln{2}/K$, in $(b)$ we substitute $T=\alpha K\log_{2}{N}$ and used the fact that $e^{-\ln 2}\leq\left(1-\ln{2}/n\right)^{n-1}$ for any integer $n>0$.
-$(c)$ follows since $e^{-x}\geq\left(1-x/n\right)^{n}$ for any integer $n>0$ and $x>0$.
-Subsequently, the probability of success satisfies
-$\mathbb{E}\left[P_{s}^{(DD)}\right]=1-\mathbb{E}\left[P_{e}^{(DD)}\right]\geq 1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}.$
-Thus, an upper bound on the expected number of defective items that the DD algorithm successfully detects can be established as follows:
-$\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]=\mathbb{E}\left[P_{s}^{(DD)}\right]K\geq\left(1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}\right)K.$
-This completed the theorem proof. ∎
-The proposition below provides a sufficient condition on $\gamma$, such that, MSGT achieves performance that outperforms ML.
-This can be assumed since LVA used for this step is an optimal MAP estimator [32].
-Proposition 1.
-Let $\eta$ denote the average success probability of ML decoder.
-For GT with correlated prior information and $\gamma\geq 1$, MSGT achieve an average success probability equal or greater than $\eta$, if
-$T\geq\log{\eta}+\log{\binom{N}{K}}-\log{\binom{N-K}{K(\gamma-1)}}.$ (3)
+Fix an arbitrary $k\in\{K_{\min},\ldots,K_{\max}\}$.
+We follow a standard counting argument in the spirit of list-decoding converses (cf.
+[54]), adapted here to decoders that output a bounded-size item set.
+For each outcome vector $Y\in\{0,1\}^{T}$, define
+$N(Y)\triangleq\sum_{\mathcal{K}:\,|\mathcal{K}|=k}\mathbb{1}\{\mathcal{K}\subseteq\mathcal{L}(Y)\},$
+namely, the number of $k$-supports fully contained in the candidate set $\mathcal{L}(Y)$.
+Since $|\mathcal{L}(Y)|$ is integer, we have $|\mathcal{L}(Y)|\leq\lfloor\gamma K_{\max}\rfloor$, and hence
+$N(Y)=\binom{|\mathcal{L}(Y)|}{k}\leq\binom{\lfloor\gamma K_{\max}\rfloor}{k}$.
+Moreover, for any true support $\mathcal{K}$ with $|\mathcal{K}|=k$, the event $\{\hat{\mathcal{K}}=\mathcal{K}\}$ necessarily implies that all defective items must survive Stage 2, and hence $\mathcal{K}\subseteq\mathcal{L}(Y)$.
+Averaging uniformly over all such supports and summing over all outcomes yields
+$\eta_{k}\leq\frac{1}{\binom{N}{k}}\sum_{Y}N(Y),$
+where the inequality follows since $\mathbb{1}\{\hat{\mathcal{K}}=\mathcal{K}\}\leq\mathbb{1}\{\mathcal{K}\subseteq\mathcal{L}(Y)\}$ and $\Pr(Y\mid\mathcal{K})\leq 1$.
+Using the bound on $N(Y)$ and the fact that there are $2^{T}$ possible outcomes, we obtain
+$\eta_{k}\leq\frac{2^{T}}{\binom{N}{k}}\binom{\lfloor\gamma K_{\max}\rfloor}{k}$. (10)
+Rearranging (10) completes the proof.
+Remark 3.
+The success probability $\eta_{k}$ in Proposition 1 is defined as the probability of exact support recovery, uniformly averaged over all supports of cardinality $k$.
+This metric is not the prior-dependent success probability of the MAP/LVA decoder under the correlated prior considered in this work.
+It is introduced solely to enable a prior-free counting argument.
+Importantly, this proposition does not characterize the performance of MAP decoding under the correlated prior.
+Rather, it provides a necessary combinatorial limitation that applies to any decoding rule whose Stage 2 output is restricted to a candidate item set of size at most ${\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$.
+Since this limitation is independent of the prior distribution over supports, it remains valid when MAP is employed with an arbitrary correlated prior.
+The bound is therefore used solely to motivate the scaling of $\gamma$ and $T$, and not as a direct performance guarantee.
+One of the key features of the proposed 2SDGT algorithm is its low and feasible complexity in practical regimes compared to ML or MAP-based GT decoders.
+Both ML and MAP involve exhaustive searches, resulting in a complexity of $\mathcal{O}\left(\binom{N}{K}KN\log_{2}N\right)$ operations [8].
+The theorem below and the subsequent Remark characterize the computational complexity of 2SDGT.
+Theorem 4.
+Consider a group test for a population of $N$ items, with an unknown number of defectives $K$ satisfying $K_{\min}\leq K\leq K_{\max}$, and a Bernoulli testing matrix designed using $K_{\max}$.
+The computational complexity of the 2SDGT algorithm is bounded by $\color[rgb]{0,0,0}\mathcal{O}\bigl(L\gamma^{K_{\max}}K_{\max}N\log_{2}N\bigr)$} operations.
 Proof.
-We follow the lines of the success probability analysis in [45, Theorem 1].
-Consider a list decoder that, given the test outcome $Y$, outputs $\mathcal{L}(Y)$, a list of items estimated as defective, such that $\lvert\mathcal{L}(Y)\lvert\geq K$.
-Let $N(Y)$ denote the number of possible defective sets for which, given $Y$, the list decoding is successful: $N(Y)\triangleq\sum_{\mathcal{K}}\mathbb{1}_{\{\mathcal{L}(Y)\cap\mathcal{K}=K\}}$.
-Assuming no prior information is given, $N(Y)$ is bounded by the counting bound
-$N(Y)\leq\binom{N-K}{L-K}.$ (4)
-The success probability of the decoder can be written as
-$P_{s}=\frac{1}{\binom{N}{K}}\sum_{Y}N(Y)=\frac{2^{T}}{\binom{N}{K}}N(Y).$ (5)
-By substituting (4) into (5), taking the logarithm and rearranging, we have that
-$T\geq\log{P_{s}}+\log{\binom{N}{K}}-\log{\binom{N-K}{L-K}}.$
-It is important to note that unlike [45, Theorem 1], we halt the success probability analysis here without deriving a converse, and instead focus on the connection between $T$ and $\gamma$.
-This analysis is valid for any list decoder, particularly for an ML decoder [17, 53], allowing us to set $P_{s}=\eta$.
-Since MSGT uses MAP, the non-uniform and correlated prior information allows us to achieve better recovery performance for the same $T$ and $L$ compared to ML [^2] In Section IV-C, we provide sufficiency bound for any MAP decoder with an arbitrary underlying correlation between infected items that outperform the ML sufficiency bound and illustrate it for practical scenarios in Section V..
-Moreover, since for any $\gamma\geq 1$, $\gamma K$ is the maximum list size of estimated defective items we allow in MSGT.
-By substituting $L=\gamma K$, we establish a lower bound for $\gamma$ that guarantees MSGT’s success probability remains at least equal to that of the ML estimator.
-∎
-One of the key features of the proposed MSGT algorithm is its low and feasible complexity in practical regimes compared to ML or MAP-based GT decoders.
-Both ML and MAP involve exhaustive searches, resulting in a complexity of $\mathcal{O}(\binom{N}{K}KN\log_{2}N)$ operations [17].
-The theorem below and the subsequent remark characterize the computational complexity of MSGT.
-Theorem 4.
-Consider a group test for a population of $N$ items, of which $K$ are defective and a Bernoulli testing matrix.
-The computational complexity of the MSGT algorithm is bounded by $\mathcal{O}\left(L\gamma^{K}KN\log_{2}N\right)$ operations.
-Proof.
-We begin by analyzing the complexity of each step of the proposed MSGT solution given in Algorithm 1, and finally, we sum everything up to determine the total complexity.
-In Stage 1, the complexity of DND is $\mathcal{O}\left(KN\log_{2}{N}\right)$ as analyzed in [54, Remark 6].
+We begin by analyzing the complexity of each step of the proposed 2SDGT solution given in Algorithm 1, and finally sum everything up to determine the total complexity.
+The complexity of DND is $\color[rgb]{0,0,0}\mathcal{O}\bigl(K_{\max}N\log_{2}{N}\bigr)$}, as analyzed in [19, Remark 6].
 Then, for each positive entry of the test result vector $Y^{T}$, the DD algorithm counts the number of possibly defective items that participate in the corresponding pool test.
-That requires $K\left|\mathcal{P}^{(S_{1,1})}\right|\log_{2}{N}$ computations, and for simplification, we bound it by the DND complexity, i.e., $\mathcal{O}\left(KN\log_{2}N\right)$.
-In Stage 2, parallel LVA requires $L$ times more computations than the VA [32].
+This can be implemented with at most ${\color[rgb]{0,0,0}K_{\max}}\bigl|\mathcal{P}^{(S_{1,1})}\bigr|\log_{2}{N}$ operations, which we bound by the DND complexity, i.e., by $\color[rgb]{0,0,0}\mathcal{O}\bigl(K_{\max}N\log_{2}N\bigr)$}.
+Hence, the overall complexity of Stage 1 is $\color[rgb]{0,0,0}\mathcal{O}\bigl(K_{\max}N\log_{2}N\bigr)$}.
+Parallel LVA requires $L$ times more computations than the Viterbi Algorithm (VA) [56].
 The VA calculates all the possible transition probabilities for each step in the sequence.
-In GT, this sequence is the order’s items sequence, where with the suggested algorithm, it is enough to consider only the $\left|\mathcal{P}^{\left(S_{1,2}\right)}\right|$ items as the sequence steps.
-The optional states are basically either “non-defective” or “defective”, so there are four possible transitions in each step of the trellis proposed herein for GT.
-Nevertheless, this algorithm can be implemented to leverage additional memory to decide the state of each item based on the preceding $\tau$ items.
-Consequently, LVA takes $2^{2\tau}L\left|\mathcal{P}^{(S_{1,2})}\right|$ computations.
-For the average case, we use the expectation bounds from Theorems 1 and 3, such that:
-$\mathbb{E}\left[\left|\mathcal{P}^{\left(S_{1}\right)}\right|\right]=\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]-\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]\leq\left(N-K\right)P_{e,a}^{(DND)}+N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}K.$
-Accordingly, the number of the required computations for LVA step in MSGT is bounded by
-$2^{2\tau}L\left[\left(N-K\right)P_{e,a}^{(DND)}+N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}K\right].$ (6)
-For the worst-case, we assume that the DD step does not affect the possible detected items set, thus $\left|\mathcal{P}^{\left(S_{1,2}\right)}\right|\leq\left|\mathcal{P}^{(S_{1,1})}\right|$.
-To further bound the expression, we utilize both the error probability of the worst case as given in Theorem 2, thus $\left|\mathcal{P}^{\left(DND\right)}\right|\leq K+(N-K)P_{e,u}^{(DND)}$.
-By substituting Theorem 2 it follows that the number of possibly defective items in the worst-case is upper bounded by
-$\begin{split}K&+\left(N-K\right)\left(P_{e,a}^{(DND)}+\left(1+\frac{g^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1}\right),\end{split}$ (7)
-for $g=\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)$.
-Now, since the two expressions multiplying $(N-K)$ in (6) and (7) are error probabilities, both of them can be roughly bounded by one.
-Hence, the number of computations of LVA in the average case and worst-case becomes $\mathcal{O}(LN)$.
-In the next step in Stage 2, we filter the LVA results.
-We sum each sequence $\mathbf{Z_{l}}$ with a complexity of $\mathcal{O}(N)$.
-If the result is in the range $\left[K,\gamma K\right]$, we extract all the combinations of size $K-\left|\mathcal{P}^{(DD)}\right|$.
-Thus, this stage is done in at most $\mathcal{O}\left(L\left(N+\binom{\gamma K}{K-\left|\mathcal{P}^{(DD)}\right|}\right)\right)$ computations.
-Finally, in the MAP step of Stage 2, the algorithm goes over at most $L$ combinations of size $K-\left|\mathcal{P}^{(DD)}\right|$ out of no more than $\gamma K$ possibly defective items in each sequence.
-Then, for each combination, the group test is applied.
-Therefore, the MAP stage requires $\binom{L\gamma K}{K}KN\log_{2}{N}$ computations.
-Substituting the bounds of the binomial coefficient $\left(\frac{\gamma K}{K}\right)^{K}\leq\binom{\gamma K}{K}\leq\left(\frac{e\gamma K}{K}\right)^{K}$, it follows that the complexity of MAP stage in the proposed MSGT is $\mathcal{O}\left(L\gamma^{K}KN\log_{2}{N}\right)$.
-To conclude, the complexity of MSGT is the sum of all the steps, i.e., $\mathcal{O}\big{(}NK\log_{2}{N}+LN+L\binom{\gamma K}{K-\left|\mathcal{P}^{(DD)}\right|}+L\gamma^{K}KN\log_{2}{N}\big{)}$.
-As $N$ grows, the dominant term is the complexity of the MAP step.
-Thus, the complexity of the MSGT algorithm is bounded by $\mathcal{O}\left(\gamma^{K}KN\log_{2}{N}\right)$ operations, which completes the proof.
-∎
+In GT, this sequence is the items’ ordering, and with the proposed algorithm it is enough to consider only the $\bigl|\mathcal{P}^{(S_{1,2})}\bigr|$ items as the sequence length.
+The possible states are “non-defective” and “defective”, so there are four transitions in each step of the GT trellis.
+More generally, the algorithm can be implemented to leverage additional memory and decide the state of each item based on the preceding $\tau$ items.
+Consequently, LVA takes at most
+$2^{2\tau}L\bigl|\mathcal{P}^{(S_{1,2})}\bigr|$
+elementary operations. For the purpose of an upper bound on the complexity, it suffices to note that for every realization
+$0\,\leq\,\bigl|\mathcal{P}^{(DD)}\bigr|\,\leq\,\bigl|\mathcal{P}^{(S_{1,1})}\bigr|\,\leq\,N,$
+and hence
+$\bigl|\mathcal{P}^{(S_{1,2})}\bigr|\;=\;\bigl|\mathcal{P}^{(S_{1,1})}\bigr|-\bigl|\mathcal{P}^{(DD)}\bigr|\;\leq\;\bigl|\mathcal{P}^{(S_{1,1})}\bigr|\;\leq\;N.$
+Therefore, we have
+$2^{2\tau}L\bigl|\mathcal{P}^{(S_{1,2})}\bigr|\;\leq\;2^{2\tau}LN\;=\;\mathcal{O}(LN),$
+so the complexity of the LVA step in 2SDGT is $\mathcal{O}(LN)$.
+In the next step, we filter the LVA results.
+We sum each sequence $\mathbf{Z}_{\ell}$ with complexity $\mathcal{O}(N)$ per sequence.
+If the sum lies in the range $K_{\min},{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}]$, we extract all combinations of sizes $k-\bigl|\mathcal{P}^{(DD)}\bigr|$ for all $k\in[K_{\min},K_{\max}]$.
+Hence, this step requires at most
+$\color[rgb]{0,0,0}\mathcal{O}\!\left(L\left(N+\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k-\left|\mathcal{P}^{(DD)}\right|}\right)\right)$.}
+This term will be dominated by the subsequent MAP step.
+Finally, in the MAP step of Stage 2, the algorithm considers at most $L$ sequences.
+In each sequence, there are at most ${\color[rgb]{0,0,0}\lfloor}{\color[rgb]{0,0,0}\gamma K_{\max}}{\color[rgb]{0,0,0}\rfloor}$ possibly defective items, and we examine all combinations of sizes $k-\bigl|\mathcal{P}^{(DD)}\bigr|$ for $k\in[K_{\min},K_{\max}]$.
+Thus, the number of candidate supports per sequence is at most
+$\color[rgb]{0,0,0}\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k-\left|\mathcal{P}^{(DD)}\right|}\;\leq\;\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k}$.}
+For each candidate support, we apply the group test metric, whose computation can be implemented with $\color[rgb]{0,0,0}\mathcal{O}\bigl(K_{\max}N\log_{2}N\bigr)$} operations.
+Therefore, the complexity of the MAP stage is bounded by
+$\color[rgb]{0,0,0}L\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k}K_{\max}N\log_{2}N.$}
+Using the standard bounds
+$\Big(\tfrac{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{K_{\max}}\Big)^{K_{\max}}\;\leq\;\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{K_{\max}}\;\leq\;\Big(\tfrac{e{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{K_{\max}}\Big)^{K_{\max}},$
+we obtain
+$L\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k}K_{\max}N\log_{2}N\;=\;\mathcal{O}\bigl(L\gamma^{K_{\max}}K_{\max}N\log_{2}N\bigr).$
+Taking all stages together, the overall complexity of 2SDGT is
+$\mathcal{O}\Big(K_{\max}N\log_{2}N+LN+L\!\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k-\left|\mathcal{P}^{(DD)}\right|}+L\gamma^{K_{\max}}K_{\max}N\log_{2}N\Big).$
+As $N$ grows, this sum is dominated by the MAP term, and we conclude that the complexity of the 2SDGT algorithm is bounded by
+$\color[rgb]{0,0,0}\mathcal{O}\bigl(L\gamma^{K_{\max}}K_{\max}N\log_{2}N\bigr)$},
+which completes the proof. ∎
 Remark 4.
-If we skip the LVA step, MSGT converges to the MAP estimator. Thus, the MAP’s complexity is
-$\mathcal{O}\left(\displaystyle\binom{\left|\mathcal{P}^{(S_{1,2})}\right|}{K-\left|\mathcal{P}^{(DD)}\right|}KN\log_{2}N\right),$
-when the DND and DD are executed as prior steps, and otherwise it is $\mathcal{O}\left(\binom{N}{K}KN\log_{2}N\right)$.
-Note that from Theorem 4 and Remark 1, it follows that the proposed MSGT algorithm performs $\mathcal{O}\left(\frac{1}{L}\left(\frac{N}{\gamma K}\right)^{K}\right)$ times fewer computational operations compared to MAP.
+Under the interval-sparsity formulation $K_{\min}\leq K\leq K_{\max}$, skipping the LVA step results in an exhaustive MAP search over all defective sets with cardinality in $K_{\min},K_{\max}]$, whose complexity scales as
+$\mathcal{O}\!\left(\sum_{k=K_{\min}}^{K_{\max}}\binom{N}{k}\,K_{\max}N\log_{2}N\right),$
+which remains combinatorial in the effective sparsity level.
+Note that from Theorem 4 and Remark 4, it follows that the proposed 2SDGT algorithm performs
+$\mathcal{O}\!\left(({1}/{L})\left({N}/{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}\right)^{K_{\max}}\right),$
+times fewer computational operations compared to exhaustive MAP search over $K_{\min},K_{\max}]$.
+The MAP baseline reported herein corresponds to MAP decoding performed over the reduced candidate set obtained after the Stage 1 pruning step, rather than exhaustive MAP decoding over all $\binom{N}{K}$ supports.
+This choice reflects the fact that exhaustive MAP decoding is computationally infeasible for the problem dimensions considered and would not provide a meaningful practical benchmark.
+The reported complexity, therefore, accounts only for the enumeration and likelihood evaluation over the retained candidate supports, enabling a fair comparison with the proposed 2SDGT algorithm.
+The Bernoulli test-design parameter $p=\ln(2)/K_{\max}$ and the subsequent complexity and performance guarantees in this section are derived under a design sparsity interval $(K_{\min},K_{\max})$, rather than assuming knowledge of the realized sparsity $K$.
+This distinction is crucial: although the realized number of defectives enters the analysis of Stage 1 through concentration bounds, Stage 2 and the associated list-decoding and counting guarantees depend only on the design interval.
+Specifically, the LVA filtering rule in Algorithm 1 retains trajectories whose support size satisfies $K_{\min}\leq|V^{(\ell)}|\leq\lfloor\gamma K_{\max}\rfloor$, independently of the true realization of $K$.
+As a result, a moderate mismatch between the realized sparsity and the design upper bound affects performance only through the tail probability $\Pr(K<K_{\min})+\Pr(K>K_{\max})$, which is explicitly separated in the overall error bound in (1).
+Importantly, the counting argument in Proposition 1 and the complexity bound in Theorem 4 remain valid regardless of the exact value of $K$, as they depend solely on the candidate-set budget $\lfloor\gamma K_{\max}\rfloor$ and not on the decoder’s knowledge of the realized sparsity.
+This formulation enables 2SDGT to operate reliably under random or imperfectly known sparsity levels without redesigning the decoding pipeline, while preserving provable complexity control.
 III-C Discussion
-To the best of our knowledge, MSGT is the first GT algorithm to effectively leverage Markovian prior statistics.
-Unlike numerous previous approaches, MSGT utilizes initial probabilities and transition matrices without necessitating specific adjustments for new use cases.
+To the best of our knowledge, 2SDGT is the first GT algorithm to effectively leverage Markovian prior statistics.
+Unlike numerous previous approaches, 2SDGT uses the same decoding flow for any prior model, without requiring problem-specific redesign.
+In contrast, prior non-adaptive GT algorithms that incorporate structural assumptions—such as graph-constrained GT [17], model-based and graph-based GT [39], and contact-based GT [34]—must often adapt their test design or graph structure when the population model changes.
+2SDGT instead relies only on the given initial distribution and transition matrices, while the algorithmic pipeline remains unchanged.
 The algorithm offers the flexibility to be fine-tuned to optimize its performance in accordance with the available computational resources and the number of tests, $T$.
-The simple reduction of the search space in Stage 1 enables MSGT to handle challenging regimes with a small number of tests.
-Stage 2, and the LVA step in particular, contributes to its high success probability.
-Furthermore, using the parallel implementation of LVA, rather than the iterative one, keeps the complexity low [32].
+The simple reduction of the search space in Stage 1 enables 2SDGT to handle challenging regimes with a small number of tests.
+Stage 2, particularly the LVA step, contributes to its high success probability.
+Additionally, using the parallel implementation of LVA, rather than the iterative one, keeps the complexity low [56].
 It is important to note that, as explained in [56], achieving optimal results is ensured only with a very large $L$, inevitably leading to complexity equivalent to MAP’s.
 However, as we empirically demonstrate in the following section, results equivalent to MAP’s can be achieved with reasonable complexity.
-Moreover, it is shown that MSGT addresses practical regimes, e.g., in COVID-19 detection [55] ($T=48$ for $(N,K)=(384,4)$), in communications [56] ($(N,K)=(10^{5},6)$), and in GT quantizer [8] ($(N,K)=(1024,16)$ [57]).
-Another aspect of novelty in this work is the integration of the VA into the GT problem.
-In the context of Markovian priors, one can think of the population’s sequence of items as a sequence of observations stemming from a hidden Markov process within a given Markov model over $N$ steps.
+Moreover, it is shown that 2SDGT addresses practical regimes, e.g., in COVID-19 detection [57] ($T=48$ for $(N,K)=(384,4)$), in communications [45] ($(N,K)=(10^{5},6)$), and in GT quantizer [20] ($(N,K)=(1024,16)$ [36]).
+Another aspect of novelty in this research is the integration of the Viterbi Algorithm into the GT problem.
+In the context of Markovian priors, one can think of the population sequence of items as a sequence of observations stemming from a hidden Markov process within a given Markov model over $N$ steps.
 In that case, the selection of a Viterbi decoder becomes natural, offering an optimal and efficient decoding solution.
-However, the most likely sequence of items does not necessarily include $K$ defective items.
-Particularly in sparse signal scenarios, which is the focus of GT, the most likely sequence typically involves the minimum number of defectives that explain the observations.
-As a result, VA may not necessarily detect more defective items than already known and may detect even more than $K$.
+However, the most likely sequence of items does not necessarily include $K_{\max}$ defective items.
+Particularly, in sparse signal scenarios, which is the focus of GT, the most likely sequence typically involves the minimum number of defectives that explain the observations.
+As a result, VA may not necessarily detect more defective items than already known and may detect even more than $K_{\max}$.
 To address this, we employ LVA, which produces a list of the $L$ most likely sequences, such that choosing an appropriate value for $L$ guarantees a successful recovery.
-Like many previous works, MSGT assumes precise knowledge of $K$.
-In practical use cases, this assumption relies on using an accurate estimator for $K$ employing $\mathcal{O}\left(\log_{2}(N)\right)$ tests [58], but the estimation might be erroneous.
-For the sake of practical completeness, it is worth noting that modifications can be made to handle incorrect estimates of $K$, albeit with increased computational complexity.
-The authors of [4] suggest altering the ML estimator to consider all possible sets, without restricting the number of defective items, and show that the probability of success is almost unaffected.
-MSGT relies on knowing $K$ only for the MAP estimation step.
-Thus, a similar adaptation can be applied.
-In MSGT, the MAP estimator should consider all possible sets from LVA and cannot stop with the first set that explains the outcome $Y$.
-Hence, such modification will increase the complexity compared to [4].
+Unlike works that assume a deterministic or perfectly known $K$, the model adopted in this work allows the realized number of defectives $K$ to be random and unknown to the decoder.
+To accommodate this uncertainty without redesigning the algorithmic pipeline, 2SDGT operates with a design interval $(K_{\min},K_{\max})$, which is used consistently for test design, Stage 2 filtering, and candidate enumeration.
+Specifically, in Stage 2, the decoder retains only trajectories whose estimated support size satisfies $K_{\min}\leq|\mathbf{V}^{(l)}|\leq{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$, and enumerates candidate subsets of all cardinalities $k\in[K_{\min},K_{\max}]$ within these supports.
+This modification eliminates the need to know the exact realization of $K$ while preserving the counting-based guarantees and complexity bounds derived in this Section.
+The overall recovery error probability under random $K$ is controlled by conditioning on the typical event $\mathcal{E}_{\mathrm{typ}}\triangleq\{K_{\min}\leq K\leq K_{\max}\}$, and adding the probability of its complement.
+In particular,
+$P_{e}\;\leq\;P_{e}\big|\mathcal{E}_{\mathrm{typ}}\;+\;\Pr(K<K_{\min})\;+\;\Pr(K>K_{\max}).$
+When $\{U_{i}\}_{i=1}^{N}$ are generated by an ergodic finite-state Markov source (or, more generally, a strongly mixing process) [24, 29], the random variable $K=\sum_{i=1}^{N}U_{i}$ satisfies concentration bounds analogous to Chernoff inequalities [40, 49].
+That is, the concentration bounds invoked below hold under the additional assumption that the $\tau$-order Markov prior is ergodic (irreducible and aperiodic), or more generally strongly mixing; this assumption is standard in correlated-source models and is used solely to control the tail behavior of $K$.
+In particular, letting $\mu=\mathbb{E}[K]$, there exist constants $c_{1},c_{2}>0$ such that for any $0<\varepsilon\leq 1$, $\Pr\!\left(K>(1+\varepsilon)\mu\right)\leq c_{1}\exp\!\left(-c_{2}\,\varepsilon^{2}\,\mu\right)$, and $\Pr\!\left(K<(1-\varepsilon)\mu\right)\leq c_{1}\exp\!\left(-c_{2}\,\varepsilon^{2}\,\mu\right)$.
+Therefore, choosing $K_{\min}=\lfloor(1-\varepsilon)\mu\rfloor$, and $K_{\max}=\lceil(1+\varepsilon)\mu\rceil$, one can obtain that both tail probabilities decay exponentially in $\mu$.
+As a result, for sufficiently sparse regimes ($\mu=\mathcal{O}(N^{\theta})$ with $\theta<1$), the probability that $K$ falls outside the design interval $(K_{\min},K_{\max})$ is negligible, while 2SDGT retains its correctness and feasible computational complexity guarantees.
+If the above ergodicity or strong mixing assumptions do not hold, alternative concentration bounds or tail estimates for dependent random variables may be employed to control $\Pr(K<K_{\min})$ and $\Pr(K>K_{\max})$, albeit potentially with weaker decay rates.
 IV MAP Analytical Bound for GT with General Correlated Prior Statistics
-In this section, we provide a sufficiency bound for the number of required pooled tests for any MAP decoder with an arbitrary underlying correlation between infected items.
-This bound also applies to the LVA-based MSGT algorithm, which can approach MAP decoder performance for a sufficiently large list [32].
+In this section, we revisit the classical information-theoretic sufficiency condition for MAP decoding in group testing and adapt it to the setting considered here, which involves general correlated prior statistics and a constrained typical pattern in which only $K_{\min}\leq K\leq K_{\max}$ out of $N$ items are defective.
+This bound also applies to the LVA-based 2SDGT algorithm, which can approach MAP decoder performance for a sufficiently large list [56].
 We calculate this bound by extending the known result of ML decoding for GT problems [8], which bounds the probability of error of a ML decoder using a Gallager-type bound [29] for multiple access channel (MAC) channels [30].
 Recall that a noiseless GT setup can be considered as a noiseless MAC setting with $N$ “users”, each having a single codeword.
-While these type of bounds are well explored in the literature for MAC channels with correlated prior statistics [36, 37, 38, 39, 40, 41, 42], they do not expand naturally to GT, since GT can be thought as a special case where only “$K$” users transmit at any given point.
+While these type of bounds are well explored in the literature for MAC channels with correlated prior statistics [58, 69, 70, 12, 11, 51, 52], they do not expand naturally to GT, since GT can be thought as a special case where only “KK” users transmit simultaneously at any given point.
 This key difference necessitates dedicated analysis as follows in this section.
-In particular, to gain from the correlated prior statistics information, unlike ML decoders and bounds e.g., in [17, 14].
+In particular, to gain from the correlated prior statistical information, unlike ML decoders and bounds e.g., in [8, 53, 4].
+That is, unlike prior works that analyze ML decoding under a uniform prior over defective sets of fixed cardinality, our analysis considers MAP decoding with arbitrary, possibly correlated priors, requiring explicit weighting of competing defective sets according to their prior probabilities.
 We first introduce new notations and definitions in Subsection IV-A, then provide the probability of error to bound the MAP performance in Subsection IV-B, characterizing in Subsection IV-C the maximal number of tests required to guarantee reliability as given in Theorem 5.
-Finally, in Subsection IV-D, we evaluate this bound in Gilbert-Elliott (GE) model [60] and illustrate it for a practical regime in Section V.
+In Subsection IV-D, we evaluate this bound in Gilbert-Elliott model [33] and illustrate it for a practical regime in Section V.
+Throughout those subsection, for essay of notation and to follow the existing GT literature, we first analyze the performance of a MAP decoder conditioned on a fixed realization $K=k$.
+Accordingly, all quantities below (including the index set $\left[\binom{N}{K}\right$], the random variable $W$, and its distribution $P_{W}$) are interpreted under the conditional prior $P(\cdot\mid K=k)$, and the resulting bound applies to the conditional error probability $P_{e}(k)\triangleq\Pr\!\big(\widehat{\mathcal{K}}\neq\mathcal{K}\,\big|\,K=k\big)$.
+Finally, in Subsection IV-E, we extend the analysis to support the random-$K$ model (see Section II) with tail probability (See Section III-C).
 IV-A Definitions and notations
@@
 In the following, $W$ might be omitted for convenience of read.
-The $\omega$-th set is denoted by $S_{\omega}$.
+The $\omega$-th set is denoted by $S_{\omega}$ [^4] Note that there is a one-to-one mapping from $\omega$ to $S_{\omega}$, so in the MAP bound analysis herein, we use both $P_{W}\left(S_{\omega}\right)$ and $P_{W}\left(\omega\right)$ interchangeably..
 Note that $\lvert S_{\omega}\rvert=K$.
 The rows of $\mathbf{X}$ that correspond to the items in $S_{\omega}$ are denoted by $\mathbf{X}_{S_{\omega}}$.
-The output For a given $\omega\neq\omega^{*}$, the true positives (TP)/ false negatives (FN)/ false positives (FP) are denoted by $TP\left(\omega,\omega^{*}\right),FN\left(\omega,\omega^{*}\right),FP\left(\omega,\omega^{*}\right)$.
+For a given $\omega\neq\omega^{*}$, the true positives (TP)/ false negatives (FN)/ false positives (FP) are denoted by $TP\left(\omega,\omega^{*}\right),FN\left(\omega,\omega^{*}\right),FP\left(\omega,\omega^{*}\right)$.
 Note that $\lvert FN\left(\omega,\omega^{\prime}\right)\rvert=\lvert FP\left(\omega,\omega^{\prime}\right)\rvert$.
@@
 IV-C Sufficiency GT MAP Bound
-Let $(S_{2},S_{1})$ denote a partition of the defective set $S$ into disjoint sets $S_{2}$ and $S_{1}$, with cardinalities $i$ and $K-i$, respectively.
-This partition allows for categorizing error occurrences into distinct groups, where in group $i$, $K-i$ defective items are already identified correctly, and the primary error scenario arises from trying to detect the remaining $i$ defective items.
+Let $(S_{2},S_{1})$ denote a partition of the defective set $S$ into disjoint sets $S_{2}$ and $S_{1}$, with cardinalities $i$ and KiK-i, respectively.
+This partition allows for categorizing error occurrences into distinct groups, where in group $i$, KiK-i defective items are already identified correctly, and the primary error scenario arises from trying to detect the remaining $i$ defective items.
 Theorem 5.
@@
 then, as $N\rightarrow\infty$ the average error probability approaches zero.
+Remark 5.
+While Theorem 5 is stated for a MAP decoder conditioned on a fixed realization $K=k$, it forms the analytical foundation for the random-$K$ model adopted throughout the paper.
+As discussed in Section III-C and presented in Section IV-E, the overall error probability under random $K$ is controlled by conditioning on a typical sparsity interval and adding exponentially decaying tail probabilities.
+Finally, since the proposed 2SDGT algorithm approaches MAP performance for sufficiently large list sizes, this bound also characterizes the fundamental test-complexity regime in which 2SDGT operates effectively.
 The sufficiency proof of Theorem 5 is twofold.
@@
 Consider a group test with a MAP decoder for a population with correlated prior statistics. The error probability $P(E_{i}^{j})$ in (13) is bounded by
-$P(E_{i}^{j})\leq 2^{-T\left(E_{0}(\rho)-\frac{\frac{1}{\binom{K}{i}}E_{s,j}(\rho,P_{W})}{T}\right)},$
+$P(E_{i}^{j})\leq 2^{-T\Big(E_{0}(\rho)-{\color[rgb]{0,0,0}\left(1/\binom{K}{i}\right)E_{s,j}(\rho,P_{W})}/{T}\Big)},$
 where the error exponent $E_{0}(\rho)$ is given by
@@
 for $0\leq\rho\leq 1$, where $\mathcal{T}\mathcal{P}_{i}=\left\{S_{1}:\lvert S_{1}\rvert=K-i\right\}$, $\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)=\left\{S_{2}:\lvert S_{2}\rvert=i,\;S_{2}\cap S_{1}=\emptyset\right\}$, and $S_{\omega}=S_{1}\cup S_{2}$.
-Proof of Lemma 1: The proof is given in Appendix B. ∎
+Proof of Lemma 1: The proof is given in Appendix C. ∎
 Proof of Theorm 5.
@@
 For completeness, we show as in [8] that
-$\frac{\partial E_{o}}{\partial\rho}|_{\rho=0}$
-$=$ $\sum_{Y}\sum_{X_{TP}}[\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\log P\left(Y,X_{TP}|X_{FP}\right)-\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})]$
-$=$ $\sum_{Y}\sum_{X_{TP}}\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})\log\frac{p(Y,X_{TP}|X_{FP})}{\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})}$
-$=$ $I(X_{FP};X_{TP},Y).$
-And also, in Appedix C, that
-$\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}=H(P_{S_{2}|S_{1}}).$
-This is obtained directly from [62] as also elaborated in [42, Chap. 1]. Note that $E_{o}(0)=0$ and $E_{s,j}(0,P_{W})=0$, hence we have
+$\hskip-24.18501pt\frac{\partial E_{o}}{\partial\rho}|_{\rho=0}$
+$\hskip-19.91684pt=$ $\hskip-14.22636pt\sum_{Y}\sum_{X_{TP}}[\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\log P\left(Y,X_{TP}|X_{FP}\right)-\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})]$
+$\hskip-19.91684pt=$ $\hskip-14.22636pt\sum_{Y}\sum_{X_{TP}}\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})\log\frac{p(Y,X_{TP}|X_{FP})}{\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})}$
+$\hskip-19.91684pt=$ $\hskip-14.22636ptI(X_{FP};X_{TP},Y).$
+And also, in Appedix D, that
+$\hskip-24.18501pt\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}=H(P_{S_{2}|S_{1}}).$
+This is obtained directly from [59] as also elaborated in [52, Chap. 1]. Note that $E_{o}(0)=0$ and $E_{s,j}(0,P_{W})=0$, hence we have
 $T\emph{f}(\rho)=$ $T\rho\Bigg(I\left(X_{FP};X_{TP},Y\right)-\frac{H\left(P_{S_{2}|S_{1}}\right)}{T}\Bigg)-\log\frac{1}{\binom{K}{i}}+T\frac{\rho^{2}}{2}E_{o}^{\prime\prime}\left(\psi\right)-T\frac{\rho^{2}}{2}E_{s,j}^{\prime\prime}\left(\psi\right),$
@@
 for some $\varepsilon>0$, we have
-$T\emph{f}(\rho)\geq-\log\frac{1}{\binom{K}{i}}+T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)+\frac{\rho}{2}\left(E_{o}^{\prime\prime}(\psi)-E_{s,j}^{\prime\prime}\left(\psi\right)\right)\Bigg{)}$
+$\hskip-14.22636ptT\emph{f}(\rho)\geq-\log\frac{1}{\binom{K}{i}}+T\rho\Bigg(I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)+\frac{\rho}{2}\left(E_{o}^{\prime\prime}(\psi)-E_{s,j}^{\prime\prime}\left(\psi\right)\right)\Bigg)$
 Moreovver, note that $E_{o}^{\prime\prime}\left(\psi\right)$ is negative [29]. However, it is independent of the other constants and $T$, hence choosing
@@
 Claim 1 ([8]).
-For large $K$, and under a fixed input distribution for the testing matrix $\left(\frac{\ln(2)}{K},1-\frac{\ln(2)}{K}\right)$, the mutual information between $X_{FP}$ and $(X_{TP},Y)$ is bounded by
+Under a fixed input distribution for the testing matrix $\left(\frac{\ln(2)}{K},1-\frac{\ln(2)}{K}\right)$, the mutual information between $X_{FP}$ and $(X_{TP},Y)$ is bounded by
 $I(X_{FP};X_{TP},Y)\geq\frac{i}{K}.$
-Hence, by applying Clain 1 in Eq. (11), we achieves the sufficiency bound on $T$ as provided in Theorem 5,
+Claim 1 relies exclusively on the independence of the Bernoulli test design used to construct the testing matrix $X$.
+No independence is assumed for the defective-item prior.
+In particular, the result holds for arbitrary (possibly correlated) priors on $\mathbf{U}$, including the trellis-based model considered in this paper.
+All prior dependence is accounted for in the MAP decoding step through the enumeration and weighting of competing defective supports, while Claim 1 characterizes only the conditional likelihood of the test outcomes given a fixed support.
+Hence, by applying Claim 1 in Eq.
+(14), we achieves the sufficiency bound on $T$ as provided in Theorem 5,
 $(1+\varepsilon)\max_{i=1,\ldots,K}\frac{K}{i}H\left(P_{S_{2}|S_{1}}\right),$
@@
 (a) Upper bound for possibly defective items after DND for $N=10000$, and $K=15$.
-(b) Lower bound for definitely defective items after DD for $N=10000$, and $K=15$.
+(b) Lower bound for definitely defective items after $D$ for $N=10000$, and $K=15$.
 (c) Minimum $\gamma$ parameter to satisfies (9) for $N=500$, and $K=3$.
@@
 IV-D Gilbert Elliott calculations
-Consider a GE model [60].
+Consider a Gilbert-Elliott model [33].
 It is a Markov process with two states.
@@
 Recall that $P_{W}\left(S_{\omega}\right)$ calculates the probability to get $S_{\omega}$ as the infected set, given that there are $K$ infected items.
-Let $\mathcal{G}\mathcal{E}_{N}$ denote a GE process that runs for $N$ steps.
+Let $\mathcal{G}\mathcal{E}_{N}$ denote a Gilbert-Elliott process that runs for $N$ steps.
 Let $\left\{U_{i}\right\}_{i=1}^{N}$ denote the states of the process during the $N$ steps.
@@
 While our analysis is conditioned on the assumption that exactly $K$ items are infected, a general $\mathcal{G}\mathcal{E}_{N}$ process may result in any number of infected items between 0 and $N$.
-We therefore analyze and simulate a special class of GE processes; GE processes that reach the infected state exactly $K$ times after $N$ steps.
-Such a GE process that is guaranteed to produce exactly $K$ infected items is denoted by $\mathcal{G}\mathcal{E}_{N,K}$.
+We therefore analyze and simulate a special class of Gilbert-Elliott processes; Gilbert-Elliott processes that reach the infected state exactly $K$ times after $N$ steps.
+Such a Gilbert-Elliott process that is guaranteed to produce exactly $K$ infected items is denoted by $\mathcal{G}\mathcal{E}_{N,K}$.
 Then $P_{W}\left(S_{\omega}\right)$ of $\mathcal{G}\mathcal{E}_{N,K}$ is given as follows
 $P_{W}\left(S_{\omega}\right)$ $=$ $P\left(\mathcal{G}\mathcal{E}_{N,K}\text{ results in }S_{\omega}\right)$ (15)
-=$\overset{(a)}{=}$ $P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\lvert\;\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)$
-=$\overset{(b)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega},\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}$
-=$\overset{(c)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)},$
+$\overset{(a)}{=}$ $P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\lvert\;\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)$
+$\overset{(b)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega},\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}$
+$\overset{(c)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)},$
 where (a) follows from the definition of $\mathcal{G}\mathcal{E}_{N,K}$, (b) from the definition of conditional probability, and (c) follows since $S_{\omega}$ contains exactly $K$ infected items.
 IV-D1 Numerator of Eq.(15)
-The numerator evaluates the probability that a standard GE process results in a specific pattern of states, where state one is achieved only at the states in $S_{\omega}$.
-Recall that GE is a Markov chain, and so $P\left(S_{\omega}\right)=P\left(U_{1}\right)\cdot\prod_{i=2}^{N}P\left(U_{i}\lvert U_{i-1}\right)$.
+The numerator evaluates the probability that a standard Gilbert-Elliott process results in a specific pattern of states, where state one is achieved only at the states in $S_{\omega}$.
+Recall that Gilbert-Elliott is a Markov chain, and so $P\left(S_{\omega}\right)=P\left(U_{1}\right)\cdot\prod_{i=2}^{N}P\left(U_{i}\lvert U_{i-1}\right)$.
 This can be readily calculated using the definitions of $q,s,\pi$, as described at the beginning of this section
@@
 We create a new Markov chain that represents $\sum_{i=1}^{n}U_{i}$ as $n$ advances from $1$ to $N$.
-This Markov chain keeps track of how many times the GE process was in state 1.
+This Markov chain keeps track of how many times the Gilbert-Elliott process was in state 1.
 Hence this Markov chain has $2\left(N+1\right)$ states.
-State $n,n^{\prime}\in\left\{0,\ldots,N\right\}$ denotes a state where the GE process was in the infected state (state 1) $n$ times, and the current state is 0,1, respectively.
+State $n,n^{\prime}\in\left\{0,\ldots,N\right\}$ denotes a state where the Gilbert-Elliott process was in the infected state (state 1) $n$ times, and the current state is 0,1, respectively.
 An example of this Markov chain is shown in Fig.
-6.
+7.
 For convenience, the states $0^{\prime},N$ are also listed, even though they cannot be achieved.
-This Markov chain’s states can be arranged in a vector, where the $j$-th state represents $\sum_{i=1}^{N}U_{i}=\left\lfloor j/2\right\rfloor$, and even/odd $j$ represents the current state is not/is 00.
+This Markov chain’s states can be arranged in a vector, where the $j$-th state represents $\sum_{i=1}^{N}U_{i}=\left\lfloor j/2\right\rfloor$, and even/odd $j$ represents the current state is not/is 0.
 For instance, in the example of Fig.
-6, the states $6/7$ represent states 2/$2/2^{\prime}$ in the figure.
+7, the states $6/7$ represent states $2/2^{\prime}$ in the figure.
 The initial state vector is given as by $\left(1-\pi,0,0,\pi,0,\ldots,0\right)$.
@@
 (c) $N=10000$ and $K=13$.
-Figure 5: Success probability of MSGT, MAP, ML and DD over 1000 iterations.
+Figure 6: Success probability of 2SDGT, MAP, ML, SCOMP, and DD over 1000 iterations.
 A comparison to ML and MAP is not presented in (b) and (c), as they are not feasible for populations of those sizes due to the computational complexity burden.
@@
 $\left(1-\pi,0,0,\pi,0,\ldots,0\right)\cdot P^{N-1}.$ (16)
-To conclude the calculation, the probability to have exactly $K$ infected items after $N$ steps can be obtained by summing items $12K+$ 1 and $22K+$ 2 of the vector in (13).
-Figure 6: Markov chain for the number of infected items of a GE model, $N=4$.
-In the next section, we evaluate the proposed sufficient bound given in Theorem 5 for MAP decoder in the GE setting considered herein.
+To conclude the calculation, the probability to have exactly $K$ infected items after $N$ steps can be obtained by summing items $2K+1$ and $2K+2$ of the vector in (16).
+Figure 7: Markov chain for the number of infected items of a Gilbert-Elliott model, $N=4$.
+In the next section, we evaluate the proposed sufficient bound given in Theorem 5 for MAP decoder in the Gilbert-Elliott setting considered herein.
+IV-E Extension to random-$K$ with tail probability
+Here we extend the analysis of the sufficient bound for fixed defectives $K=k$ to the model given in Section II, where the defectives $K$ is a random variable.
+The analyses in the subsection before are conditional on the realization $K=k$ and yield a conditional MAP error bound $\Pr(\widehat{\mathcal{K}}\neq\mathcal{K}\mid K=k)$.
+By the law of total probability, the overall error probability satisfies
+$\Pr(\widehat{\mathcal{K}}\neq\mathcal{K})=\sum_{k=K_{\min}}^{K_{\max}}\Pr(K=k)\Pr(\widehat{\mathcal{K}}\neq\mathcal{K}\mid K=k)+\Pr\!\big(K\notin[K_{\min},K_{\max}]\big).$
+Consequently,
+$\Pr(\widehat{\mathcal{K}}\neq\mathcal{K})\leq\max_{K_{\min}\leq k\leq K_{\max}}\Pr(\widehat{\mathcal{K}}\neq\mathcal{K}\mid K=k)+\Pr\!\big(K\notin[K_{\min},K_{\max}]\big).$
+Hence, if the condition of Theorem 5 holds uniformly for all $k\in[K_{\min},K_{\max}]$ and the tail probability $\Pr(K\notin[K_{\min},K_{\max}])$ is negligible, then reliable recovery is guaranteed.
 V Numerical Evaluation
-This section assesses the performance of the proposed MSGT algorithm by numerical study.
+This section assesses the performance of the proposed 2SDGT algorithm by numerical study.
 First, in Subsection V-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III.
-Then, in Subsection V-2, we contrast the performance of MSGT with those of DD, ML, and MAP in a practical regime of $N$ and $K$.
-To generate the correlated prior information between adjacent items, we use the special GE process that reaches the infected state exactly $K$ times after $N$ steps, as described and analyzed in Section IV-D.
-The GE model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$.
-These characteristics align well with the inputs required by the MSGT algorithm.
-In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [43, 44], or in sparse signal recovery in signal-processing with fixed vector size of input samples [9, 10, 11, 13, 64]), we show that the low computational complexity MSGT algorithm can reduce the number of pool tests by at least $25\%$.
+Then, in Subsection V-2, we contrast the performance of 2SDGT with those of DD, ML, and MAP in a practical regime of $N$ and $K$.
+Although Section II adopts a random-$K$ model (with a design interval $(K_{\min},K_{\max})$), the simulations in this section are conducted for a fixed realized value of $K$ for comparison with the existing solutions in the GT literature [8, 4, 3].
+This is also consistent with the theory, since the analytical results can be applied conditionally on the event $\{K=k\}$ for a typical value $k\in[K_{\min},K_{\max}]$, while the probability of atypical realizations $\{K\notin[K_{\min},K_{\max}]\}$ is controlled separately via exponential tail bounds (Chernoff-type concentration) as discussed in Section II and Section III-C.
+To generate the correlated prior information between adjacent items, we use the special Gilbert-Elliott process that reaches the infected state exactly $K$ times after $N$ steps, as described and analyzed in Section IV-D.
+The Gilbert-Elliott model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$.
+These characteristics align well with the inputs required by the 2SDGT algorithm.
+In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [44, 9], or in sparse signal recovery in signal-processing with fixed vector size of input samples [32, 63, 1, 26, 42]), we show that the low computational complexity 2SDGT algorithm can reduce the number of pooled tests by at least $25\%$.
 (a) Number of computational operations.
-(b) Correlation matrix $\mathbf{\Phi}_{i}$ for $\tau=2$.
-(c) Long memory priors - exact recovery.
-Figure 7: (a) Number of computational operations in MSGT and MAP as given in Theorem 4 and Remark 1, respectively.
-(b) Example of correlation matrix $\mathbf{\Phi}_{i}$.
-(c) Probability of success of MSGT with exact prior statistics of 3-memory-steps Markov process, and with limited prior statistics assuming the Markov process has only one memory step.
+(b) Long memory priors - exact recovery.
+Figure 8: (a) Number of computational operations in 2SDGT and MAP as given in Theorem 4 and Remark 4, respectively.
+(b) Probability of success of 2SDGT with exact prior statistics of 3-memory-steps Markov process, and with limited prior statistics assuming the Markov process has only one memory step.
 $N=1024,K=8$, 1000 iterations.
@@
 In Fig.
-4(a) we show the concentration of $\left|\mathcal{P}^{(S_{1,1})}\right|$, as obtained from the simulation, along with the bound on its expectation and on the worst-case that were calculated in Theorems 1 and 2 respectively.
-Note that the worst-case scenario regarding MSGT is when LVA filters the correct set of defective items.
+5(a) we show the concentration of $\left|\mathcal{P}^{(S_{1,1})}\right|$, as obtained from the simulation, along with the bound on its expectation and the upper deviation bound that were calculated in Theorems 1 and 2 respectively.
+Under the random-$K$ model, these theoretical bounds are interpreted conditionally on the realized value of $K$ (as done here), while the probability that $K$ falls outside the design interval $(K_{\min},K_{\max})$ is negligible by concentration and is accounted for separately in the overall error analysis.
+Note that the worst-case scenario regarding 2SDGT is when LVA filters the correct set of defective items.
 That may happen if the number of possibly defective items exceeds the threshold $\gamma K$.
-Since in MSGT, we only allow this deviation from the average and ignore the case of exceeding this threshold, our upper bound for the worst-case does not cover all potential realizations of $\left|\mathcal{P}^{(S_{1,1})}\right|$.
+Since in 2SDGT, we only allow this deviation from the average and ignore the case of exceeding this threshold, our upper bound for the worst-case does not cover all potential realizations of $\left|\mathcal{P}^{(S_{1,1})}\right|$.
 Similarly, Fig.
-4(b) demonstrates the concentration of $\left|\mathcal{P}^{(DD)}\right|$, as acquired through simulation, and the lower bound on its expectation as given in Theorem 3.
+5(b) demonstrates the concentration of $\left|\mathcal{P}^{(DD)}\right|$, as acquired through simulation, and the lower bound on its expectation as given in Theorem 3.
 Fig.
-4(c) illustrates the numerically computed lower bound for $\gamma$, derived from the inequality provided in Proposition 1.
+5(c) illustrates the numerically computed lower bound for $\gamma$, derived from the inequality provided in Proposition 1.
 For this simulation, we fix $N=500$, $K=3$, and calculate $\gamma$ value for a specific range of $T$ relative to the upper bound of ML.
-As explained in Section III, our conjecture asserts that any value of $\gamma$ surpassing this lower bound guarantees that MSGT performance will be at least on par with that of ML.
-Therefore, whenever computational resources allow, it is advisable to choose the value of $\gamma$ corresponding to the lower bound.
-This approach was followed in the subsequent simulations, and the practical outcomes presented in Subsection V-2 provide empirical support for our conjecture.
-Fig.7(a) compares the number of potential combinations to be examined in the MAP step, with and without the execution of the LVA step in MSGT.
-It can be observed that the LVA step performs an extensive filtering process, which allows MSGT to remain feasible even when executing MAP is no longer possible, especially in a regime below ML upper bound.
+Fig.8(a) compares the number of potential combinations to be examined in the MAP step, with and without the execution of the LVA step in 2SDGT.
+It can be observed that the LVA step performs an extensive filtering process, which allows 2SDGT to remain feasible even when executing MAP is no longer possible, especially in a regime below ML upper bound.
 The converse of the GT problem with general prior statistics was developed by Li et al.
@@
 $\left(1-P_{e}\right)H(\mathbf{U})=\sum_{i=1}^{N}H\left(U_{i}\lvert U_{1},\ldots,U_{i-1}\right).$
-The GE model considered in our numerical evaluations is a stationary Markov chain with $\tau=1$.
-Thus, $P\left(U_{i}\lvert U_{0},\ldots,U_{i-1}\right)=P\left(U_{i}\lvert U_{i-1}\right)$, for $i\in\{2,\ldots,N\}$.
+The Gilbert-Elliott model considered in our numerical evaluations is a stationary Markov chain with $\tau=1$. Thus,
+$P\left(U_{i}\lvert U_{0},\ldots,U_{i-1}\right)=P\left(U_{i}\lvert U_{i-1}\right),\quad\text{for}\quad i\in\{2,\ldots,N\}.$
 Substituting those priors, it follows that the converse of our problem is
@@
 5 and Fig.
-5.
-In Fig 8, we compare the lower and the upper bounds with GE model.
+6.
+In Fig 9, we compare the lower and the upper bounds with Gilbert-Elliott model.
 The lower and upper bounds for GT with correlated prior statistics are given in (17) and in Section IV-D, respectively, while for GT without prior statistical information, the bounds are as given in [8].
-Figure 8: Lower and upper bounds for GT with GE model (left). MSGT and corresponding MAP upper bound for GT with GE model (right). For MAP and ML upper bounds $\varepsilon=0.6$.
+Figure 9: Lower and upper bounds for GT with GE model (left). 2SDGT and corresponding MAP upper bound for GT with GE model (right). For MAP and ML upper bounds $\varepsilon=0.6$.
 V-2 Algorithm Evaluation
-We demonstrate the performance of MSGT using simulation.
-The population is sampled from GE model, and the regime is $K=\mathcal{O}\left(N^{\theta}\right)$ with $\theta\leq 1/3$.
-The GE parameters serve as our prior statistics, but in practice, we ignore samples where the number of defective items does not match $K$.
+We demonstrate the performance of 2SDGT using simulation.
+The population is sampled from Gilbert-Elliott model, and the regime is $\mu=\mathcal{O}\left(N^{\theta}\right)$ with $\theta\leq 1/3$.
+The Gilbert-Elliott parameters serve as our prior statistics, but in practice, we ignore samples where the number of defective items does not match $K$.
+Equivalently, the reported success probabilities are conditional on a fixed realized $K$, which matches the way the analytical guarantees are applied on the typical event $K\in[K_{\min},K_{\max}]$ in the random-$K$ formulation.
 In addition, although in the complexity analysis, we considered Bernoulli encoder for simplification, here we use a near-constant column weight encoder that optimizes DND’s performance [37], with $\frac{\ln{2}}{K}T$ tests sampled randomly for each item.
-The parameter $\gamma$ was chosen to satisfy (3) and $L=500$ was chosen empirically.
+In all experiments, the parameter $\gamma$ is selected to satisfy a target upper bound $\eta$ on the probability that the number of surviving candidates after Stage 1 exceeds $\gamma K_{\max}$ (see (9)), and $L=500$ was chosen empirically.
+We fix $\eta=10^{-2}$, which was found to provide a good tradeoff between computational complexity and recovery performance across all tested regimes.
+The resulting value of $\gamma$ is computed according to the analytical bounds derived in Section III-B and is held fixed across Monte Carlo trials.
 Fig.
-5 compares MSGT to MAP, ML, and DD algorithms.
+6 compares 2SDGT to MAP, ML, DD, and SCOMP (an algorithm that starts with DD and sequentially marks extra items as defective, ensuring the result is a satisfying set [3]) algorithms.
 We run DND and DD before ML and MAP for feasible runtime and memory consumption.
@@
 Note that for $N=1024,10000$ it is no longer possible to compare the performance since ML and MAP become infeasible (Fig.
-5(b),5(c)).
-Finally, MSGT was also tested with more complex probabilistic models.
+6(b),6(c)).
+Finally, 2SDGT was also tested with more complex probabilistic models.
 We sample the population based on a 3-memory-steps Markov process ($\mathbf{\Phi}_{i}$ is a $8\times 2$ matrix for all $i$).
-We execute MSGT using these prior statistics and also execute it with limited prior statistics, assuming that the process has only one memory step ($\mathbf{\Phi}_{i}$ is $2\times 2$).
-For example of $\mathbf{\Phi}_{i}$ representing a 2-memory-steps process, see Fig 7(b).
+We execute 2SDGT using these prior statistics and also execute it with limited prior statistics, assuming that the process has only one memory step ($\mathbf{\Phi}_{i}$ is $2\times 2$).
+For example of $\mathbf{\Phi}_{i}$ representing a 2-memory-steps process, see Fig 1.
 The results are shown in Fig.
-7(c).
+8(b).
 It is evident that utilizing the prior of long memory improves the success probability by $10\%$ in this scenario.
-We do note again that for the practical regime tested as in [55, 56, 8, 57], i.e., $N=1024$ and $K=8$, it is no longer possible to compare the performance since, unlike the efficient proposed MSGT algorithm, ML and MAP decoders for GT become infeasible.
+We do note again that for the practical regime tested as in [57, 45, 20, 36], i.e., $N=1024$ and $K=8$, it is no longer possible to compare the performance since, unlike the efficient proposed 2SDGT algorithm, ML and MAP decoders for GT become infeasible.
 VI Conclusions and Future Work
-In this work, we propose a MAP decoder for GT problems with arbitrary correlated sources.
+In this work, we propose a MAP decoder for GT problems with general correlated priors that admit a trellis representation.
 We analyze the performance of the MAP decoders for this case and provide, to the best of our knowledge, the only analysis that goes below $T_{ML}$, as the existing solutions in the literature do not offer an efficient and general solution for this regime.
-Our proposed MSGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of potential defective combinations.
+Our proposed 2SDGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of potential defective combinations.
 This approach yields an efficient computational solution that, as demonstrated in our simulation result for practical scenarios, can approach the minimum number of tests as MAP algorithm.
@@
 Appendix A Recovery Algorithms
-For the completeness of our proposed Algorithm 1, we present here all the algorithms that we use as integral components within MSGT, and provide detailed explanations for each.
+For the completeness of our proposed Algorithm 1, we present here all the algorithms that we use as integral components within 2SDGT, and provide detailed explanations for each.
 A-A Definitely Not Defective (DND)
@@
 It systematically compares each column of the $\mathbf{X}$ with $Y^{T}$.
-If a specific column contains a value of $1$ while its corresponding test result is 00, we say that this column cannot explain the test result.
+If a specific column contains a value of $1$ while its corresponding test result is 0, we say that this column cannot explain the test result.
 Since each column represents one item in the population, if $\mathbf{X}(i)$ cannot explain $Y^{T}$, then DND marks this item as definitely non-defective.
@@
 Output: $\mathcal{P}^{(DND)}$
-1:$\mathcal{P}^{(DND)}\leftarrow\emptyset$
+[…]
 2:for $i$ s.t. $\mathbf{Y}_{i}=0$ do $\mathcal{P}^{(DND)}\leftarrow\mathcal{P}^{(DND)}\cup\{j\lvert X_{i,j}=1\}$
@@
 We skip the last step of DD in [3], and in our implementation, DD returns only the set of definitely defective items, denoted by $\mathcal{P}^{(DD)}$.
-Algorithm 3 Definite Defective (DD) [19]
+Algorithm 3 Definite Defective ($D$) [3]
 Input: $\mathbf{X},Y,\mathcal{P}^{(S_{1,1})}$
 Output: $\mathcal{P}^{(DD)}$
-1:$\mathcal{P}^{(DD)}\leftarrow\emptyset$
+[…]
 2:for $i$ s.t. $\mathbf{Y}_{i}=1$ do
@@
 A-C Priors Update
-The algorithm translates the information gathered during the first stage of MSGT into the state space.
-In particular, if an element is detected as defective, its probability of being defective is updated to $1$.
-Additionally, the transition probabilities are modified so that the transition to the defective state becomes deterministic.
+The updatePriors procedure incorporates the information obtained in Stage 1 into the prior model.
+When Stage 1 determines that an item is certainly defective or non-defective, we treat this as fixed information and update the prior accordingly.
+If the item lies in the first $\tau$ positions, we modify the initial distribution so that it assigns probability only to configurations consistent with the known value.
+If the item appears at a later position, we restrict the transition probabilities to this item to the detected state only.
 Algorithm 4 updatePriors
-Input: $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N},\mathcal{P}^{(S_{1,2})},\mathcal{P}^{(DD)}$
-Output: $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$
-1:for each $i\in\mathcal{P}^{(\mathrm{DD})}$ do
-2:    $\pi_{i}\leftarrow 0$
-3:    if $i\neq 1$ then
-4:         $\boldsymbol{\Phi}_{i-1}[:,0]\leftarrow\boldsymbol{0}$
-5:         $\boldsymbol{\Phi}_{i-1}[:,1]\leftarrow\boldsymbol{1}$
-6:    end if
-7:end for
-8:return $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$
+Input: $\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N},\mathcal{P}^{(S_{1,2})},\mathcal{P}^{(DD)}$
+Output: $\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}$
+1:for each $i\in\mathcal{P}^{(S_{1,2})}\cup\mathcal{P}^{(DD)}$ do
+2:  Let $u^{\star}\leftarrow 0$ if $i\in\mathcal{P}^{(S_{1,2})}$, otherwise $u^{\star}\leftarrow 1$.
+3:  if $i\leq\tau$ then
+$\pi^{\prime}(u_{1},\ldots,u_{\tau})\leftarrow\frac{\pi(u_{1},\ldots,u_{\tau})\,\mathbf{1}\{u_{i}=u^{\star}\}}{\sum_{(v_{1},\ldots,v_{\tau})}\pi(v_{1},\ldots,v_{\tau})\,\mathbf{1}\{v_{i}=u^{\star}\}}.$
+4:  else
+$\Phi^{\prime}_{i}(u_{i}\mid u_{i-1},\ldots,u_{i-\tau})\leftarrow\begin{cases}1,&u_{i}=u^{\star},\\ 0,&u_{i}\neq u^{\star},\end{cases}$
+5:  end if
+6:end for
 A-D List Viterbi Algorithm (LVA)
@@
 In the algorithm we suggest herein, the key differences are: (1) the population sequence in GT replaces the time sequence in classical LVA as given in [56], and (2) we use the aggregated sequence likelihood instead of the general cost function presented in the original paper.
-In particular, in the suggested algorithm, as we traverse the trellis diagram, we are iteratively maximizing the likelihood of the sequence representing the status of the population, tested with correlated and non-uniform prior statistics, denoted by $\Psi$ (see line 15).
+In particular, in the suggested algorithm, as we traverse the trellis diagram, we are iteratively maximizing the likelihood of the sequence representing the status of the population, tested with correlated and non-uniform prior statistics, denoted by $\Psi$ (see line 16).
 Algorithm 5 List Viterbi Algorithm [56]
-Input: $L,\tau,\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N},\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$
+Input: $L,\tau,{\color[rgb]{0,0,0}\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}}$
 Output: $\mathbf{Z}$}
@@
 2:for $s\leftarrow 1$ to $2^{\tau}$ do
-3:    $\Psi(1,s,1)\leftarrow\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}\left(s\right)$triangleright initial probability for each
-                                     state
-4:    $\xi\left(n,s,l\right)\leftarrow s$triangleright previous state of each state
-5:    for $l\leftarrow 2$ to $L$ do
-6:         $\Psi(1,s,l)\leftarrow 0$
-7:         $\xi\left(n,s,l\right)\leftarrow s$
-8:    end for
-9:end for
-10:Recursion:
-11:for $n\leftarrow 2$ to $N$ do
-12:    for $s_{2}\leftarrow 1$ to $2^{\tau}$ do
-13:         for $l\leftarrow 1$ to $L$ do
-14:             $\triangleright$ find the probability of the $l$-most likely
-                   previous state and its rank
-15:             $\Psi(n,s_{2},l)=\max^{(l)}_{{\begin{subarray}{c}s_{1}\in\{1,...,2^{\tau}\}\\ k\in\{1,..,L\}\end{subarray}}}\newline \hskip 103.50017pt\left\{\Psi(n-1,s_{1},k)\Phi_{n}\left(s_{1},s_{2}\right)\right\}$
-16:             $\left({s_{1}}^{*},k^{*}\right)=\arg\max^{(l)}_{{\begin{subarray}{c}s_{1}\in\{1,...,2^{\tau}\}\\ k\in\{1,..,L\}\end{subarray}}}\newline \hskip 103.50017pt\left\{\Psi(n-1,s_{1},k)\Phi_{n}\left(s_{1},s_{2}\right)\right\}$
-17:             $\xi\left(n,s_{2},l\right)\leftarrow{s_{1}}^{*}$
-18:             $\chi\left(n,l\right)\leftarrow k^{*}$
+3:  $\color[rgb]{0,0,0}\Psi(\tau,s,1)\leftarrow\pi(s)$} $\triangleright$ initial probability for each state
+4:  ${\color[rgb]{0,0,0}\xi(\tau,s,1)}\leftarrow s$ $\triangleright$ previous state of each state
+5:  $\color[rgb]{0,0,0}\chi(\tau,s,1)\leftarrow 1$}
+6:  for $l\leftarrow 2$ to $L$ do
+7:    $\Psi(\tau,s,l)\leftarrow 0$
+8:    ${\color[rgb]{0,0,0}\xi(\tau,s,l)}\leftarrow s$
+9:    $\chi(\tau,s,l)\leftarrow 1$
+10:  end for
+11:end for
+12:Recursion:
+13:for $n\leftarrow\tau+1$ to NN do
+14:  for $s_{2}\leftarrow 1$ to $2^{\tau}$ do
+15:    for $l\leftarrow 1$ to $L$ do
+16:
+$\Psi(n,s_{2},l)\;=\;\max\nolimits^{(l)}_{\begin{subarray}{c}s_{1}\in\{1,\dots,2^{\tau}\}\\ k\in\{1,\dots,L\}\end{subarray}}\Big\{\Psi(n-1,s_{1},k)\;\Phi_{n}(s_{1},s_{2})\Big\}$
+$(\,s_{1}^{\ast},\;k^{\ast}\,)\;=\;\arg\max\nolimits^{(l)}_{\begin{subarray}{c}s_{1}\in\{1,\dots,2^{\tau}\}\\ k\in\{1,\dots,L\}\end{subarray}}\Big\{\Psi(n-1,s_{1},k)\;\Phi_{n}(s_{1},s_{2})\Big\}$
+17:     $\xi(n,s_{2},l)\leftarrow s_{1}^{\ast}$
+18:     $\chi(n,s_{2},l)\leftarrow k^{\ast}$
 19:    end for
@@
 23:for $l\leftarrow 1$ to $L$ do
-24:    $\mathbf{Z}_{l,N}\leftarrow\arg\max_{s\in\{1,...,2^{\tau}\}}\left\{\Psi(N,s,l)\right\}$
-25:    $l_{N}\leftarrow\chi\left(n,l\right)$
-26:    for $n\leftarrow N-1$ to $1$ do
-27:         $\mathbf{Z}_{l,n}\leftarrow\xi\left(n+1,\mathbf{Z}_{l_{n+1},n+1},l_{n+1}\right)$
-28:         $l_{n}\leftarrow\chi\left(n+1,l_{n+1}\right)$
+24:  $Z_{l,N}\leftarrow\arg\max_{s}\Psi(N,s,l)$
+25:  $l_{N}\leftarrow{\color[rgb]{0,0,0}l}$
+26:  for $n\leftarrow N-1$ to $\tau$ do
+27:    $Z_{l,n}\leftarrow\xi\big(n+1,\;Z_{l,n+1},\;l_{n+1}\big)$
+28:    $l_{n}\leftarrow\chi\big(n+1,\;Z_{l,n+1},\;l_{n+1}\big)$
 29:  end for
@@
 In the recursion stage, we iterate from the second item to the last, and for each item, compute all the $2^{\tau}\times 2^{\tau}$ possible transition probabilities between all the possible states he could be in and the states of its predecessor.
-For each item, state, and rank, we set in $\Psi$ the $L$-most likely probabilities of the overall sequences from the first item until the current item and state (line 15), and we set in $\xi$ the corresponding previous state and in $\chi$ the corresponding rank (lines 17-18).
+For each item, state, and rank, we set in $\Psi$ the $L$-most likely probabilities of the overall sequences from the first item until the current item and state (line 16), and we set in $\xi$ the corresponding previous state and in $\chi$ the corresponding rank (lines 17-18).
 Here, $\max^{(l)}$ denotes the $l$-largest value.
@@
 If $\tau>1$, and a further processing is applied to map the states $\{1,...,2^{\tau}\}$ to the binary states $\{0,1\}$, representing “defective” and “non-defective”.
-A-E Maximum A Posteriori (MAP)
+A-E Candidate Enumeration via getAllCombinations
+In Stage 2 of 2SDGT (Algorithm 1), after obtaining a list of candidate population trajectories from the LVA, we must extract candidate defective sets to be evaluated by the MAP estimator.
+This task is performed by the procedure getAllCombinations, as shown in Algorithm 6.
+Given a set of indices $\mathbf{V}^{(l)}\subseteq\mathcal{N}$ corresponding to items estimated as defective in the $l$-th LVA trajectory, the goal of getAllCombinations is to enumerate all subsets whose cardinality lies within a prescribed range.
+This enables 2SDGT to handle uncertainty in the true number of defectives while keeping the candidate list size computationally feasible.
+Under the random-$K$ model in Section II, the algorithm does not assume knowledge of the realized $K$.
+Instead, Stage 2 operates with a design interval $(K_{\min},K_{\max})$, and getAllCombinations enumerates all subsets whose cardinality satisfies $K_{\min}\;\leq\;|c|\;\leq\;K_{\max},c\subseteq\mathbf{V}^{(l)}$.
+This modification ensures robustness to fluctuations in the realized number of defectives while preserving correctness on the typical event $\{K\in[K_{\min},K_{\max}]\}$.
+We note that the procedure is purely combinatorial and does not rely on probabilistic assumptions and that the upper bound ${\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$ ensures that the total number of candidate sets remains manageable, as analyzed in Section III.
+In practice, if $|\mathbf{V}^{(l)}|<K_{\min}$, the procedure returns the empty set, and the corresponding trajectory is discarded.
+Algorithm 6 getAllCombinations
+Input: $\mathbf{V}$, $K_{\min}$, $K_{\max}$
+Output: $\mathcal{C}$
+[…]
+2:for $k=K_{\min}$ to $K_{\max}$ do
+3:  for each subset $c\subseteq\mathbf{V}$ such that $|c|=k$ do
+4:    $\mathcal{C}\leftarrow\mathcal{C}\cup\{c\}$
+5:  end for
+6:end for
+7:return $\mathcal{C}$
+A-F Maximum A Posteriori (MAP)
 The MAP estimator [29], as shown in Algorithm 7, returns the set $c\in\mathbf{C}$ with the highest maximum a posteriori probability among all the sets, that is, the set $c$ that obtains the maximum $P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)$.
@@
 If the set $c$ and the given testing matrix $\mathbf{X}$ cannot explain for $Y$, then this probability is equal to zero.
-The second probability corresponds to the prior probability of the overall defective set $c$, calculated using $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N},\left\{\mathbf{\Phi}_{i}\right\}_{i=1}^{N}$.
+The second probability corresponds to the prior probability of the overall defective set $c$, calculated using $\color[rgb]{0,0,0}\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}$}.
 The MAP algorithm returns the estimated defective set, denoted by $\hat{\mathcal{K}}$.
 Algorithm 7 Maximum A Posteriori (MAP) [29]
-Input: $\mathbf{X},Y,\mathbf{C},\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N},\left\{\mathbf{\Phi}_{i}\right\}_{i=1}^{N}$
+Input: $\mathbf{X},Y,\mathbf{C},{\color[rgb]{0,0,0}\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}}$
 Output: $\hat{\mathcal{K}}$
 1:return $\arg\max_{c\in\mathbf{C}}{P\left(Y\lvert\mathbf{X},c\right)P\left(c\right)}$
-Appendix B Probability of Error (Proof of Lemma 1)
-Let $i\in\left[K\right],j\in\left[\binom{N-K}{i}\right]$. Let ~$\tilde{\mathcal{I}}^{j}_{dif}$ denote the set of all indices $\omega$ that result in a the set of FN of size $i$ as in $j$, i.e.
-~$\tilde{\mathcal{I}}^{j}_{dif}=\left\{\omega\in\left[\binom{N}{K}\right]:\quad FN\left(\omega,\omega^{*}\right)=\Omega_{i}\left(j\right)\right\}.$ (15)
+Appendix B Proof of Theorem 3
+Let $M_{0}$ denote the number of negative tests, and let $G$ denote the number of non-defective items that never appear in any negative test.
+We follow the $(M_{0},G)$ conditioning framework for DD given in [3, Appendix A.C].
+Throughout the proof, we condition on a fixed realization of the number of defectives $k$; the Bernoulli design parameter is $p=\ln(2)/K_{\max}$, but the argument below holds for any $p\in(0,1)$.
+For a fixed defective item $i$, let $L_{i}$ be the number of positive tests that: (i) contain item $i$, and (ii) contain no other item from $\mathcal{P}^{(S_{1,1})}$.
+We condition on $(M_{0},G)=(m_{0},g)$ and analyze a fixed defective item $i$.
+Let $\mathsf{Pos}$ denote the event that a given test is positive, and let $\mathsf{SoloDef}(i)$ denote the event that the test contains item $i$ and no other defective items.
+Under a Bernoulli(p)(p) design with $k$ defectives, $q_{0}\triangleq\Pr(\text{test negative})=(1-p)^{{\color[rgb]{0,0,0}k}}$ denotes the probability that a test is negative, and $q_{1}\triangleq\Pr(\mathsf{SoloDef}(i))=p(1-p)^{{\color[rgb]{0,0,0}k}-1}$ denotes the probability that a test contains defective $i$ and no other defectives, so that
+$a\triangleq\Pr(\mathsf{SoloDef}(i)\mid\mathsf{Pos})=\frac{\Pr(\mathsf{SoloDef}(i))}{\Pr(\mathsf{Pos})}=\frac{q_{1}}{1-q_{0}},$
+is the effective probability that a test includes item $i$ and no other defective items, given that the test is positive.
+Given $G=g$, the probability that none of the $g$ surviving non-defective candidates is included in a test is $(1-p)^{g}$.
+Hence, for each positive test, the probability that it contains item $i$ and no other item from the Stage 1 candidate set equals $a(1-p)^{g}$.
+Since there are $T-m_{0}$ positive tests and the Bernoulli design is independent across tests, it follows that
+$L_{i}\,\big|\,(M_{0}=m_{0},G=g)\sim\mathrm{Bin}\!\big(T-m_{0},\ a(1-p)^{g}\big).$
+Thus, we have
+$\color[rgb]{0,0,0}P(L_{i}>0\mid M_{0}=m_{0},G=g)=1-P(L_{i}=0\mid M_{0}=m_{0},G=g).$}
+For a binomial random variable, $P(L_{i}=0)=(1-\theta)^{n}$ with $n=T-m_{0}$ and $\theta=a(1-p)^{g}$, as detailed in Appendix B-A, we obtain
+$P(L_{i}>0\mid M_{0}=m_{0},G=g)\;=\;1-\bigl(1-a(1-p)^{g}\bigr)^{T-m_{0}}.$ (18)
+By symmetry of the items, the expected number of items identified by the DD step is
+$\mathbb{E}\big[\,|\mathcal{P}^{(\mathrm{DD})}|\,\big|\,K={\color[rgb]{0,0,0}k}\big]\;=\;{\color[rgb]{0,0,0}k}\,P(L_{1}>0).$ (19)
+We now apply Lemma 2, which states that for any $u\in[0,1]$ and integer $n\geq 1$
+$1-(1-u)^{n}\geq nu-\frac{n(n-1)}{2}u^{2}.$
+Applying this with $u=a(1-p)^{g}$ and $n=T-m_{0}$ to (18) yields
+$P(L_{1}>0\mid M_{0}=m_{0},G=g)$ $\geq(T-m_{0})\,a(1-p)^{g}-\frac{(T-m_{0})(T-m_{0}-1)}{2}\,a^{2}(1-p)^{2g}.$ (20)
+Taking expectation of both sides of (20) with respect to $(M_{0},G)$, substituting into (19), and by linearity, we obtain
+$\mathbb{E}\big[\,|\mathcal{P}^{(\mathrm{DD})}|\,\big|\,K={\color[rgb]{0,0,0}k}\big]$ $={\color[rgb]{0,0,0}k}\,\mathbb{E}\big[P(L_{1}>0\mid M_{0},G)\big]$
+$\geq{\color[rgb]{0,0,0}k}\left\{\,\mathbb{E}\big[(T-M_{0})a(1-p)^{G}\big]-\,\mathbb{E}\left[\frac{(T-M_{0})(T-M_{0}-1)}{2}\,a^{2}(1-p)^{2G}\right]\right\}$
+$={\color[rgb]{0,0,0}k}\left\{a\,\mathbb{E}\big[(T-M_{0})(1-p)^{G}\big]-\frac{a^{2}}{2}\,\mathbb{E}\big[(T-M_{0})(T-M_{0}-1)(1-p)^{2G}\big]\right\}.$ (21)
+Conditioned on $M_{0}=m_{0}$, each non-defective item appears in none of the $m_{0}$ negative tests with probability $(1-p)^{m_{0}}$. Hence, the number $G$ of such non-defectives follows
+$G\mid M_{0}=m_{0}\;\sim\;\mathrm{Bin}(N-{\color[rgb]{0,0,0}k},\,(1-p)^{m_{0}}).$
+Let $\rho=(1-p)^{m_{0}}$ denote this probability. Then, the probability generating function of a binomial variable implies
+$\mathbb{E}\big[(1-p)^{G}\mid M_{0}=m_{0}\big]=\mathbb{E}\big[(1-p)^{G}\big]=\left(1-p\cdot\rho\right)^{N-{\color[rgb]{0,0,0}k}}=\left(1-p(1-p)^{m_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}.$ (22)
+Similarly, using the identity $(1-p)^{2G}=\left((1-p)^{2}\right)^{G}$, we obtain
+$\mathbb{E}\big[(1-p)^{2G}\mid M_{0}=m_{0}\big]=\left(1-\big[1-(1-p)^{2}\big]\cdot\rho\right)^{N-{\color[rgb]{0,0,0}k}}=\left(1-(2p-p^{2})(1-p)^{m_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}.$ (23)
+Taking expectations over $M_{0}$ (using (22) and (23)), we obtain
+$\mathbb{E}\big[(T-M_{0})(1-p)^{G}\big]=\mathbb{E}\left[(T-M_{0})\left(1-p(1-p)^{M_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}\right],$
+and
+$\mathbb{E}\big[(T-M_{0})(T-M_{0}-1)(1-p)^{2G}\big]=\mathbb{E}\left[(T-M_{0})(T-M_{0}-1)\left(1-(2p-p^{2})(1-p)^{M_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}\right].$
+To simplify these expressions, we now apply the following bounds. First, for all $M_{0}\geq 0$, since $(1-p)^{M_{0}}\in[0,1]$ and $p>0$, we have
+$1-p(1-p)^{M_{0}}\geq 1-p.$
+Then, since $x\mapsto x^{N-{\color[rgb]{0,0,0}k}}$ is monotonically increasing on $[0,1]$, it follows that
+$\left(1-p(1-p)^{M_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}\geq(1-p)^{N-{\color[rgb]{0,0,0}k}}.$
+Therefore,
+$\mathbb{E}\big[(T-M_{0})(1-p)^{G}\big]$ $=\mathbb{E}\left[(T-M_{0})\left(1-p(1-p)^{M_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}\right]$
+$\geq\mathbb{E}[T-M_{0}]\cdot(1-p)^{N-{\color[rgb]{0,0,0}k}}=T(1-q_{0})(1-p)^{N-{\color[rgb]{0,0,0}k}},$ (24)
+For the second term, we simply upper-bound everything, such that, $(T-M_{0})(T-M_{0}-1)\leq T^{2}$ and $(1-p)^{2G}\leq 1$, so
+$\mathbb{E}\big[(T-M_{0})(T-M_{0}-1)(1-p)^{2G}\big]\leq T^{2}.$ (25)
+Substituting (24) and (25) into (B), we conclude the proof.
+B-A Proof of (18)
+Set $n\triangleq T-m_{0}$ and $\theta\triangleq a(1-p)^{g}$.
+Conditioned on $(M_{0},G)=(m_{0},g)$, there are exactly $n$ positive tests, and in each positive test, the event “test witnesses defective item ii” occurs with probability $\theta$, independently across tests.
+Therefore,
+$L_{i}\,\big|\,(M_{0}=m_{0},G=g)\sim\mathrm{Bin}(n,\theta).$
+By definition, the event $\{L_{i}=0\}$ means that none of the $n$ positive tests is a witness test for item $i$.
+Since these $n$ witness events are independent and each fails with probability $1-\theta$, we obtain
+$P(L_{i}=0\mid M_{0}=m_{0},G=g)$
+[…]
+$\overset{(b)}{=}\prod_{t=1}^{n}P(\text{no witness in test $t$}\mid M_{0}=m_{0},G=g)$
+$\overset{(c)}{=}\prod_{t=1}^{n}(1-\theta)$
+$=(1-\theta)^{n}$
+$=\bigl(1-a(1-p)^{g}\bigr)^{T-m_{0}},$
+where (a) follows since this is the definition of $L_{i}=0$, (b) due to the independence across tests, and (c) as each test is a witness w.p. $\theta$. Finally, we have
+$P(L_{i}>0\mid M_{0}=m_{0},G=g)=1-P(L_{i}=0\mid M_{0}=m_{0},G=g)=1-\bigl(1-a(1-p)^{g}\bigr)^{T-m_{0}},$
+which yields (18).
+B-B Binomial Bound
+Lemma 2.
+Let $u\in[0,1]$ and $n\in\mathbb{N}$. Then
+$1-(1-u)^{n}\;\geq\;nu-\frac{n(n-1)}{2}u^{2}.$
+Proof.
+The function $f(u)=(1-u)^{n}$ is analytic on $[0,1]$, and has a Taylor expansion around $u=0$
+$(1-u)^{n}=1-nu+\frac{n(n-1)}{2}u^{2}+R_{3}(u),$ (26)
+where $R_{3}(u)$ denotes the third-order remainder in the Taylor expansion. For $n\geq 3$, the third derivative of $f$ is
+$f^{(3)}(u)=-n(n-1)(n-2)(1-u)^{n-3}\leq 0\quad\text{for all }u\in[0,1].$
+By the Lagrange form of the remainder, there exists $\xi\in(0,u)$ such that
+$R_{3}(u)=\frac{f^{(3)}(\xi)}{3!}\,u^{3}\leq 0.$ (27)
+For $n=1,2$, the expansion is exact and $R_{3}(u)\equiv 0$. Now, we multiply the identity in (26) by 1-1 and reorder terms to obtain
+$1-(1-u)^{n}=nu-\frac{n(n-1)}{2}u^{2}-R_{3}(u).$
+Since $R_{3}(u)\leq 0$ for all $u\in[0,1]$ as given in (27), we have $R_{3}(u)\geq 0$, and therefore
+$1-(1-u)^{n}\geq nu-\frac{n(n-1)}{2}u^{2},$
+which completes the proof. ∎
+Appendix C Probability of Error (Proof of Lemma 1)
+Let $i\in\left[K\right],j\in\left[\binom{N-K}{i}\right$]. Let $\tilde{\mathcal{I}}^{j}_{dif}$ denote the set of all indices $\omega$ that result in a the set of FN of size $i$ as in $j$, i.e.
+$\tilde{\mathcal{I}}^{j}_{dif}=\left\{\omega\in\left[\binom{N}{K}\right]:\quad FN\left(\omega,\omega^{*}\right)=\Omega_{i}\left(j\right)\right\}.$ (28)
 Recall $\omega^{*}$ denotes the index of the set of the $K$ actual infected items.
@@
 For $S_{\omega^{\star}}=\left\{1,2,3\right\}$, let $j$ be the index such that $\Omega_{2}\left(j\right)=\left\{2,3\right\}$.
-In this case, ~$\tilde{\mathcal{I}}^{j}_{dif}$ has all indices that point to $\left\{1,4,5\right\},\left\{1,4,6\right\},\left\{1,4,7\right\},\left\{1,5,6\right\},\left\{1,5,7\right\},\left\{1,6,7\right\}$.
+In this case, $\tilde{\mathcal{I}}^{j}_{dif}$ has all indices that point to $\left\{1,4,5\right\},\left\{1,4,6\right\},\left\{1,4,7\right\},\left\{1,5,6\right\},\left\{1,5,7\right\},\left\{1,6,7\right\}$.
 Let $\zeta_{\omega^{\prime}},\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}$ denote the event that $S_{\omega^{\prime}}$ is more likely than $S_{\omega^{\star}}$ (see Eq.
@@
 $P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)=P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)=P\left(\mathbf{X}_{TP}\right).$
-By substituting Eq. (17) in Eq. (B), it follows that
-$P\left(E_{i}^{j}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$ (18)
-$\leq\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
-$=\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}$
-$\overset{(a)}{\leq}\left[\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}\right]^{\rho}$
-$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}},$
-$\qquad\forall s>0,\rho\in\left[0,1\right]$
+By substituting Eq. (30) in Eq. (C), it follows that
+$\hskip-8.5359ptP\left(E_{i}^{j}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$ (31)
+$\hskip-8.5359pt\leq\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
+$\hskip-8.5359pt=\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}$
+$\hskip-8.5359pt\overset{(a)}{\leq}\left[\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}\right]^{\rho}$
+$\hskip-8.5359pt=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}},$
+$\hskip-8.5359pt\hskip 18.49988pt\forall s>0,\rho\in\left[0,1\right]$
 where $(a)$ follows from the fact that the probability $P\left(E_{i}\lvert w^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)\leq 1$, hence it holds that if it is upperbounded by $U$, then $P\left(E_{i}\lvert w^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)\leq U^{\rho}$ for any $\rho\in[0,1]$.
 Using the law of total probability, we have
-$P\left(E_{i}^{j}\lvert W=\omega^{*}\right)$ (19)
-$=\sum_{\mathbf{X}_{S_{\omega^{\star}}}}\sum_{Y^{T}}P\left(\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
-$=\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(\mathbf{X}_{TP},Y^{T}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
-$\overset{(a)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)\cdot P\left(\mathbf{X}_{FN}\right)\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}}$
-$\overset{(b)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right))}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right))^{\frac{\rho}{1+\rho}}}$
-$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},$
+$\hskip-8.5359ptP\left(E_{i}^{j}\lvert W=\omega^{*}\right)$ (32)
+$\hskip-8.5359pt=\sum_{\mathbf{X}_{S_{\omega^{\star}}}}\sum_{Y^{T}}P\left(\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
+$\hskip-8.5359pt=\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(\mathbf{X}_{TP},Y^{T}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
+$\hskip-8.5359pt\overset{(a)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)\cdot P\left(\mathbf{X}_{FN}\right)\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}}$
+$\hskip-8.5359pt\overset{(b)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right))}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right))^{\frac{\rho}{1+\rho}}}$
+$\hskip-8.5359pt=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$\hskip-8.5359pt=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$\hskip-8.5359pt=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},$
 in $(a)$ we substitute Eq.
@@
 where
-$\begin{split}&\mathcal{T}\mathcal{P}_{i}=\left\{S_{1}:\lvert S_{1}\rvert=K-i\right\},\\ &\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)=\left\{S_{2}:\lvert S_{2}\rvert=i,\;S_{2}\cap S_{1}=\emptyset\right\},\\ &\text{and}\\ &\omega\in\left[\binom{N}{K}\right]\;s.t.\;S_{\omega}=S_{1}\cup S_{2}.\end{split}$
+$\begin{split}&\mathcal{T}\mathcal{P}_{i}=\left\{S_{1}:\lvert S_{1}\rvert=K-i\right\},\\ &\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)=\left\{S_{2}:\lvert S_{2}\rvert=i,\;S_{2}\cap S_{1}=\emptyset\right\},\\ &\hskip-36.98866pt\text{and}\\ &\omega\in\left[\binom{N}{K}\right]\;s.t.\;S_{\omega}=S_{1}\cup S_{2}.\end{split}$
 Proof.
@@
 The overall number of combinations is given by
-$\binom{N}{K-i}\cdot\binom{N-\left(K-i\right)}{i}$
+$\hskip-17.07182pt\binom{N}{K-i}\cdot\binom{N-\left(K-i\right)}{i}$
 $=$ $\frac{N!}{\left(K-i\right)!\left(N-\left(K-i\right)\right)!}\cdot\frac{\left(N-\left(K-i\right)\right)!}{i!\left(N-K\right)!}$
@@
 Now, by substituting Eq. (32) in Eq. (13), it follows that
-$P\left(E_{i}^{j}\right)=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)P\left(E_{i}^{j}|W=\omega^{*}\right)$
-$\leq\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{1-\frac{\rho}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-=$\overset{(a)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}\right]{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-$\overset{(b)}{\leq}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
-=$\overset{(c)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}\left[\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}\right]^{T}$
-=$\overset{(d)}{=}2^{-T\left(E_{0}(\rho)-\frac{\frac{1}{\binom{K}{i}}E_{s,j}(\rho,P_{W})}{T}\right)},$
+$\hskip-8.5359ptP\left(E_{i}^{j}\right)=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)P\left(E_{i}^{j}|W=\omega^{*}\right)$
+$\hskip-8.5359pt\leq\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$\hskip-8.5359pt=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{1-\frac{\rho}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$\hskip-8.5359pt=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$\hskip-8.5359pt\overset{(a)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}\right]{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$\hskip-8.5359pt\overset{(b)}{\leq}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
+$\hskip-8.5359pt\overset{(c)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}\left[\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}\right]^{T}$
+$\hskip-8.5359pt\overset{(d)}{=}2^{-T\left(E_{0}(\rho)-\frac{\frac{1}{\binom{K}{i}}E_{s,j}(\rho,P_{W})}{T}\right)},$
 where (a) follows from Lemma 3 with $\alpha=\frac{1}{1+\rho}$, (b) follows from the fact that for a fixed
@@
 This completed the error probability proof.
-Appendix C Derivative of $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$
+Appendix D Derivative of $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$
 In this section, we show the derivative of $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$ for $\rho=0$.
-This result is obtained directly from [62] as also elaborated in [42, Chap.
+This result is obtained directly from [59] as also elaborated in [52, Chap.
 1].
@@
 Thus,
-$\frac{\partial}{\partial\rho}\left(A\right)=A\frac{\partial}{\partial\rho}\left(\log\left(A\right)\right)$
-$=A\frac{\partial}{\partial\rho}\left(\left(1+\rho\right)\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)\right)$
-$=A\left(\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)+(1+\rho)\frac{\partial\tilde{f}\left(\rho,S_{1}\right)}{\partial\rho}\right)$
-=$\overset{(a)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\left(\log(\tilde{f}\left(\rho,S_{1}\right))+(1+\rho)\frac{\partial\tilde{f}\left(\rho,S_{1}\right)}{\partial\rho}\right)$
-=$\overset{(b)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
-=$\overset{(c)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
-=$\overset{(d)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
-where (a) follows replacing $A$, (b) by derivative of $\tilde{f}\left(\rho,S_{1}\right)$, and (c) replacing $\tilde{f}\left(\rho,S_{1}\right)$.
+$\hskip-8.5359pt\frac{\partial}{\partial\rho}\left(A\right)=A\frac{\partial}{\partial\rho}\left(\log\left(A\right)\right)$
+$\hskip-8.5359pt=A\frac{\partial}{\partial\rho}\left(\left(1+\rho\right)\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)\right)$
+$\hskip-8.5359pt=A\left(\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)+(1+\rho)\frac{\partial\tilde{f}\left(\rho,S_{1}\right)}{\partial\rho}\right)$
+$\hskip-8.5359pt\overset{(a)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\left(\log(\tilde{f}\left(\rho,S_{1}\right))+(1+\rho)\frac{\partial\tilde{f}\left(\rho,S_{1}\right)}{\partial\rho}\right)$
+$\hskip-8.5359pt\overset{(b)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\Biggr(\log(\tilde{f}\left(\rho,S_{1}\right))-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr)$
+$\hskip-8.5359pt\overset{(c)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{1+\rho}\Biggr(\log(\tilde{f}\left(\rho,S_{1}\right))-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr)$
+$\hskip-8.5359pt\overset{(d)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}\Biggr(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr).$
+where (a) follows replacing AA, (b) by derivative of $\tilde{f}\left(\rho,S_{1}\right)$, and (c) replacing $\tilde{f}\left(\rho,S_{1}\right)$.
 (d) follows by applying that $\log a-\frac{1}{\left(1+\rho\right)}\log b=\log\frac{a}{b^{\frac{1}{\left(1+\rho\right)}}}$.
@@
 Hence.
-from Eq.(C), we have
-$\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}=\frac{\partial\log\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A}{\partial\rho}=\frac{\partial\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A}{\partial\rho}$
-$=$ $\frac{1}{\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}$
+from Eq.(D), we have
+$\hskip-22.76228pt\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}=\frac{\partial\log\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A}{\partial\rho}=\frac{\partial\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A}{\partial\rho}$
+$=$ $\hskip 0.0pt\frac{1}{\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}$
 $\Biggr(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr).$
-Next, setting $\rho=0$ in Eq. (C), and since
+Next, setting $\rho=0$ in Eq. (D), and since
 $\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}A|_{\rho=0}=\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)=1,$
 we have
-$\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}$
+$\hskip-14.22636pt\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}$
 $=$ $\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\Biggr(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)}{P_{W}\left(S_{1},S_{2}\right)}\Biggr)$